### v2

# Insert-Only versus Insert-Delete in Dynamic Query Evaluation

Abstract.
We study the dynamic query evaluation problem: Given a natural join query $Q$ and a stream of updates, we would like to construct a data structure that supports constant-delay enumeration of the query output after each update.
We show that a stream of $N$ insert-only updates to an initially empty database can be executed in total time $\mathcal{O}(N^{\mathsf{w}(Q)})$, where $\mathsf{w}(Q)$ is the fractional hypertree width of $Q$. This matches the complexity of the static query evaluation problem for $Q$ and a database of size $N$. One corollary is that the amortized time per single-tuple insert is constant for acyclic joins.
In contrast, we show that a sequence of $N$ inserts and deletes to $Q$ can be executed in total time $\widetilde{\mathcal{O}}(N^{\mathsf{w}(\widehat{Q})})$, where $\widehat{Q}$ is obtained from $Q$ by extending every relational atom with extra variables that represent the “lifespans” of tuples in $Q$. We show that this reduction is optimal in the sense that the static evaluation runtime of $\widehat{Q}$ provides a lower bound on the total update time of $Q$. Our approach recovers the optimal single-tuple update time for known queries such as the hierarchical and Loomis-Whitney join queries.
incremental view maintenance; optimality; intersection joins
1. Introduction
Answering queries under updates to the underlying database, also called dynamic query evaluation, is a fundamental problem in data management, with recent work addressing it from both systems and theoretical perspectives. There have been recent efforts on building systems for dynamic query evaluation, such as DBToaster (Koch et al., 2014), DynYannakakis (Idris et al., 2017, 2020), and F-IVM (Nikolic and Olteanu, 2018; Nikolic et al., 2020; Kara et al., 2023b). By allowing tuples to carry payloads with elements from rings (Koch, 2010), such systems can maintain complex analytics over database queries, such as linear algebra computation (Nikolic et al., 2014), collection programming (Koch et al., 2016), and machine learning models (Kara et al., 2023b). Beyond conjunctive queries, there has been work on intersection joins under updates (Tao and Yi, 2022), and dynamic computation for expressive languages such as Datalog (Motik et al., 2019), Differential Datalog (Abadi et al., 2015), or DBSP (Budiu et al., 2023). The descriptive complexity of various recursive Datalog queries under updates, such as reachability, has also been investigated (Datta et al., 2015, 2018; Schwentick et al., 2020).
While this problem attracted continuous interest in academia and industry over the past decades, it is only relatively recently that the first results on the fine-grained complexity and in particular on the optimality of this problem have emerged. These efforts aim to mirror the breakthrough made by the introduction of worst-case optimal join algorithms in the database theory literature (Atserias et al., 2013; Veldhuizen, 2014; Ngo et al., 2018). Beyond notable yet limited explorations, understanding the optimality of maintenance for the entire language of conjunctive queries remains open. Prime examples of progress towards the optimality of query maintenance are the characterizations of queries that admit (worst-case or amortized) constant time per single-tuple update (insert or delete): the $q$-hierarchical queries (Berkholz et al., 2017; Idris et al., 2017; Kara et al., 2023b; Wang et al., 2023), queries that become $q$-hierarchical in the presence of free access patterns (Kara et al., 2023a) or by rewriting under functional dependencies (Kara et al., 2023b), or queries on update sequences whose enclosureness is bounded by a constant (Wang and Yi, 2020). The $\delta_{1}$-hierarchical queries (Kara et al., 2020b, 2023c, 2023a) and the triangle queries (Kara et al., 2019, 2020a) admit optimal, albeit non-constant, update time conditioned on the Online Matrix-Vector Multiplication (OMv) conjecture (Henzinger et al., 2015).
In this paper, we complement the above work in five ways.
First, we give reductions from the dynamic query evaluation problem to the static query evaluation problem and vice versa. These reductions use a translation of the time dimension in the dynamic problem into a spatial dimension in the static problem, so that the maintenance under a stream of updates corresponds to taking the intersection of intervals representing the “lifespans” of tuples in the update stream. Our reductions link the complexities of the dynamic and static problems and allow to transfer both algorithms and lower bounds from one problem to the other.
Second, we pinpoint the complexity gap between the insert-only and the insert-delete settings in the dynamic evaluation of any (full conjunctive) query $Q$. Given a stream of updates to an initially empty database, our goal is to execute these updates one by one, while supporting constant-delay enumeration of the query result after each update. In the insert-only setting, we show that a stream of inserts to $Q$ can be executed one insert at a time in the same overall time as if the entire stream was executed as one bulk update (which corresponds to static evaluation). In contrast, in the insert-delete setting, a stream of both inserts and deletes to $Q$ can be executed one update at a time in the same overall time as the static evaluation of another query $\widehat{Q}$, which is obtained by extending every relational atom in $Q$ with extra variables representing the “lifespans” of tuples of $Q$. This query $\widehat{Q}$ may have a higher complexity than $Q$. Our reduction from the maintenance of $Q$ to the static evaluation problem for $\widehat{Q}$ is nevertheless optimal: Lower bounds on the latter problem imply lower bounds on the former. This shows that the maintenance of $Q$ in the insert-delete setting can be more expensive than in the insert-only setting.
Third, we devise new dynamic query evaluation algorithms that use our reductions. They improve or recover best known update times, e.g., amortized constant update time for any acyclic query in the insert-only setting and for any hierarchical query in the insert-delete setting. They also achieve optimal $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$ amortized update time for Loomis-Whitney queries in the insert-delete setting, where $\mathcal{D}$ is the database size at update time.
Fourth, our algorithms support constant-delay enumeration of both the full query output and the change to the output (delta enumeration) after each update. In particular, they map an input stream of updates to an output stream of updates, so they are closed under updates. In practice, the change to the output after an update is often much smaller than the full output, thus using delta enumeration can be more efficient than full enumeration.
Fifth, we prove more general lower bounds on the update time for any join query $Q$. Prior works, e.g., (Berkholz et al., 2017; Kara et al., 2019), typically rely on fine-grained complexity conjectures, such as the OMv conjecture (Henzinger et al., 2015), to prove such lower bounds and only for specific queries. In contrast, our lower bounds are stated in terms of any lower bounds on the static query evaluation problem, regardless of what these lower bounds are. Hence, our lower bounds can be thought of as unconditional in the sense that they do not rely on any conjecture that might be true or false. They also apply to any query.
1.1. Motivating Examples
We exemplify our results for the following three join queries:
$Q_{p}(A,B,C,D)$ $=R(A,B)\wedge S(B,C)\wedge T(C,D)$
(1) $Q_{\triangle}(A,B,C)$ $=R(A,B)\wedge S(B,C)\wedge T(A,C)$
$Q_{4}(A,B,C,D)$ $=R(A,B,C)\wedge S(B,C,D)\wedge T(A,C,D)\wedge U(A,B,D)$
[…]
We show that $Q_{p}$ admits single-tuple updates: in amortized constant time in the insert-only setting; and in amortized $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$ time in the insert-delete setting ($\tilde{O}$ notation hides a $\text{\sf polylog}(|\mathcal{D}|)$ factor). Both times are optimal, the latter is conditioned on the OMv conjecture. The amortized constant time per insert does not contradict the non-constant lower bound for non-hierarchical queries without self-joins (Berkholz et al., 2017): that lower bound proof relies on a reduction from OMv that requires both inserts and deletes. Furthermore, it is non-trivial to achieve amortized constant time per insert, while also allowing constant-delay enumeration after each insert; such a result was only shown recently in (Wang et al., 2023). The lazy maintenance approach, which only updates the input data, is not sufficient to achieve this result. This is because, while it does indeed take constant insert time, it still requires the computation of the query output before the enumeration of the first output tuple. This computation takes linear time using a factorized approach (Kara et al., 2023b) and needs to be repeated for every enumeration request.
The triangle query $Q_{\triangle}$ and the Loomis-Whitney-4 query $Q_{4}$ need $\mathcal{O}(|\mathcal{D}|^{1/2})$ and $\mathcal{O}(|\mathcal{D}|^{1/3})$, respectively, amortized single-tuple insert times. The total times of $\mathcal{O}(|\mathcal{D}|^{3/2})$ and $\mathcal{O}(|\mathcal{D}|^{4/3})$, respectively, match the worst-case optimal times for the static evaluation of these two queries (Ngo et al., 2018). In the insert-delete setting, single-tuple updates take $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$ amortized time for both queries, and this is optimal based on the OMv conjecture.
2. Problem Setting
We consider natural join queries, or equivalently, full conjunctive queries and refer to them as queries. Sec. 4 defines them.
We investigate the incremental view maintenance (IVM) of a query $Q$ under single-tuple updates to an initially empty database $\mathcal{D}$. A single-tuple update is an insert or a delete of a tuple into a relation in $\mathcal{D}$. We consider four variants of the IVM problem depending on the update and enumeration mode. With regard to updates, we distinguish between the insert-only setting, where we only allow single-tuple inserts, and the insert-delete setting, where we allow both single-tuple inserts and single-tuple deletes. With regard to enumeration, we distinguish between the full enumeration and the delta enumeration, where, after each update, the full query result and respectively the change to the query result at the current time can be enumerated with constant delay. Sec. 4 introduces the notion of constant-delay enumeration. The four IVM variants are parameterized by a query $Q$ and take as input a database $\mathcal{D}$, whose relations are initially empty, and a stream $\delta_{1}\mathcal{D},\delta_{2}\mathcal{D},\ldots,\delta_{N}\mathcal{D}$ of $N$ single-tuple updates to $\mathcal{D}$. Neither $N$ nor the updates are known in advance. The updates arrive one by one. Let $\mathcal{D}^{(0)}\stackrel{{\scriptstyle\text{def}}}{{=}}\emptyset,\mathcal{D}^{(1)},\ldots,\mathcal{D}^{(N)}$ be the sequence of database versions, where for each $\tau\in[N]$, the database version $\mathcal{D}^{(\tau)}$ results from applying the update $\delta_{\tau}\mathcal{D}$ to the previous version $\mathcal{D}^{(\tau-1)}$.
The task of the first IVM variant is to support constant-delay enumeration of the full query result $Q(\mathcal{D}^{(\tau)})$ after each update:
Problem: $\text{IVM}^{\pm}[Q]$ Parameter: Query $Q$ Given: An initially empty database $\mathcal{D}$ and a stream $\delta_{1}\mathcal{D},\ldots,\delta_{N}\mathcal{D}$ of $N$ single-tuple updates to $\mathcal{D}$. $N$ is not known in advance. Task: Support constant-delay enumeration of $Q(\mathcal{D}^{(\tau)})$ after each update $\delta_{\tau}\mathcal{D}$
The task of the second IVM variant is to support constant-delay enumeration of the change to the query result after each update:
Problem: $\text{IVM}^{\pm}_{\delta}[Q]$ Parameter: Query $Q$ Given: An initially empty database $\mathcal{D}$ and a stream $\delta_{1}\mathcal{D},\ldots,\delta_{N}\mathcal{D}$ of $N$ single-tuple updates to $\mathcal{D}$. $N$ is not known in advance. Task: Support constant-delay enumeration of the change $\delta_{\tau}Q(\mathcal{D})$ after each update $\delta_{\tau}\mathcal{D}$
The change $\delta_{\tau}Q(\mathcal{D})$ is the difference between the query result right after update $\tau$ and the query result right after update $\tau-1$.
The other two variants, denoted as $\text{IVM}^{+}[Q]$ and $\text{IVM}^{+}_{\delta}[Q]$, are identical to $\text{IVM}^{\pm}[Q]$ and $\text{IVM}^{\pm}_{\delta}[Q]$, respectively, except that the updates $\delta_{\tau}\mathcal{D}$ are restricted to be single-tuple inserts (no deletes are allowed). In all four variants, we are interested in the update time, i.e., the time to process a single-tuple update.
Example 2.1 ().
Consider the triangle query $Q_{\triangle}$ from Eq. (1). Suppose we have the update stream of length $8$ given in the second column of Table 1. The third and fourth column show the full result and the delta result, respectively, after each update.
$\tau$ $\delta_{\tau}\mathcal{D}$ $Q(\mathcal{D}^{(\tau)})$ $\delta_{\tau}Q(\mathcal{D})$
[…]
3 $+T(a_{1},c_{1})$$\{(a_{1},b_{1},c_{1})\}$$\{+Q(a_{1},b_{1},c_{1})\}$
[…]
Table 1. A stream of 8 updates to $Q_{\triangle}$ in Eq. (1). The last two columns show the full and delta result after each update.
We obtain lower bounds on the update time for our IVM problems by reductions from the following static query evaluation problem:
Problem: $\text{Eval}[Q]$ Parameter: Query $Q$ Given: Database $\mathcal{D}$ Task: Compute $Q(\mathcal{D})$ For this problem, we are interested in the time to compute $Q(\mathcal{D})$.
Definition 2.2 ($\mathsf{w}(Q)$ and $\mathsf{\omega}(Q)$).
Given a query $Q$, let $\mathsf{w}(Q)$ denote the fractional hypertree width of $Q$ (see Section 4). Let $\mathsf{\omega}(Q)$ denote the smallest exponent $\alpha$ such that $\text{Eval}[Q]$ has an algorithm with runtime $\mathcal{O}(|\mathcal{D}|^{\alpha+o(1)}+|Q(\mathcal{D})|)$ on any input database $\mathcal{D}$. Given a union of queries $Q=\bigvee_{i}Q_{i}$, we define $\mathsf{w}(Q)$ and $\mathsf{\omega}(Q)$ to be the maximum of $\mathsf{w}(Q_{i})$ and $\mathsf{\omega}(Q_{i})$, respectively, over all $i$.  [^1] The function $\omega(Q)$ is defined analogously to the constant $\omega$ which is the best exponent for matrix multiplication.
The function $\mathsf{\omega}(Q)$ is not known in general. However, we know that $\mathsf{\omega}(Q)\leq\mathsf{w}(Q)$; see e.g. (Abo Khamis et al., 2016). More strongly, $\mathsf{\omega}(Q)$ is upper bounded by the submodular width of $Q$ (Marx, 2013; Abo Khamis et al., 2017)222 (Abo Khamis et al., 2017) provides a query evaluation algorithm with runtime $\mathcal{O}(|\mathcal{D}|^{\mathrm{subw}(Q)}\cdot\text{\sf polylog}{|\mathcal{D}|}+|Q(\mathcal{D})|)=\mathcal{O}(|\mathcal{D}|^{\mathrm{subw}(Q)+o(1)}+|Q(\mathcal{D})|)$, thanks to the extra $+o(1)$ in the exponent. $\mathrm{subw}(Q)$ denotes the submodular width of $Q$. . For acyclic queries, we have $\mathsf{\omega}(Q)=1$ (Yannakakis, 1981). For the triangle query, $\mathsf{\omega}(Q_{\triangle})\geq\frac{4}{3}$ modulo the 3SUM conjecture (Patrascu, 2010). In this paper, we prove several lower bounds on the IVM problems in terms of the function $\mathsf{\omega}(Q)$. These lower bounds are not conditioned on any fine-grained complexity conjecture (like the OMv conjecture (Henzinger et al., 2015), stated in Appendix A). Instead, they apply no matter what the value of $\mathsf{\omega}(Q)$ turns out to be. In that sense, they can be thought of as unconditional.
Complexity Measures
For all problems introduced above, we consider the query to be fixed. Moreover, we say that the amortized update time for the $\tau$-th single-tuple update is $f(N,|\mathcal{D}^{(\tau)}|)$ for some function $f$, if the total time to process all $N$ updates in the stream is upper bounded by $\sum_{\tau\in[N]}f(N,|\mathcal{D}^{(\tau)}|)$. In the insert-only setting, the database size $|\mathcal{D}^{(\tau)}|$ at time $\tau$ is always equal to $\tau$. Hence, for $\text{IVM}^{+}[Q]$ and $\text{IVM}^{+}_{\delta}[Q]$, we ignore the database size and only measure the update time as a function of the length $N$ of the update stream. In the insert-delete setting, the database size at time $\tau$ is upper bounded by $\tau$ and could be much smaller than $\tau$. As a result, for $\text{IVM}^{\pm}[Q]$ and $\text{IVM}^{\pm}_{\delta}[Q]$, it is more natural to ignore $N$ and only measure the update time as a function of $|\mathcal{D}^{(\tau)}|$. For the four IVM problems, the constant enumeration delay is not allowed to be amortized, regardless of whether the update time is amortized or not. For the problem Eval, we measure the computation time as a function of the size of the input database $\mathcal{D}$. We use classical $\mathcal{O}$-notation to express worst-case bounds and $\widetilde{\mathcal{O}}$-notation to hide a polylogarithmic factor in $N$, $|\mathcal{D}^{(\tau)}|$, or respectively $|\mathcal{D}|$.
Multivariate Extensions
For the purpose of stating our results, we introduce the following concept.
Definition 2.3 (Multivariate extension $\widehat{Q}$ of a query $Q$).
Consider a query $Q(\bm{X})=R_{1}(\bm{X}_{1})\wedge\ldots\wedge R_{k}(\bm{X}_{k})$ where $\bm{X}=\bigcup_{i\in[k]}\bm{X}_{i}$, and fresh variables $Z_{1},\ldots,Z_{k}$ that do not occur in $Q$. (In this paper, we drop head variables $\bm{X}$ and write $Q$ instead of $Q(\bm{X})$ for brevity.) Let $\Sigma_{k}$ be the set of permutations of the set $[k]$. For any permutation $\bm{\sigma}=(\sigma_{1},\ldots,\sigma_{n})\in\Sigma_{k}$, we denote by $\widehat{Q}_{\bm{\sigma}}$ the query that results from $Q$ by extending the schema of each atom $R_{\sigma_{i}}(\bm{X}_{\sigma_{i}})$ with the variables $Z_{1},\ldots,Z_{i}$:
$\widehat{Q}_{\bm{\sigma}}=\widehat{R}_{\sigma_{1}}(Z_{1},\bm{X}_{\sigma_{1}})$ $\wedge\widehat{R}_{\sigma_{2}}(Z_{1},Z_{2},\bm{X}_{\sigma_{2}})$
(2) $\wedge\cdots\wedge\widehat{R}_{\sigma_{k}}(Z_{1},\ldots,Z_{k},\bm{X}_{\sigma_{k}})$
The multivariate extension $\widehat{Q}$ of $Q$ is the union of $\widehat{Q}_{\bm{\sigma}}$ over all permutations $\bm{\sigma}\in\Sigma_{k}$:
(3) $\widehat{Q}=\bigvee_{\bm{\sigma}\in\Sigma_{k}}\widehat{Q}_{\bm{\sigma}}$
We call each $\widehat{Q}_{\bm{\sigma}}$ a component of $\widehat{Q}$.
Example 2.4 ().
The multivariate extension $\widehat{Q}_{\triangle}$ of $Q_{\triangle}$ from Eq. (1) consists of 6 components $\widehat{Q}_{123},\widehat{Q}_{132},\widehat{Q}_{213},\widehat{Q}_{231},\widehat{Q}_{312}$ and $\widehat{Q}_{321}$. The first component, $\widehat{Q}_{123}$, orders the input relations as $(R,S,T)$ and applies Eq. (2) by adding $\{Z_{1}\}$ to $R$, $\{Z_{1},Z_{2}\}$ to $S$, and $\{Z_{1},Z_{2},Z_{3}\}$ to $T$, thus resulting in $\widehat{R}(Z_{1},A,B)$, $\widehat{S}(Z_{1},Z_{2},B,C)$, and $\widehat{T}(Z_{1},Z_{2},Z_{3},A,C)$ respectively. Below, we add the subscript $123$ to each relation to indicate the component for later convenience:
(4) $\hskip-11.38092pt\widehat{Q}_{123}=\widehat{R}_{123}(Z_{1},A,B)\wedge\widehat{S}_{123}(Z_{1},Z_{2},B,C)\wedge\widehat{T}_{123}(Z_{1},Z_{2},Z_{3},A,C)$
The other components are given in Eq. (48).
3. Overview of Main Results
In this section, we overview our main results and discuss their implications. Recall that we use “query” to refer to a “natural join query” (or equivalently “full conjunctive query”). Also recall that $\mathsf{w}(Q)$ and $\mathsf{\omega}(Q)$ denote the fractional hypertree width of $Q$ and the smallest exponent for $\text{Eval}[Q]$, respectively, as per Definition 2.2. The notions of acyclicity, hierarchical, and Loomis-Whitney queries are reviewed in Sec. 4. Complexity measures are explained in Sec. 2. In our results, the constant enumeration delay is not amortized regardless of whether the update time is amortized. For all our IVM problems, $N$ is the number of single-tuple updates, while $|\mathcal{D}^{(\tau)}|$ is the database size at update time $\tau$.
3.1. Insert-Only Setting
In the insert-only setting, we show that off-the-shelf worst-case optimal join algorithms, e.g., LeapFrog TrieJoin (Veldhuizen, 2014), can be used to achieve the best known amortized update time for the dynamic query evaluation of arbitrary join queries, so including the cyclic queries, while supporting constant delay for full/delta enumeration. We give an upper bound on the single-tuple update time in terms of the fractional hypertree width $\mathsf{w}(Q)$ of $Q$:
Theorem 3.1 ().
[…]
Sec. 5 overviews our algorithm that meets the upper bound in Theorem 3.1. It involves worst-case optimal join algorithms and tree decompositions. Amortization is necessary as some inserts may be costly, while many others will necessarily be relatively cheaper, so the average insert cost matches the cost of computing a factorized representation of the query result divided by the number of inserts.
Since every acyclic query has a fractional hypertree width of one, Theorem 3.1 implies that every acyclic query can be maintained with amortized constant time per insert:
Corollary 3.2 (Theorem 3.1).
[…]
Corollary 3.2 recovers the amortized constant update time for acyclic joins from prior work (Wang et al., 2023). It also shows that the insert-only setting can be computationally cheaper than the insert-delete setting investigated in prior work (Berkholz et al., 2017): In the insert-delete setting, the non-hierarchical acyclic queries cannot admit $\mathcal{O}(|\mathcal{D}|^{1/2-\gamma})$ update time for any database $\mathcal{D}$ and $\gamma>0$ (Berkholz et al., 2017), conditioned on the Online Matrix-Vector Multiplication (OMv) conjecture.
We accompany the above upper bound with a lower bound on the insertion time in terms of the lower bound $\mathsf{\omega}(Q)$ on the static evaluation of $Q$. This lower bound is not conditioned on any fine-grained complexity conjecture (like OMv). Instead, it holds no matter what the value of $\omega(Q)$ is:
Theorem 3.3 ().
[…]
3.2. Insert-Delete Setting
In the insert-delete setting, our approach can maintain arbitrary join queries, and in particular any cyclic join query, with update times that can be asymptotically lower than recomputation. We give an upper bound on the update time for a query $Q$ in terms of the fractional hypertree width of the multivariate extension $\widehat{Q}$ of $Q$ (Definition 2.3):
Theorem 3.4 ().
[…]
As a corollary, we prove a stronger upper bound where $N$ is replaced by $|\mathcal{D}^{(\tau)}|$, the current database size at update time $\tau$. Note that $|\mathcal{D}^{(\tau)}|\leq N$ and could be unboundedly smaller than $N$:
Corollary 3.5 ().
[…]
Sec. 7.1 overviews our algorithm that meets the upper bound in Theorem 3.4. It involves intersection joins (Sec. 6). The following statements shed light on the relationship between a query $Q$ and its multivariate extension $\widehat{Q}$ (their proofs are deferred to Appendix E):
  • $Q$ is hierarchical if and only if its multivariate extension $\widehat{Q}$ is acyclic, which is equivalent to $\mathsf{w}(\widehat{Q})=1$.
  • If $Q$ is non-hierarchical, then $\mathsf{w}(\widehat{Q})\geq\frac{3}{2}$.
  • If $Q$ is Loomis-Whitney of any degree, then $\mathsf{w}(\widehat{Q})=\frac{3}{2}$.
  • For any query $Q$, we have $\mathsf{w}(Q)\leq\mathsf{w}(\widehat{Q})\leq\mathsf{w}(Q)+1$.
Immediate corollaries of these statements are that, in the insert-delete setting, our approach needs: (1) amortized $\widetilde{\mathcal{O}}(1)$ update time for hierarchical queries; and (2) amortized $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$ update time for the triangle query, which is the Loomis-Whitney query of degree 3. These update times mirror those given in prior work on $q$-hierarchical queries (Berkholz et al., 2017) and the full triangle query (Kara et al., 2020a), albeit our setting is restricted to full queries and initially empty databases and our update times are amortized and have a polylog factor.
A close analysis of the lower bound proof for non-hierarchical queries in prior work (Berkholz et al., 2017) reveals that for any non-hierarchical query $Q$ and any $\gamma>0$, there is no algorithm that solves $\text{IVM}^{\pm}[Q]$ or $\text{IVM}^{\pm}_{\delta}[Q]$ with $\widetilde{\mathcal{O}}(N^{1/2-\gamma})$ amortized update time, unless the OMv-conjecture fails (Proposition E.5 in Appendix E). Following Theorem 3.4 and the fact that $\mathsf{w}(\widehat{Q})=\frac{3}{2}$ for any Loomis-Whitney query $Q$, we conclude that both $\text{IVM}^{\pm}[Q]$ and $\text{IVM}^{\pm}_{\delta}[Q]$ can be solved with $\widetilde{\mathcal{O}}(N^{1/2})$ amortized update time, which is optimal, unless the OMv-conjecture fails.
[…]
We complement the upper bound on the update time in Theorem 3.4 with a lower bound in terms of the lower bound $\mathsf{\omega}(\widehat{Q}_{\bm{\sigma}})$ for the static evaluation of any component $\widehat{Q}_{\bm{\sigma}}$ of $\widehat{Q}$. This lower bound is not conditioned on any complexity conjecture. However, it only applies to the $\text{IVM}^{\pm}_{\delta}[Q]$ problem. Sec. 7.2 gives the high-level idea:
Theorem 3.6 ().
[…]
Since $|\mathcal{D}^{(\tau)}|\leq N$, the above immediately implies that $\text{IVM}^{\pm}_{\delta}[Q]$ cannot be solved with amortized update time $\widetilde{\mathcal{O}}(|\mathcal{D}^{(\tau)}|^{\mathsf{\omega}(\widehat{Q}_{\bm{\sigma}})-1-\gamma})$.
4. Preliminaries
We use $\mathbb{N}$ to denote the set of natural numbers including 00. For $n\in\mathbb{N}$, we define $[n]\stackrel{{\scriptstyle\text{def}}}{{=}}\{1,2,\ldots,n\}$. In case $n=0$, we have $[n]=\emptyset$.
Data Model
A schema is a tuple of variables. We treat schemas and sets of variables interchangeably, assuming a fixed ordering of variables. The domain of a variable $X$ is denoted by $\text{\sf Dom}(X)$. A value tuple $\bm{t}$ over schema $\bm{X}=(X_{1},\ldots,X_{n})$ is an element from $\text{\sf Dom}(\bm{X})\stackrel{{\scriptstyle\text{def}}}{{=}}\text{\sf Dom}(X_{1})\times\cdots\times\text{\sf Dom}(X_{n})$. We denote by $\bm{t}(X)$ the $X$-value of $\bm{t}$. A relation over schema $\bm{X}$ is a finite set of value tuples over the same schema. The size $|R|$ of a relation $R$ is the number of its tuples. A database $\mathcal{D}$ is a finite set of relations. Its size $|\mathcal{D}|$ is given by the sum of the sizes of its relations.
Queries
We consider natural join queries or full conjunctive queries. We refer to them as queries for short. A query has the form
(5) $Q(\bm{X})=R_{1}(\bm{X}_{1})\wedge\cdots\wedge R_{k}(\bm{X}_{k}),$
where each $R_{i}$ is a relation symbol, each $\bm{X}_{i}$ is a tuple of variables, and $\bm{X}=\bigcup_{i\in[k]}\bm{X}_{i}$. We call each $R_{i}(\bm{X}_{i})$ an atom of $Q$. We denote by $\text{\sf at}(Q)=\{R_{i}(\bm{X}_{i})\mid i\in[k]\}$ the set of all atoms of $Q$ and by $\text{\sf at}(X)$ the set of atoms $R_{i}(\bm{X}_{i})$ with $X\in\bm{X}_{i}$. The set of all variables of $Q$ is denoted by $\text{\sf vars}(Q)=\bigcup_{i\in[k]}\bm{X}_{i}$. Since the set $\bm{X}$ is always equal to $\text{\sf vars}(Q)$ for all queries we consider in the paper, we drop $\bm{X}$ from the query notation and abbreviate $Q(\bm{X})$ with $Q$ in Eq. (5). We say that $Q$ is without self-joins if every relation symbol appears in at most one atom.
A tuple $\bm{t}$ over schema $\bm{X}$ is in the result of $Q$ if there are tuples $\bm{t}_{i}\in R_{i}$ for $i\in[k]$ such that for all $X\in\text{\sf vars}(Q)$, the set $\{\bm{t}_{i}(X)\mid i\in[k]\wedge X\in\bm{X}_{i}\}$ consists of the single value $\bm{t}(X)$. Given a query $Q$ and a database $\mathcal{D}$, we denote by $Q(\mathcal{D})$ the result of $Q$ over $\mathcal{D}$.
A union of queries is of the form $Q=\bigvee_{i\in[n]}Q_{i}$ where each $Q_{i}$, called a component of $Q$, is a query over the same set of variables as $Q$. The result of $Q$ is the union of the results of its components.
Query Classes
Acyclic queries are reviewed in Appendix A.
Hierarchical queries are a sub-class of acyclic queries. A query is called hierarchical if for any two variables $X$ and $Y$, it holds $\text{\sf at}(X)\subseteq\text{\sf at}(Y)$, $\text{\sf at}(Y)\subseteq\text{\sf at}(X)$, or $\text{\sf at}(X)\cap\text{\sf at}(Y)=\emptyset$ (Suciu et al., 2011).
Loomis-Whitney queries generalize the triangle query from a clique of degree $3$ to higher degrees (Loomis and Whitney, 1949). A Loomis-Whitney query of degree $k\geq 3$ has $k$ variables $X_{1},\ldots,X_{k}$ and has, for each subset $\bm{Y}\subset\{X_{1},\ldots,X_{k}\}$ of size $k-1$, one distinct atom with schema $\bm{Y}$.
A union of queries is acyclic (hierarchical, Loomis-Whitney) if each of its components is acyclic (hierarchical, Loomis-Whitney). Example A.1 showcases the query classes mentioned above.
Single-tuple Updates and Deltas
We denote an insert of a tuple $\bm{t}$ into a relation $R$ by $+R(\bm{t})$ and a delete of $\bm{t}$ from $R$ by $-R(\bm{t})$. The insert $+R(\bm{t})$ inserts $\bm{t}$ into $R$ if $\bm{t}$ is not contained in $R$, and the delete $-R(\bm{t})$ deletes $\bm{t}$ from $R$ if it is contained in $R$. Each insert or delete has a timestamp $\tau$, which is a natural number that indicates the specific time at which the insert or delete occurs. An update $\delta_{\tau}R$ to $R$ at time $\tau$ is the set of all inserts and deletes to $R$ at time $\tau$. An update $\delta_{\tau}R$ is applied to $R$ by first applying all the deletes in $\delta_{\tau}R$ and then applying all the inserts. We use $R\biguplus\delta_{\tau}R$ to denote the result of applying an update $\delta_{\tau}R$ to a relation $R$. We use $R^{(\tau)}$ to refer to the specific version of a relation $R$ at timestamp $\tau$. In particular, $R^{(\tau)}$ is the result of applying $\delta_{\tau}R$ to the previous version $R^{(\tau-1)}$, i.e.  $R^{(\tau)}=R^{(\tau-1)}\biguplus\delta_{\tau}R$. An update $\delta_{\tau}\mathcal{D}$ to a database $\mathcal{D}$ is the set of all inserts and deletes to the relations in $\mathcal{D}$ at time $\tau$. Similarly, we use $\mathcal{D}^{(\tau)}$ to refer to the specific version of $\mathcal{D}$ at time $\tau$, i.e., $\mathcal{D}^{(\tau)}=\mathcal{D}^{(\tau-1)}\biguplus\delta_{\tau}\mathcal{D}$. If $\delta_{\tau}\mathcal{D}$ consists of a single insert or delete, i.e., $|\delta_{\tau}\mathcal{D}|=1$, it is called a single-tuple update. In particular, if it consists of a single insert (or delete), we call it a single-tuple insert (or single-tuple delete). Given a query $Q$ over a database $\mathcal{D}$, we use $\delta_{\tau}Q(\mathcal{D})$ to denote the change in $Q(\mathcal{D})$ at time $\tau$. Specifically, $\delta_{\tau}Q(\mathcal{D})$ is the difference between $Q(\mathcal{D}^{(\tau)})$ and $Q(\mathcal{D}^{(\tau-1)})$:
$\delta_{\tau}Q(\mathcal{D})\stackrel{{\scriptstyle\text{def}}}{{=}}$$\{+Q(\bm{t})\mid\bm{t}\in Q(\mathcal{D}^{(\tau)})\text{ and }\bm{t}\notin Q(\mathcal{D}^{(\tau-1)})\}$
(6) $\cup$ $\{-Q(\bm{t})\mid\bm{t}\in Q(\mathcal{D}^{(\tau-1)})\text{ and }\bm{t}\notin Q(\mathcal{D}^{(\tau)})\}$
When the time $\tau$ is clear from the context, we drop the annotation $(\tau)$ and write $R,\delta R,\mathcal{D},\delta\mathcal{D},Q(\mathcal{D})$, and $\delta Q(\mathcal{D})$. In the insert-only setting, since all deltas are positive, we ignore the signs and see a single-tuple update as just a tuple, $\delta R$ as just a relation, and so on.
Constant-Delay Enumeration
Given a query $Q$ and a database $\mathcal{D}$, an enumeration procedure for $Q(\mathcal{D})$ (or $\delta_{\tau}Q(\mathcal{D})$) outputs the elements in $Q(\mathcal{D})$ (or $\delta_{\tau}Q(\mathcal{D})$) one by one in some order and without repetition. The enumeration delay of the procedure is the maximum of three times: the time between the start of the enumeration process and the output of the first element, the time between outputting any two consecutive elements, and the time between outputting the last element and the end of the enumeration process (Durand and Grandjean, 2007). The enumeration delay is constant if it does not depend on $|\mathcal{D}|$.
Width Measures
The AGM-bound and fractional hypertree width of join queries are defined as usual. Their definitions are deferred to Appendix A. If a query is a union of join queries, its AGM-bound (or fractional hypertree width) is the maximum AGM-bound (or fractional hypertree width) over its components.
5. IVM: Insert-Only Setting
In this section, we give an overview of our algorithm for $\text{IVM}^{+}[Q]$ that meets the upper bound of Theorem 3.1. We leave the details and proofs (including the extension to $\text{IVM}^{+}_{\delta}[Q]$) to Appendix B. We start with the following lemma: (Recall notation from Sec. 4.)
Lemma 5.1 ().
Given a query $Q$, an initially empty database $\mathcal{D}^{(0)}$, and a stream of $N$ single-tuple inserts, we can compute the new output tuples $\delta_{\tau}Q(D)$ after every insert $\delta_{\tau}\mathcal{D}$, where the total computation time over all inserts is $\mathcal{O}(N+\text{\sf AGM}(Q,\mathcal{D}^{(N)}))$.
Note that the total computation time above is the same as the AGM bound (Atserias et al., 2013) of $Q$ over the final database $\mathcal{D}^{(N)}$. In particular, even in the static setting where $\mathcal{D}^{(N)}$ is given upfront, the output size $|Q(\mathcal{D}^{(N)})|$ could be as large as $\text{\sf AGM}(Q,\mathcal{D}^{(N)})$ in the worst-case. Moreover, worst-case optimal join algorithms cannot beat this runtime (Ngo et al., 2018; Veldhuizen, 2014; Ngo, 2018). The above lemma is proved based on the query decomposition lemma (Ngo et al., 2014; Ngo, 2018), which says the following. Let $Q$ be a query and let $\bm{Y}\subseteq\text{\sf vars}(Q)$. Then, the AGM bound of $Q$ can be decomposed into a sum of AGM bounds of “residual” queries: one query $Q\Join\bm{y}$ for each tuple $\bm{y}$ over the variables $\bm{Y}$. Appendix B gives a proof of the query decomposition lemma and an example.
Example 5.2 (for Lemma 5.1).
Consider the triangle query $Q_{\triangle}$ from Eq. (1). Lemma 5.1 says that given a stream of $N$ single-tuple inserts into $R,S$, and $T$, we can compute the new output tuples after every insert in a total time of $\mathcal{O}(N^{3/2})$. To achieve this, we need to maintain two indices for $R(A,B)$: one index that sorts $R$ first by $A$ and then by $B$, while the other sorts $R$ first by $B$ and then by $A$. For every insert to $R$, we update the two indices simultaneously. Similarly, we maintain two indices for each of $S$ and $T$.
Following notation from Sec. 4, let $R^{(0)}\stackrel{{\scriptstyle\text{def}}}{{=}}\emptyset,R^{(1)},\ldots,R^{(N)}$ be the sequence of versions of $R$ after each insert, and the same goes for $S$ and $T$. Suppose that the $\tau$-th insert in the stream is inserting a tuple $(a,b)$ into $R(A,B)$. The new output tuples that are added due to this insert correspond to the output of the following query: (Below, we drop the database instance $\mathcal{D}$ from the notation $\delta_{\tau}Q_{\triangle}(\mathcal{D})$ since $\mathcal{D}$ is clear from the context.)
$\delta_{\tau}Q_{\triangle}(A,B,C)\stackrel{{\scriptstyle\text{def}}}{{=}}\sigma_{B=b}S^{(\tau)}(B,C)\wedge\sigma_{A=a}T^{(\tau)}(A,C)$
The output size of this query is upper bounded by:
$\min(|\sigma_{B=b}S^{(\tau)}(B,C)|,|\sigma_{A=a}T^{(\tau)}(A,C)|)\leq$
(7) $\sqrt{|\sigma_{B=b}S^{(\tau)}(B,C)|\cdot|\sigma_{A=a}T^{(\tau)}(A,C)|}$
Moreover, the above query can be computed in time within $\mathcal{O}(1)$ factor from the quantity in Eq. (7). To achieve this time, we need to use the index for $S(B,C)$ that is indexed by $B$ first, and we also need the index for $T(A,C)$ that is indexed by $A$ first. Because $S^{(\tau)}\subseteq S^{(N)}$ and $T^{(\tau)}\subseteq T^{(N)}$, the quantity in Eq. (7) is bounded by:
(8) $\sqrt{|\sigma_{B=b}S^{(N)}(B,C)|\cdot|\sigma_{A=a}T^{(N)}(A,C)|}$
The query decomposition lemma (specifically Eq. (39) in Example B.2) says that the sum of the quantity in Eq. (8) over all $(a,b)\in R^{(N)}$ is bounded by $\sqrt{|R^{(N)}|\cdot|S^{(N)}|\cdot|T^{(N)}|}\leq N^{3/2}$. Hence, all inserts into $R$ take time $\mathcal{O}(N^{3/2})$.
Now suppose that the $\tau$-th insert in the stream is inserting a tuple $(b,c)$ into $S(B,C)$. To handle this insert, we compute the query:
$\delta_{\tau}Q_{\triangle}(A,B,C)\stackrel{{\scriptstyle\text{def}}}{{=}}\sigma_{B=b}R^{(\tau)}(A,B)\wedge\sigma_{C=c}T^{(\tau)}(A,C)$
To compute this query in the desired time, we need to use the index for $T(A,C)$ that is indexed by $C$ first. This is why we need to maintain two indices for $T(A,C)$. The same goes for $R$ and $S$.
Instead of the AGM-bound, the upper bound in Theorem 3.1 is given in terms of the fractional hypertree width of $Q$. To achieve this bound, we take an optimal tree decomposition of $Q$ and maintain a materialized relation for every bag in the tree decomposition using Lemma 5.1. In order to support constant-delay enumeration of the output, we “calibrate” the bags by semijoin reducing adjacent bags with one another. The following example illustrates this idea.
Example 5.3 (for Theorem 3.1).
Suppose we want to solve $\text{IVM}^{+}[Q]$ for the following query $Q$ consisting of two adjacent triangles: One triangle over $\{A,B,C\}$ and the other over $\{B,C,D\}$.
(9) $Q(A,B,C,D)=R(A,B)\wedge S(B,C)\wedge T(A,C)\wedge U(B,D)\wedge V(C,D)$
The above query $Q$ has a fractional hypertree width of $\frac{3}{2}$. In particular, one optimal tree decomposition of $Q$ consists of two bags: One child bag $B_{1}=\{A,B,C\}$ and another root bag $B_{2}=\{B,C,D\}$. Theorem 3.1 says that given a stream of $N$ single-tuple inserts, $Q$ can be updated in a total time of $\mathcal{O}(N^{3/2})$ where we can do constant-delay enumeration of the output after every insert. To achieve this, we maintain the following query plan:
$Q_{1}(A,B,C)$ $=R(A,B)\wedge S(B,C)\wedge T(A,C)$
$P_{1}(B,C)$ $=Q_{1}(A,B,C)$
$Q_{2}^{\prime}(B,C,D)$ $=P_{1}(B,C)\wedge S(B,C)\wedge U(B,D)\wedge V(C,D)$
By Lemma 5.1, $Q_{1}$ above can be updated in a total time of $\mathcal{O}(N^{3/2})$, as shown in Example 5.2. Also, $P_{1}$ can be updated in the same time because it is just a projection of $Q_{1}$. Regardless of the size of $P_{1}$, the query $Q_{2}^{\prime}$ can be updated in a total time of $\mathcal{O}(N^{3/2})$ because its AGM bound is upper bounded by the input relations $S,U$ and $V$. To do constant-delay enumeration of $Q(A,B,C,D)$, we enumerate $(b,c,d)$ from $Q_{2}^{\prime}$, and for every $(b,c)$, we enumerate the corresponding $A$-values from $Q_{1}$. Note that at least one $A$-value must exist because $Q_{2}^{\prime}$ includes $P_{1}$, which is the projection of $Q_{1}$.
In the above example, we only needed to do the calibration bottom-up from the leaf $Q_{1}$ to the root $Q_{2}^{\prime}$. We show in Appendix B.4 that this is not sufficient for $\text{IVM}^{+}_{\delta}[Q]$. Instead, we also need to do an extra calibration top-down. For example, suppose we have an insert of a tuple $(a,b)$ into $R$ above. In order to enumerate the new output tuples corresponding to $(a,b)$, we have to start our enumeration from $Q_{1}$ as the root. For that purpose, $Q_{1}$ also needs to be calibrated with $Q_{2}^{\prime}$. See Example B.3.
6. Technical Tools: Intersection Joins
Before we explain our upper and lower bounds for IVM in the insert-delete setting in Sec. 7, we review in this section some necessary background on intersection joins in the static setting.
Definition 6.1 (Intersection queries (Khamis et al., 2022)).
An intersection query is a (full conjunctive) query that contains interval variables, which are variables whose domains are discrete intervals over the natural numbers. For instance, the discrete interval $[3,7]$ consists of the natural numbers from $3$ to $7$. For better distinction, we denote an interval variable by $[X]$ and call the variables that are not interval variables point variables. A value of an interval variable $[X]$ is denoted by $[x]$, which is an interval. Two or more intervals join if they intersect. In particular, the semantics of an intersection query $Q=R_{1}(\bm{X}_{1})\wedge\cdots\wedge R_{k}(\bm{X}_{k})$ is defined as follows. A tuple $\bm{t}$ over the schema $\text{\sf vars}(Q)$ is in the result of $Q$ if there are tuples $\bm{t}_{i}\in R_{i}$ for $i\in[k]$ such that for all $X\in\text{\sf vars}(Q)$, it holds:
  • if $X$ is a point variable, then the set of points $\{\bm{t}_{i}(X)\mid i\in[k]\wedge X\in\bm{X}_{i}\}$ consists of a single value, which is the point $\bm{t}(X)$;
  • if $[X]$ is an interval variable, then the intervals in the set $\{\bm{t}_{i}([X])\mid i\in[k]\wedge[X]\in\bm{X}_{i}\}$ have a non-empty intersection, which is the interval $\bm{t}([X])$.
For the purpose of this paper, we are only interested in a special class of intersection queries, defined below:
Definition 6.2 (The time extension $\overline{Q}$ of a query $Q$).
Let $Q=R_{1}(\bm{X}_{1})\wedge\cdots\wedge R_{k}(\bm{X}_{k})$ be a query where $\text{\sf vars}(Q)$ are all point variables, and let $[Z]$ be a new interval variable. The time extension $\overline{Q}$ of $Q$ is an intersection query that results from $Q$ by adding $[Z]$ to the schema of each atom:
(10) $\overline{Q}=\overline{R}_{1}([Z],\bm{X}_{1})\wedge\cdots\wedge\overline{R}_{k}([Z],\bm{X}_{k})$
A query $\overline{Q}$ is called a time extension if it is the time extension of some query $Q$.
We call $\overline{Q}$ the “time extension” of $Q$ because in Sec. 7.1, we use the interval $[Z]$ in $\overline{Q}$ to represent the “lifespan” of tuples in $Q$.
Example 6.3 ().
Consider the following intersection query, which is the time extension of $Q_{\triangle}$ from Eq. (1):
(11) $\overline{Q}_{\triangle}([Z],A,B,C)=\overline{R}([Z],A,B)\wedge\overline{S}([Z],B,C)\wedge\overline{T}([Z],A,C)$
In the above, $[Z]$ is an interval variable while $A,B,C$ are point variables. Consider the database instance $\overline{\mathcal{D}}$ depicted in Figure 1(d). The output of $\overline{Q}_{\triangle}$ on this instance is depicted in Figure 1(d).
Figure 1. A database instance $\overline{\mathcal{D}}$ for the query $\overline{Q}_{\triangle}$ in Eq. (11).
$[Z]$ $A$ $B$
$[1,8]$ $a_{1}$ $b_{1}$
(a) $\overline{R}$
$[Z]$ $B$ $C$
$[2,5]$ $b_{1}$ $c_{1}$
$[4,6]$ $b_{2}$ $c_{1}$
(b) $\overline{S}$
$[Z]$ $A$ $C$
$[3,7]$ $a_{1}$ $c_{1}$
(c) $\overline{T}$
$[Z]$ $A$ $B$ $C$
$[3,5]$ $a_{1}$ $b_{1}$ $c_{1}$
(d) $\overline{Q}_{\triangle}$
A database instance $\overline{\mathcal{D}}$ for the query $\overline{Q}_{\triangle}$ in Eq. (11).
Figure 1. A database instance $\overline{\mathcal{D}}$ for the query $\overline{Q}_{\triangle}$ in Eq. (11).
Let $\overline{Q}$ be the time extension of a query $Q$. Prior work reduces the evaluation of $\overline{Q}$ to the evaluation of $\widehat{Q}$, where $\widehat{Q}$ is the multivariate extension of $Q$ (Eq. (3)), whose variables are all point variables (Khamis et al., 2022). Moreover, it is shown in (Khamis et al., 2022) that this reduction is optimal (up to a $\widetilde{\mathcal{O}}(1)$ factor). In particular, the query $\overline{Q}$ is exactly as hard as the hardest component $\widehat{Q}_{\bm{\sigma}}$ in $\widehat{Q}$. This is shown by a backward reduction from (the static evaluation of) each component $\widehat{Q}_{\bm{\sigma}}$ to $\overline{Q}$. We explain both reductions in detail in Appendix C and briefly summarize the forward reduction below. We rely on both reductions to prove upper and lower bounds for IVM in the insert-delete setting (Sec. 3.2).
Reduction from $\overline{Q}$ to $\widehat{Q}$
Suppose we want to evaluate $\overline{Q}$ on a given database instance $\overline{\mathcal{D}}$. We construct a segment tree, $\mathcal{T}_{N}$, over the interval variable $[Z]$, where the parameter $N$ is roughly the maximum number of different $Z$-values. A segment tree $\mathcal{T}_{N}$ is a balanced binary tree with $N$ leaves. Each node corresponds to an interval. The root corresponds to the interval $[1,N]$, its left and right child correspond to the intervals $[1,N/2]$ and $[N/2+1,N]$ respectively, and so on. Each node can be represented by a bitstring of length at most $\log_{2}N$. Figure 2 in Appendix C depicts the segment tree $\mathcal{T}_{8}$. Each interval $[z]$ can be broken down into at most $\mathcal{O}(\log N)$ nodes in $\mathcal{T}_{N}$. We refer to these nodes as the canonical partition of $[z]$, and denote them by $\text{\sf CP}_{N}([z])$. For example, the canonical partitions of $[2,8]$ and $[2,5]$ in the segment tree $\mathcal{T}_{8}$ are:
(12) $\text{\sf CP}_{8}([2,8])=\{001,01,1\},\quad\quad\text{\sf CP}_{8}([2,5])=\{001,01,100\}$
The canonical partition of the intersection of some intervals is roughly the intersection of their canonical partitions (Lemma C.1). Two (or more) nodes in the segment tree correspond to overlapping intervals if and only if the corresponding bitstrings have a prefix-suffix relationship (Lemma C.1). For example, the nodes labeled by $1$ and $100$ in Figure 2 correspond to two overlapping intervals, namely $[5,8]$ and $[5,5]$, because the string $1$ is a prefix of $100$. Therefore, testing intersections can be reduced to testing whether some bitstrings form a chain of prefixes. Armed with this idea, we convert $\overline{\mathcal{D}}$ into a database instance $\widehat{\mathcal{D}}$ for $\widehat{Q}$, called the canonical partition of $\overline{\mathcal{D}}$ and denoted by $\text{\sf CP}_{N}(\overline{\mathcal{D}})$. There is a mapping between the outputs $\overline{Q}(\overline{\mathcal{D}})$ and $\widehat{Q}(\widehat{\mathcal{D}})$ (Theorem C.5).
Example 6.4 ().
Suppose we want to reduce $\overline{Q}_{\triangle}$ from Eq. (11) to $\widehat{Q}_{\triangle}$ from Example 2.4. In particular, we are given the database instance $\overline{\mathcal{D}}$ from Figure 1(d) and want to compute the corresponding instance $\widehat{\mathcal{D}}$ for $\widehat{Q}_{\triangle}$. We show how to construct $\widehat{R}_{123},\widehat{S}_{123},\widehat{T}_{123}$ for $\widehat{Q}_{123}$ in Eq. (4). The remaining $\widehat{Q}_{132},\ldots,\widehat{Q}_{321}$ are similar. The query $\widehat{Q}_{123}$ is meant to test whether $\overline{R},\overline{S},\overline{T}$ contain three intervals $[z^{R}],[z^{S}],[z^{T}]$ whose canonical partitions contain bitstrings $z^{R},z^{S},z^{T}$ where $z^{R}$ is a prefix of $z^{S}$ which is a prefix of $z^{T}$. This can only happen if there are bitstrings $z_{2},z_{3}$ where $z^{S}=z^{R}\circ z^{2}$ and $z^{T}=z^{S}\circ z^{3}$, where $\circ$ denotes string concatenation. We define:
(13) $\widehat{R}_{123}$ =$\stackrel{{\scriptstyle\text{def}}}{{=}}\{(z_{1},a,b)\mid\exists[z]:([z],a,b)\in\overline{R}\wedge z_{1}\in\text{\sf CP}_{N}([z])\}$
$\widehat{S}_{123}$ =$\stackrel{{\scriptstyle\text{def}}}{{=}}\{(z_{1},z_{2},b,c)\mid\exists[z]:([z],b,c)\in\overline{S}\wedge(z_{1}\circ z_{2})\in\text{\sf CP}_{N}([z])\}$
$\widehat{T}_{123}$ =$\stackrel{{\scriptstyle\text{def}}}{{=}}\{(z_{1},z_{2},z_{3},a,c)\mid\exists z:([z],a,c)\in\overline{T}\wedge$
$\quad\quad\quad\quad(z_{1}\circ z_{2}\circ z_{3})\in\text{\sf CP}_{N}([z])\}$
Let $\varepsilon$ denote the empty string. Applying the above to $\overline{\mathcal{D}}$ from Figure 1(d), $\widehat{R}_{123},\widehat{S}_{123},\widehat{T}_{123}$ respectively contain the tuples: $(\varepsilon,a_{1},b_{1})$, $(\varepsilon,01,b_{1},c_{1})$, $(\varepsilon,01,\varepsilon,a_{1},c_{1})$, which join together producing the output tuple $(\varepsilon,01,\varepsilon,a_{1},b_{1},c_{1})$ of $\widehat{Q}_{123}$. This is a witness to the output tuple $([3,5],a_{1},b_{1},c_{1})$ of $\overline{Q}_{\triangle}$. See Example C.2 for more details.
7. IVM: Insert-Delete Setting
In this section, we give a brief overview of how we obtain our upper and lower bounds for IVM in the insert-delete setting from Theorems 3.4 and 3.6, respectively. Details are deferred to Appendix D.
7.1. Upper bound for $\text{IVM}^{\pm}[Q]$
The following example demonstrates our algorithm for $\text{IVM}^{\pm}[Q]$ from Theorem 3.4. Algorithm 1 in Appendix D.1 shows the general algorithm. Appendix D.2 shows how to extend it to $\text{IVM}^{\pm}_{\delta}[Q]$.
Example 7.1 (for Theorem 3.4).
Suppose we want to solve the problem $\text{IVM}^{\pm}[Q_{\triangle}]$ for $Q_{\triangle}$ in Eq. (1), where we have a stream of $N$ single-tuple inserts/deletes into $R,S$ and $T$. For simplicity, assume that $N$ is known in advance (Remark D.1 explains how to remove this assumption.) We introduce a new interval variable $[Z]$ and use it to represent time. Specifically, $[Z]$ will represent the lifespan of every tuple in the database, in the spirit of temporal databases (Jensen and Snodgrass, 2018). By adding $[Z]$ to every atom, we obtain the time extension $\overline{Q}_{\triangle}$ of $Q_{\triangle}$ in Eq. (11). Suppose that the $\tau$-th update is an insert of a tuple $(a,b)$ into $R$, i.e. $+R(a,b)$. Then, we apply the insert $+\overline{R}([\tau,\infty],a,b)$ into the time extension, indicating that the tuple $(a,b)$ lives in $R$ from time $\tau$ on (since we don’t know yet its future deletion time). Now suppose that the $\tau^{\prime}$-th update (for some $\tau^{\prime}>\tau$) is a delete of the same tuple $(a,b)$ in $R$. Then, we replace the tuple $([\tau,\infty],a,b)$ in $\overline{R}$ with $([\tau,\tau^{\prime}],a,b)$. We can extract the result of $Q_{\triangle}$ at any time $\tau$ by selecting output tuples $([z],a,b,c)$ of $\overline{Q}_{\triangle}$ where the interval $[z]$ contains the current time $\tau$. Therefore, if we can efficiently maintain $\overline{Q}_{\triangle}$, we can efficiently maintain $Q_{\triangle}$. Table 2 shows the same stream of 8 updates into $Q_{\triangle}$ as Table 1 but adds the corresponding updates to $\overline{Q}_{\triangle}$. The final database $\overline{\mathcal{D}}$ after all updates have taken place is the same as the one shown in Figure 1(d).
$\tau$ $\delta_{\tau}\mathcal{D}$ $\delta_{\tau}\overline{\mathcal{D}}$
1 $+R(a_{1},b_{1})$ $+\overline{R}([1,\infty],a_{1},b_{1})$
2 $+S(b_{1},c_{1})$ $+\overline{S}([2,\infty],b_{1},c_{1})$
3 $+T(a_{1},c_{1})$ $+\overline{T}([3,\infty],a_{1},c_{1})$
4 $+S(b_{2},c_{1})$ $+\overline{S}([4,\infty],b_{2},c_{1})$
5 $-S(b_{1},c_{1})$ $-\overline{S}([2,\infty],b_{1},c_{1})$, $+\overline{S}([2,5],b_{1},c_{1})$
6 $-S(b_{2},c_{1})$ $-\overline{S}([4,\infty],b_{2},c_{1})$, $+\overline{S}([4,6],b_{2},c_{1})$
7 $-T(a_{1},c_{1})$ $-\overline{T}([3,\infty],a_{1},c_{1})$, $+\overline{T}([3,7],a_{1},c_{1})$
8 $-R(a_{1},b_{1})$ $-\overline{R}([1,\infty],a_{1},b_{1})$, $+\overline{R}([1,8],a_{1},b_{1})$
Table 2. Updates to $Q_{\triangle}$ (1) (same as Table 1) along with the corresponding updates to $\overline{Q}_{\triangle}$ (11). At the end (i.e. $\tau=8$), the database $\overline{\mathcal{D}}$ is the same as the one shown in Figure 1(d).
To maintain $\overline{Q}_{\triangle}$, we use the six components $\widehat{Q}_{123},\ldots,\widehat{Q}_{321}$ of the multivariate extension $\widehat{Q}_{\triangle}$ of $Q_{\triangle}$ from Example 2.4. In particular, we construct and maintain the relations $\widehat{R}_{123},\widehat{S}_{123}$ and $\widehat{T}_{123}$ from Eq. (13) and forth. We use them along with Theorem 3.1 to maintain $\widehat{Q}_{123}$ (and the same goes for $\widehat{Q}_{132},\ldots$). For example, suppose that the number of updates $N$ is $8$, and consider the insert $+S(b_{1},c_{1})$ at time $2$ in Table 2. This insert corresponds to inserting the following tuples into $\widehat{S}_{123}$: (Recall Eq. (12). Below, $z_{1}$ and $z_{2}$ are bitstrings and $z_{1}\circ z_{2}$ is their concatenation.)
+$+\{(z_{1},z_{2},b_{1},c_{1})\mid(z_{1}\circ z_{2})\in\text{\sf CP}_{8}([2,8])\}$
The corresponding delete $-S(b_{1},c_{1})$ at time 5 corresponds to deleting and inserting the following two sets of tuples from/to $\widehat{S}_{123}$:
$-\{(z_{1},z_{2},b_{1},c_{1})\mid(z_{1}\circ z_{2})\in\text{\sf CP}_{8}([2,8])\}$
+$+\{(z_{1},z_{2},b_{1},c_{1})\mid(z_{1}\circ z_{2})\in\text{\sf CP}_{8}([2,5])\}$
Out of the box, Theorem 3.1 is only limited to inserts. However, we can show that these pairs of inserts/deletes have a special structure that allows us to maintain the guarantees of Theorem 3.1. In particular, at time 5, when we truncate a tuple $\bm{t}$ from $[2,8]$ to $[2,5]$, even after truncation, $\bm{t}$ still joins with the same set of tuples that it used to join with before the truncation. This is because the only intervals that overlap with $[2,8]$ but not with $[2,5]$ are the ones that start after 5. But these intervals are in the future, hence they don’t exist yet in the database (see Proposition D.2). We use this idea to amortize the cost of truncations over the inserts.
For $\text{IVM}^{\pm}[Q_{\triangle}]$, our target is to do constant-delay enumeration of the full output $Q_{\triangle}$ at time $\tau$. To that end, we enumerate tuples $([z],a,b,c)$ of $\overline{Q}_{\triangle}$ where the interval $[z]$ contains the current time $\tau$. And to enumerate those, we enumerate tuples $(z_{1},z_{2},z_{3},a,b,c)$ from $\widehat{Q}_{123},\ldots,\widehat{Q}_{321}$ where $z_{1}\circ z_{2}\circ z_{3}$ corresponds to an interval in $\text{\sf CP}_{N}([z])$ that contains $\tau$. This is a selection condition over $(z_{1},z_{2},z_{3})$. By construction, the tree decomposition of any component, say $\widehat{Q}_{123}$, must contain a bag $W$ that contains the variables $\{z_{1},z_{2},z_{3}\}$. This is because, by Definition 2.3, $\widehat{Q}_{123}$ contains an atom $\widehat{T}_{123}$ that contains these variables. We designate $W$ as the root of tree decomposition of $\widehat{Q}_{123}$ and start our constant-delay enumeration from $W$. The same goes for $\widehat{Q}_{132},\ldots,\widehat{Q}_{321}$. Moreover, the enumeration outputs of $\widehat{Q}_{123},\ldots,\widehat{Q}_{321}$ are disjoint in this case. See the proof of Lemma D.3.
For $\text{IVM}^{\pm}_{\delta}[Q_{\triangle}]$, our target is to enumerate the change to the output at time $\tau$. To that end, we enumerate tuples $([z],a,b,c)$ of $\overline{Q}_{\triangle}$ where the interval $[z]$ has $\tau$ as an endpoint. Appendix D.2 explains the technical challenges and how to address them.
7.2. Lower bound for $\text{IVM}^{\pm}_{\delta}[Q]$
The following example demonstrates our lower bound for $\text{IVM}^{\pm}_{\delta}[Q]$ from Theorem 3.6, which is based on a reduction from the static evaluation of a component $\widehat{Q}_{\bm{\sigma}}$ of the multivariate extension of $Q$ to $\text{IVM}^{\pm}_{\delta}[Q]$. Algorithm 2 in Appendix D.4 details the reduction.
Example 7.2 (for Theorem 3.6).
Consider $Q_{\triangle}$ from Eq. (1) and $\widehat{Q}_{123}$ from Eq. (4). Suppose that there exists $\gamma>0$ such that $\text{IVM}^{\pm}_{\delta}[Q_{\triangle}]$ can be solved with amortized update time $\widetilde{\mathcal{O}}(N^{\alpha})$ where $N$ is the number of single-tuple updates, $\alpha\stackrel{{\scriptstyle\text{def}}}{{=}}\omega(\widehat{Q}_{123})-1-\gamma$, and $\omega$ is given by Definition 2.2. We show that this implies that $\text{Eval}[\widehat{Q}_{123}]$ can be solved in time $\widetilde{\mathcal{O}}(|\widehat{\mathcal{D}}_{123}|^{\omega(\widehat{Q}_{123})-\gamma}+|\widehat{Q}_{123}(\widehat{\mathcal{D}}_{123})|)$ on any database instance $\widehat{\mathcal{D}}_{123}$, thus contradicting the definition of $\omega(\widehat{Q}_{123})$.
Let us take an arbitrary database instance $\widehat{\mathcal{D}}_{123}=(\widehat{R}_{123},\widehat{S}_{123},\widehat{T}_{123})$ and evaluate $\widehat{Q}_{123}$ over $\widehat{\mathcal{D}}_{123}$ using the oracle that solves $\text{IVM}^{\pm}_{\delta}[Q_{\triangle}]$. Example C.6 shows how to reduce $\text{Eval}[\widehat{Q}_{123}]$ over $\widehat{\mathcal{D}}_{123}$ to the static evaluation of $\overline{Q}_{\triangle}$ from Eq. (11) over a database instance $\overline{\mathcal{D}}=(\overline{R},\overline{S},\overline{T})$ that satisfies:
$|\overline{\mathcal{D}}|=|\widehat{\mathcal{D}}_{123}|,\quad\quad|\overline{Q}_{\triangle}(\overline{\mathcal{D}})|=|\widehat{Q}_{123}(\widehat{\mathcal{D}}_{123})|.$
Moreover, $\overline{\mathcal{D}}$ can be constructed in linear time (Theorem C.9.) Next, we reduce the evaluation of $\overline{Q}_{\triangle}$ over $\overline{\mathcal{D}}$ into $\text{IVM}^{\pm}_{\delta}[Q_{\triangle}]$.
We interpret the interval $[Z]$ of a tuple in $\overline{Q}_{\triangle}$ as the “lifespan” of a corresponding tuple in $Q_{\triangle}$, similar to Section 7.1. In particular, we initialize relations $R,S,T$ for $Q_{\triangle}$. We also sort the tuples $([\alpha,\beta],a,b)$ of $\overline{R}$ based on the end points $\alpha$ and $\beta$ of their $[Z]$-intervals. Each tuple $([\alpha,\beta],a,b)$ appears twice in the order: once paired with its beginning $\alpha$ and another with its end $\beta$. Similarly, we add the tuples $([\alpha,\beta],b,c)$ of $\overline{S}$ and $([\alpha,\beta],a,c)$ of $\overline{T}$ to the same sorted list, where each tuple appears twice. Now, we go through the list in order. If the next tuple $([\alpha,\beta],a,b)$ is paired with its beginning $\alpha$, then we insert the tuple $(a,b)$ into $R$. However, if the next tuple $([\alpha,\beta],a,b)$ is paired with its end $\beta$, then we delete $(a,b)$ from $R$. The same goes for $S$ and $T$. The number of updates is $2|\overline{\mathcal{D}}|$, and the total time needed to process them is $\widetilde{\mathcal{O}}(|\overline{\mathcal{D}}|^{\alpha+1})$.
To compute the output of $\overline{Q}_{\triangle}$, we do the following. After every insert or delete in $Q_{\triangle}$, we use the oracle for $\text{IVM}^{\pm}_{\delta}[Q_{\triangle}]$ to enumerate the change in the output of $Q_{\triangle}$. Whenever the oracle reports an insert of an output tuple $(a,b,c)$ at time $\tau$ and the delete of the same tuple at a later time $\tau^{\prime}>\tau$, then we know that $([\tau,\tau^{\prime}],a,b,c)$ is an output tuple of $\overline{Q}_{\triangle}$. Therefore, the overall time needed to compute the output of $\overline{Q}_{\triangle}$ (including the total update time from before) is $\widetilde{\mathcal{O}}(|\overline{\mathcal{D}}|^{\alpha+1}+|\overline{Q}_{\triangle}(\overline{\mathcal{D}})|)=$ $\widetilde{\mathcal{O}}(|\widehat{\mathcal{D}}_{123}|^{\alpha+1}+|\widehat{Q}_{123}(\widehat{\mathcal{D}}_{123})|)$. This contradicts the definition of $\omega(\widehat{Q}_{123})$.
Appendix D.4 highlights the challenges in extending the above lower bound from $\text{IVM}^{\pm}_{\delta}$ to $\text{IVM}^{\pm}$, which remains an open problem.
8. Conclusion
This paper puts forward two-way reductions between the dynamic and static query evaluation problems, allowing the transfer of both algorithms and lower bounds between them. For the dynamic problem, it also characterizes the complexity gap between the insert-only and the insert-delete settings. The proposed algorithms recover best known results and produce new ones. Unlike existing lower bounds, the lower bounds obtained in the paper are more general as they do not rely on existing complexity conjectures.
The assumptions that the queries have all variables free and the input database be initially empty can be lifted without difficulty. To support arbitrary conjunctive queries, our algorithms need to consider free-connex tree decompositions, which ensure constant-delay enumeration of the tuples in the query result. To support non-empty initial databases, we can use a preprocessing phase to construct our data structure, where all initial tuples get the same starting timestamp. The complexities stated in the paper need to be changed to account for the size of the initial database as part of the size of the update stream.
Acknowledgements.
This work was partially supported by NSF-BSF 2109922, NSF-IIS 2314527, NSF-SHF 2312195, and UZH Global Strategy and Partnerships Funding Scheme, and was conducted while some of the authors participated in the Simons Program on Logic and Algorithms in Databases and AI.
Appendix A Missing Details in Preliminaries
Query Classes
Acyclic queries. A query is ($\alpha$-)acyclic if we can construct a tree, called join tree, such that the nodes of the tree are the atoms of the query (coverage) and for each variable, it holds: if the variable appears in two atoms, then it appears in all atoms on the path connecting the two atoms (connectivity(Brault-Baron, 2016).
We give below some examples of acyclic, hierarchical, and Loomis-Whitney queries, where the latter two are defined in Section 4. Recall that for a query $Q$ and a variable $X\in\text{\sf vars}(Q)$, we use $\text{\sf at}(X)$ to denote the set of atoms of $Q$ that contain $X$ in their schema.
Example A.1 (Query Classes).
The query $Q_{\text{h}}=R(A,B)\wedge S(A,C)$ is hierarchical. We have $\text{\sf at}(A)=\{R(A,B),S(A,C)\}$, $\text{\sf at}(B)=\{R(A,B)\}$, and $\text{\sf at}(C)=\{S(A,C)\}$. It holds $\text{\sf at}(B)\subseteq\text{\sf at}(A)$, $\text{\sf at}(C)\subseteq\text{\sf at}(A)$, and $\text{\sf at}(B)\cap\text{\sf at}(C)=\emptyset$. The query $Q_{\text{nh}}=R(A)\wedge S(A,B)\wedge T(B)$ is acyclic but not hierarchical. Indeed, we have $\text{\sf at}(A)=\{R(A),S(A,B)\}$ and $\text{\sf at}(B)=\{T(B),$ $S(A,B)\}$ and it does not hold $\text{\sf at}(A)\subseteq\text{\sf at}(B)$ nor $\text{\sf at}(B)\subseteq\text{\sf at}(A)$ nor $\text{\sf at}(A)\cap\text{\sf at}(B)=\emptyset$. A possible join tree for $Q_{nh}$ has $R(A,B)$ as root and the other two atoms as the children of the root. The triangle query (or Loomis-Whitney query of degree $3$) is of the form $Q_{\triangle}=R(A,B)\wedge S(B,C)\wedge T(A,C)$. A Loomis-Whitney query of degree $4$ is of the form $Q_{\text{lw4}}=R(A,B,C)\wedge S(B,C,D)\wedge T(C,D,A)\wedge U(D,A,B)$.
Width Measures
Consider the following linear program for a query $Q$ and a database $\mathcal{D}$:
min $\makebox[10.76385pt][r]{}\sum_{R(\bm{X})\in\,\text{\sf at}(Q)}\log(|R|)\cdot\lambda_{R(\bm{X})}$
[…]
Every feasible solution $(\lambda_{R(\bm{X})})_{R(\bm{X})\in\text{\sf at}(Q)}$ to the above program is called a fractional edge cover of $Q$ with respect to $\mathcal{D}$. The optimal objective value of the program is called the fractional edge cover number $\rho^{*}(Q,\mathcal{D})$ of $Q$ with respect to $\mathcal{D}$. The fractional edge cover number gives a bound on the size of $Q(\mathcal{D})$: $|Q(\mathcal{D})|\leq 2^{\rho^{*}(Q,\mathcal{D})}$ (Atserias et al., 2013). This bound is known as the AGM-bound and denoted as $\text{\sf AGM}(Q,D)$. We denote by $\rho^{\ast}(Q)$ the optimal objective value of the linear program that results from the above program by replacing the objective function with $\sum_{R(\bm{X})\in\,\text{\sf at}(Q)}\lambda_{R(\bm{X})}$.
Definition A.2 (Tree Decomposition).
A tree decomposition of a query $Q$ is a pair $(\mathcal{T},\chi)$, where $\mathcal{T}$ is a tree with vertices $V(\mathcal{T})$ and $\chi:V(\mathcal{T})\rightarrow 2^{\text{\sf vars}(Q)}$ maps each node $t$ of the tree $\mathcal{T}$ to a subset $\chi(t)$ of variables of $Q$ such that the following properties hold:
  1. (1)
    for every atom $R(\bm{X})\in\text{\sf at}(Q)$, the schema $\bm{X}$ is a subset of $\chi(t)$ for some $t\in V(\mathcal{T})$, and
  2. (2)
    for every variable $X\in\text{\sf vars}(Q)$, the set $\{t\mid X\in\chi(t)\}$ is a non-empty connected subtree of $\mathcal{T}$. The sets $\chi(t)$ are called the bags of the tree decomposition.
We use $\text{\sf TD}(Q)$ to denote the set of tree decompositions of a query $Q$.
Definition A.3 (The restriction $Q_{\bm{Y}}$ of a query $Q$).
For a query $Q$ and a subset $\bm{Y}\subseteq\text{\sf vars}(Q)$, we define the restriction of $Q$ to $\bm{Y}$, denoted by $Q_{\bm{Y}}$, to be the query that results from $Q$ by restricting the schema of each atom to those variables that appear in $\bm{Y}$. Formally:
$R_{\bm{Y}}(\bm{X}\cap\bm{Y})$ =$\stackrel{{\scriptstyle\text{def}}}{{=}}\pi_{\bm{X}\cap\bm{Y}}R(\bm{X}),\quad\quad\text{for all $R(\bm{X})\in\text{\sf at}(Q)$}$
(14) $Q_{\bm{Y}}(\bm{Y})$ =$\stackrel{{\scriptstyle\text{def}}}{{=}}\bigwedge_{R(\bm{X})\in\text{\sf at}(Q)}R_{\bm{Y}}(\bm{X}\cap\bm{Y})$
Definition A.4 (Fractional Hypertree Width).
Given a query $Q$ and a tree decomposition $(\mathcal{T},\chi)$ of $Q$, the fractional hypertree width of $(\mathcal{T},\chi)$ and of $Q$ are defined respectively as follows:
$\mathsf{w}(\mathcal{T},\chi)$ =$\stackrel{{\scriptstyle\text{def}}}{{=}}\max_{t\in V(\mathcal{T})}\rho^{*}(Q_{\chi(t)})$
(15) $\mathsf{w}(Q)$ =$\stackrel{{\scriptstyle\text{def}}}{{=}}\min_{(\mathcal{T},\chi)\in\text{\sf TD}(Q)}\mathsf{w}(\mathcal{T},\chi)$
We call a tree decomposition $(\mathcal{T},\chi)$ of $Q$ optimal if $\mathsf{w}(\mathcal{T},\chi)=\mathsf{w}(Q)$. The fractional hypertree width of a union $Q=\bigcup_{i\in[n]}Q_{i}$ of join queries is
$\mathsf{w}(Q)$ =$\stackrel{{\scriptstyle\text{def}}}{{=}}\max_{i\in[n]}\mathsf{w}(Q_{i})$
Variable Elimination Orders
We review the equivalence between tree decompositions and variable elimination orders, which is a folklore concept in graph and database theory; see e.g. (Abo Khamis et al., 2016) and references thereof. Given a query $Q$, a variable elimination order $\bm{\mu}$ for $Q$ is an order of the variables of $Q$ that we use to eliminate them one by one. In order to eliminate a variable $X$, we take all atoms $R(\bm{X})$ where $X\in\bm{X}$ and replace them with a new atom $R_{X}^{\bm{\mu}}(\bm{U}_{X}^{\bm{\mu}}-\{X\})$ where $\bm{U}_{X}^{\bm{\mu}}$ is the union of schemas $\bm{X}$ of all these atoms $R(\bm{X})$ that contain $X$. It is long known that for a fixed variable elimination order $\bm{\mu}$, the sets $\bm{U}_{X}^{\bm{\mu}}$ are the bags of some tree decomposition of $Q$. Moreover, one of these tree decompositions that are obtained from variable elimination orders is optimal, i.e. has a minimum fractional hypertree width. Therefore, the fractional hypertree width can be equivalently defined as follows. Let $\Sigma_{\text{\sf vars}(Q)}$ denote the set of all permutations of the variables of $Q$, and let $\bm{\mu}\in\Sigma_{\text{\sf vars}(Q)}$ be a variable elimination order for $Q$:
$\mathsf{w}(\bm{\mu})$ =$\stackrel{{\scriptstyle\text{def}}}{{=}}\max_{X\in\bm{\mu}}\rho^{*}(Q_{\bm{U}_{X}^{\bm{\mu}}})$
(16) $\mathsf{w}(Q)$ =$\stackrel{{\scriptstyle\text{def}}}{{=}}\min_{\bm{\mu}\in\Sigma_{\text{\sf vars}(Q)}}\mathsf{w}(\bm{\mu})$
In particular, it is known that (16) and (15) are equivalent; see for example (Abo Khamis et al., 2016) and references thereof. We call $\mathsf{w}(\bm{\mu})$ the fractional hypertree width of the variable elimination order $\bm{\mu}$. Moreover, we call $\bm{\mu}$ optimal if $\mathsf{w}(\bm{\mu})=\mathsf{w}(Q)$.
Online Matrix-Vector Multiplication (OMv) Conjecture
In order to state the OMv conjecture, we first state the OMv problem. Given a Boolean matrix $\bm{M}$ and $n$ column vectors $\bm{v}_{1},\ldots,\bm{v}_{n}$ that arrive one-by-one, the task is to output the vector $\bm{M}\bm{v}_{r}$ after seeing $\bm{v}_{r}$ for each $r\in[n]$:
Problem: OMv Given: $n\times n$ Boolean matrix $\bm{M}$ and a stream of $n$ $n\times 1$ Boolean column vectors $\bm{v}_{1},\ldots,\bm{v}_{n}$ Task: Compute $\bm{M}\bm{v}_{r}$ after seeing each $\bm{v}_{r}$
Conjecture A.5 (OMv (Henzinger et al., 2015)).
For any $\gamma>0$, there is no algorithm that solves the OMv problem in $\mathcal{O}(n^{3-\gamma})$ time.
Online Vector-Matrix-Vector Multiplication (OuMv) Conjecture
The OuMv problem is defined as follows. Given a Boolean matrix $\bm{M}$ and $n$ pairs of column vectors $(\bm{u}_{1},\bm{v}_{1}),\ldots,(\bm{u}_{n},\bm{v}_{n})$ that arrive one-by-one, the task is to output the product ${\bm{u}}^{\top}_{r}\bm{M}\bm{v}_{r}$ after seeing $(\bm{u}_{r},\bm{v}_{r})$ for each $r\in[n]$:
Problem: OuMv Given: $n\times n$ Boolean matrix $\bm{M}$ and a stream of $n$ pairs of $n\times 1$ Boolean column vectors $(\bm{u}_{1},\bm{v}_{1}),\ldots,(\bm{u}_{n},\bm{v}_{n})$ Task: Compute $\bm{u}_{r}^{\top}\bm{M}\bm{v}_{r}$ after seeing each $(\bm{u}_{r},\bm{v}_{r})$
Conjecture A.6 (OuMv (Henzinger et al., 2015)).
For any $\gamma>0$, there is no algorithm that solves the OuMv problem in $\mathcal{O}(n^{3-\gamma})$ time.
It is known that the OMv conjecture implies the OuMv conjecture (Henzinger et al., 2015).
Appendix B Missing Details in Section 5
In this section, we explain in detail our results on IVM in the insert-only setting, where no deletes are allowed. The main goal is to prove Theorem 3.1 (as well as Theorem 3.3). The results in this section also serve as building blocks for our results on the insert-delete setting in Section D.
B.1. Preliminaries: The Query Decomposition Lemma
We prove the query decomposition lemma (Lemma 17) and illustrate it by an example. This lemma is the core of the proof of Lemma 5.1, which in turn is a key ingredient in the proof of Theorem 3.1. Recall notation from Section A.
The query decomposition lemma basically says the following. Let $Q$ be a query and let $\bm{Y}\subseteq\text{\sf vars}(Q)$. Then, the AGM bound of $Q$ can be decomposed into a sum of AGM bounds of “residual” queries: one query $Q\Join\bm{y}$ for each tuple $\bm{y}$ over the variables $\bm{Y}$.
Lemma B.1 (Query Decomposition Lemma (Ngo et al., 2014; Ngo, 2018)).
Given a query $Q$ and a subset $\bm{Y}\subseteq\text{\sf vars}(Q)$, let $\left(\lambda_{R(\bm{X})}\right)_{R(\bm{X})\in\text{\sf at}(Q)}$ be a fractional edge cover of $Q$. Then, the following inequality holds:
(17) $\sum_{\bm{y}\in\text{\sf Dom}(\bm{Y})}\underbrace{\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R(\bm{X})\ltimes\bm{y}|^{\lambda_{R(\bm{X})}}}_{\text{AGM-bound of $Q\Join\bm{y}$}}\leq\underbrace{\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R|^{\lambda_{R(\bm{X})}}}_{\text{AGM-bound of $Q$}}$
In the above, $\bm{y}\in\text{\sf Dom}(\bm{Y})$ indicates that the tuple $\bm{y}$ has schema $\bm{Y}$. Moreover, $R(\bm{X})\ltimes\bm{y}$ denotes the semijoin of the atom $R(\bm{X})$ with the tuple $\bm{y}$.
Proof of Lemma 17.
In order to prove inequality (17), we prove the following special case where $\bm{Y}$ consists of a single variable $Y$:
(18) $\sum_{y\in\text{\sf Dom}(Y)}\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R(\bm{X})\ltimes y|^{\lambda_{R(\bm{X})}}\leq\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R|^{\lambda_{R(\bm{X})}}$
Inequality (17) follows by repeatedly applying (18) to each variable in $\bm{Y}$. To that end, note that for each “residual” query, $Q_{y}\stackrel{{\scriptstyle\text{def}}}{{=}}Q\Join y$, the following is a valid fractional edge cover, where $R_{y}(\bm{X})\stackrel{{\scriptstyle\text{def}}}{{=}}R(\bm{X})\ltimes y$:
$\left(\lambda_{R_{y}(\bm{X})}\right)_{R_{y}(\bm{X})\in\text{\sf at}(Q_{y})}$
Now, we prove (18):
$\sum_{y\in\text{\sf Dom}(Y)}\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R(\bm{X})\ltimes y|^{\lambda_{R(\bm{X})}}$ $=$
(23) $\sum_{y\in\text{\sf Dom}(Y)}\prod_{\begin{subarray}{c}R(\bm{X})\in\text{\sf at}(Q)\\ Y\in\bm{X}\end{subarray}}|\sigma_{Y=y}R(\bm{X})|^{\lambda_{R(\bm{X})}}\prod_{\begin{subarray}{c}R(\bm{X})\in\text{\sf at}(Q)\\ Y\notin\bm{X}\end{subarray}}|R|^{\lambda_{R(\bm{X})}}$ $=$
(28) $\prod_{\begin{subarray}{c}R(\bm{X})\in\text{\sf at}(Q)\\ Y\notin\bm{X}\end{subarray}}|R|^{\lambda_{R(\bm{X})}}\sum_{y\in\text{\sf Dom}(Y)}\prod_{\begin{subarray}{c}R(\bm{X})\in\text{\sf at}(Q)\\ Y\in\bm{X}\end{subarray}}|\sigma_{Y=y}R(\bm{X})|^{\lambda_{R(\bm{X})}}$$\leq$
(33) $\prod_{\begin{subarray}{c}R(\bm{X})\in\text{\sf at}(Q)\\ Y\notin\bm{X}\end{subarray}}|R|^{\lambda_{R(\bm{X})}}\prod_{\begin{subarray}{c}R(\bm{X})\in\text{\sf at}(Q)\\ Y\in\bm{X}\end{subarray}}\left(\sum_{y\in\text{\sf Dom}(Y)}|\sigma_{Y=y}R(\bm{X})|\right)^{\lambda_{R(\bm{X})}}$ $=$
(38) $\prod_{\begin{subarray}{c}R(\bm{X})\in\text{\sf at}(Q)\\ Y\notin\bm{X}\end{subarray}}|R|^{\lambda_{R(\bm{X})}}\prod_{\begin{subarray}{c}R(\bm{X})\in\text{\sf at}(Q)\\ Y\in\bm{X}\end{subarray}}|R|^{\lambda_{R(\bm{X})}}$ $=$
$\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R|^{\lambda_{R(\bm{X})}}$
Inequality (28) above follows from Hölder’s inequality, which applies specifically because the sum of $\lambda_{R(\bm{X})}$ over all atoms $R(\bm{X})$ that contain $Y$ is at least 1. ∎
Example B.2 (for Lemma 17).
Consider the triangle query:
$Q_{\triangle}(A,B,C)=R(A,B)\wedge S(B,C)\wedge T(A,C)$
By choosing $\lambda_{R(A,B)}=\lambda_{S(B,C)}=\lambda_{T(A,C)}=\frac{1}{2}$ and $\bm{Y}=\{A,B\}$, inequality (17) implies the following:
(39) $\sum_{(a,b)\in R}\sqrt{|\sigma_{B=b}S(B,C)|\cdot|\sigma_{A=a}T(A,C)|}\leq\sqrt{|R|\cdot|S|\cdot|T|}$
This inequality can be proved as follows:
$\sqrt{|R|\cdot|S|\cdot|T|}$ $=$
$\sqrt{|S|}\cdot\sqrt{\left(\sum_{a}|\sigma_{A=a}R(A,B)|\right)\left(\sum_{a}|\sigma_{A=a}T(A,C)|\right)}$$\geq$
$\sqrt{|S|}\cdot\sum_{a}\sqrt{|\sigma_{A=a}R(A,B)|\cdot|\sigma_{A=a}T(A,C)|}$ $=$
$\sum_{a}\sqrt{|\sigma_{A=a}R(A,B)|\cdot|\sigma_{A=a}T(A,C)|\cdot|S|}$ $=$
$\sum_{a}\sqrt{|\sigma_{A=a}T(A,C)|}\sqrt{\left(\sum_{b}|\sigma_{A=a,B=b}R(A,B)|\right)\left(\sum_{b}{|\sigma_{B=b}S(B,C)|}\right)}$$\geq$
$\sum_{a}\sqrt{|\sigma_{A=a}T(A,C)|}\sum_{b}\sqrt{|\sigma_{A=a,B=b}R(A,B)|\cdot{|\sigma_{B=b}S(B,C)|}}$ $=$
$\sum_{a,b}\sqrt{|\sigma_{A=a,B=b}R(A,B)|\cdot|\sigma_{B=b}S(B,C)|\cdot|\sigma_{A=a}T(A,C)|}$ $=$
$\sum_{(a,b)\in R}\sqrt{|\sigma_{B=b}S(B,C)|\cdot|\sigma_{A=a}T(A,C)|}$
The two inequalities above follow from the Cauchy-Schwarz inequality, which is a special case of Hölder’s inequality.
B.2. Proof of Lemma 5.1
Theorem 3.1 gives an upper bound on the amortized update time for IVM in the insert-only setting in terms of the fractional hypertree width of the query $Q$. Before we show how to achieve this bound in Sections B.3 and B.4, we prove in this section a weaker bound that is based on the AGM-bound of $Q$. In particular, we prove Lemma 5.1, which says the following. Given a query $Q$ (whose input relations are initially empty) and a stream of $N$ single-tuple inserts into $Q$, we can update $Q$ after every insert and report the new output tuples $\delta Q$ in a total time for all updates/reports together that does not exceed the AGM bound of $Q$. Note that here we are referring to the final AGM bound of $Q$ after all inserts have been performed.
Lemma 5.1.
Given a query $Q$, an initially empty database $\mathcal{D}^{(0)}$, and a stream of $N$ single-tuple inserts, we can compute the new output tuples, $\delta_{\tau}Q(D)$, after every insert $\delta_{\tau}\mathcal{D}$ where the total time over all inserts combined is $\mathcal{O}(N+\text{\sf AGM}(Q,\mathcal{D}^{(N)}))$.
The proof of Lemma 5.1 relies on the query decomposition lemma (Lemma 17 (Ngo et al., 2014; Ngo, 2018)). Example 5.2 demonstrates Lemma 5.1 over the triangle query $Q_{\triangle}$. Lemma 5.1 considers only the $\text{IVM}^{+}_{\delta}[Q]$ problem, but the proof can be straightforwardly extended to $\text{IVM}^{+}[Q]$.
Proof of Lemma 5.1.
For a given relation $R$ in $Q$, we use $R^{(0)}\stackrel{{\scriptstyle\text{def}}}{{=}}\emptyset,$ $R^{(1)},$ $\ldots$,… , $R^{(N)}$ to denote the sequence of versions of $R$ after each insert. Because we only have inserts here, we have $0=|R^{(0)}|\leq|R^{(1)}|\leq\ldots\leq|R^{(N)}|\leq N$. Moreover, for each $\tau\in[N]$, let $\delta_{\tau}R\stackrel{{\scriptstyle\text{def}}}{{=}}R^{(\tau)}\setminus R^{(\tau-1)}$. Note that $|\delta_{\tau}R|$ is either 1 or 0 depending on whether the $\tau$-th insert was into the relation $R$ or into another relation of $Q$. Let $\left(\lambda_{R(\bm{X})}\right)_{R(\bm{X})\in\text{\sf at}(Q)}$ be a fractional edge cover for $Q$. We will show how to perform all $N$ inserts and report $\delta_{\tau}Q$ after every insert in a total time of $\mathcal{O}(\theta)$ (over all inserts/reports combined) where
$\theta\stackrel{{\scriptstyle\text{def}}}{{=}}\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R^{(N)}|^{\lambda_{R(\bm{X})}}$
In particular, for a fixed atom $S(\bm{Y})\in\text{\sf at}(Q)$, we will show how to perform all inserts into $S$ in a total time of $\mathcal{O}(\theta)$. For any $\tau\in[N]$ where $\delta_{\tau}S\neq\emptyset$, the insertion of $\delta_{\tau}S$ into $S$ can be performed in time:
(40) $\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R^{(\tau)}(\bm{X})\ltimes\delta_{\tau}S(\bm{Y})|^{\lambda_{R(\bm{X})}}$
This can be done by solving the query
$\delta_{\tau}Q\stackrel{{\scriptstyle\text{def}}}{{=}}\bigwedge_{R(\bm{X})\in\text{\sf at}(Q)}R^{(\tau)}(\bm{X})\ltimes\delta_{\tau}S(\bm{Y})$
In particular, (40) is the AGM bound of $\delta_{\tau}Q$. We can meet this bound if we solve $\delta_{\tau}Q$ using a worst-case optimal join algorithm (Ngo et al., 2018; Veldhuizen, 2014; Ngo, 2018). For that purpose, each relation $R$ needs to be indexed with a trie (Veldhuizen, 2014) using an appropriate variable ordering for this query. [^3] We might need to maintain multiple tries (with different variable orderings) for the same relation $R(\bm{X})$ in order to handle inserts into different atoms $S(\bm{Y})$. The total runtime of all inserts into $S$ is therefore bounded by:  [^4] In order to avoid incurring an extra factor of $\log N$ in the runtime, we can switch from tries to hashtables (Veldhuizen, 2014).
$\sum_{\tau\in[N]}\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R^{(\tau)}(\bm{X})\ltimes\delta_{\tau}S(\bm{Y})|^{\lambda_{R(\bm{X})}}$$\leq$
$\sum_{\tau\in[N]}\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R^{(N)}(\bm{X})\ltimes\delta_{\tau}S(\bm{Y})|^{\lambda_{R(\bm{X})}}$$\leq$
$\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R^{(N)}|^{\lambda_{R(\bm{X})}}$ $=\theta$
The first inequality above holds because $R^{(\tau)}\subseteq R^{(N)}$, while the second inequality follows from (17). Moreover, note that (40) is an upper bound on the output size of $\delta_{\tau}Q$ (because it is the AGM-bound). From the above, we have $\sum_{\tau\in[N]}|\delta_{\tau}Q|\leq\theta$. ∎
B.3. Upper bound for $\text{IVM}^{+}[Q]$
In this section, we prove Theorem 3.1 for $\text{IVM}^{+}[Q]$, which basically says the following: Given a query $Q$ whose fractional hypertree width is ${\mathsf{w}(Q)}$ and a stream of $N$ single-tuple inserts (where all input relations are initially empty), we can maintain some data structure for $Q$ and update it after every insert where the total update time over all inserts combined is $\mathcal{O}(N^{\mathsf{w}(Q)})$. Moreover, we can use this data structure at any time to do constant-delay enumeration of the output of $Q$. In particular, fix an optimal tree decomposition of $Q$, i.e. one whose width is ${\mathsf{w}(Q)}$. The data structure that we use consists of a collection of intermediate relations: One for each bag of the chosen tree decomposition. Moreover, every bag will be “calibrated” with its children bags in the tree decomposition. We use Lemma 5.1 to update every bag within a total time of $\mathcal{O}(N^{\mathsf{w}(Q)})$. Moreover, because every bag is calibrated with its children bags, we can always do constant-delay enumeration starting from the root bag. Example 5.3 demonstrates the idea.
Theorem 3.1.
[…]
Proof of Theorem 3.1.
The case of $\text{IVM}^{+}[Q]$: Let $(\mathcal{T},\chi)$ be a tree decomposition of $Q$ whose fractional hypertree width is ${\mathsf{w}(Q)}$. (Recall notation from Section A.) Fix an orientation of $\mathcal{T}$ starting from some root node $r\in V(\mathcal{T})$. Given a node $t\in V(\mathcal{T})$, we will use $\text{\sf parent}(t),\text{\sf children}(t),\text{\sf ancestors}(t)$ and $\text{\sf decendants}(t)$ to refer to the parent, children, ancestors and descendants of $t$ in $\mathcal{T}$, respectively, under the fixed orientation of $\mathcal{T}$. For every node $t\in V(\mathcal{T})$, let $\bm{X}_{t}\stackrel{{\scriptstyle\text{def}}}{{=}}\chi(t)$ be the variables of $t$ and $Q_{t}\stackrel{{\scriptstyle\text{def}}}{{=}}Q_{\bm{X}_{t}}$ be the restriction of $Q$ to $\bm{X}_{t}$, given by Definition A.3. We call $Q_{t}$ a bag query. Let $\bm{X}\stackrel{{\scriptstyle\text{def}}}{{=}}\text{\sf vars}(Q)$. By definition, the query $Q$ is equivalent to the join of the outputs of all bag queries:
$Q(\bm{X})=\bigwedge_{t\in V(\mathcal{T})}Q_{t}(\bm{X}_{t})$
Moreover, for every node $t\in V(\mathcal{T})$, we have $\rho^{*}(Q_{t})\leq{\mathsf{w}(Q)}$. For each node $t$, let $\bm{Y}_{t}\stackrel{{\scriptstyle\text{def}}}{{=}}\bm{X}_{t}\cap\bm{X}_{p}$ where $p$ is the parent of $t$. (If $t$ is the root, then $\bm{X}_{p}$ is considered to be empty, hence $\bm{Y}_{t}$ is empty.) We construct a query plan for $Q$ by defining a collection of intermediate relations $Q_{t}^{\prime},P_{t}^{\prime}$ for every node $t\in V(\mathcal{T})$. In particular, $Q_{t}^{\prime},P_{t}^{\prime}$ are defined inductively starting from the leaves of $\mathcal{T}$ all the way to the root $r$:
(41) $Q^{\prime}_{t}(\bm{X}_{t})$ =$\stackrel{{\scriptstyle\text{def}}}{{=}}Q_{t}(\bm{X}_{t})\wedge\bigwedge_{c\in\text{\sf children}(t)}P_{c}^{\prime}(\bm{Y}_{c})$
(42) $P_{t}^{\prime}(\bm{Y}_{t})$ =$\stackrel{{\scriptstyle\text{def}}}{{=}}\pi_{\bm{Y}_{t}}Q^{\prime}_{t}(\bm{X}_{t})$
For any node $t\in V(\mathcal{T})$, we have:
(43) $Q^{\prime}_{t}(\bm{X}_{t})=\pi_{\bm{X}_{t}}\bigwedge_{d\in\text{\sf decendants}(t)\cup\{t\}}Q_{d}(\bm{X}_{d})$
The proof is by induction starting from the leaves of $\mathcal{T}$ to the root. It is similar to the bottom-up phase of the Yannakakis algorithm (Yannakakis, 1981).
The AGM-bound of each bag query $Q_{t}$ is at most $N^{{\mathsf{w}(Q)}}$. Since adding a new relation to a bag does not increase the AGM bound of the bag, the AGM-bound of each $Q^{\prime}_{t}$ is also at most $N^{{\mathsf{w}(Q)}}$. Hence, the output size of each $Q^{\prime}_{t}$( and by extension $P^{\prime}_{t}$) is at most $N^{\mathsf{w}(Q)}$. Each query $Q^{\prime}_{t}$ can have two types of inserts: Inserts coming from (the input relations of) $Q_{t}$ and inserts coming from $P^{\prime}_{c}$ for the children $c$ of $t$. By Lemma 5.1, all updates to $Q^{\prime}_{t}$ combined take time $\mathcal{O}(N^{\mathsf{w}(Q)})$. Moreover, the projection $P^{\prime}_{t}$ of $Q^{\prime}_{t}$ can also be updated in the same time.
In order to do constant-delay enumeration of the output of $Q(\bm{X})$, we start from the root bag $r$ and enumerate the output of $Q^{\prime}_{r}$. By (43), $Q^{\prime}_{r}(\bm{X}_{r})=\pi_{\bm{X}_{r}}Q(\bm{X})$. For every tuple $\bm{x}_{r}$ in the output of $Q^{\prime}_{r}$, we enumerate the corresponding tuples $\bm{x}_{c}$ in the output of $Q^{\prime}_{c}$ for every child $c$ of $r$. (Note that for every child $c$, there must be at least one tuple $\bm{x}_{c}$ that joins with $\bm{x}_{r}$, thanks to (43).) We inductively continue this process until we reach the leaves of $\mathcal{T}$. In particular, for every node $t\in V(\mathcal{T})$ and every tuple $\bm{x}_{t}$ that we enumerate from $Q^{\prime}_{t}$, we enumerate the corresponding tuples $\bm{x}_{c}$ from $Q^{\prime}_{c}$ for every child $c$ of $t$, and we rely on (43) to ensure that there is at least one tuple $\bm{x}_{c}$ for every child $c$ of $t$. ∎
B.4. Upper bound for $\text{IVM}^{+}_{\delta}[Q]$
In this section, we prove Theorem 3.1 for the $\text{IVM}^{+}_{\delta}[Q]$ problem. In particular, we explain how to modify the algorithm and proof from Section B.3, which was originally meant for the $\text{IVM}^{+}[Q]$ problem. Just like we did in Section B.3, here also we will construct an optimal tree decomposition of $Q$ that minimizes the fractional hypertree width, and we will maintain materialized relations for every bag of the tree decomposition. However, suppose that we have an insert of a tuple $\bm{t}$ into a relation $R$ and we want to enumerate the new output tuples $\delta Q$ that are added due to this particular insert. We can no longer start our constant-delay enumeration from the root of the tree decomposition, as we did in Section B.3. Instead, we have to find the bag $B$ that contains $R$, select $\bm{t}$, and start our enumeration from $B$, as if $B$ were the root of the tree decomposition. In order to be able to enumerate the output of $Q$ starting from any bag $B$ of the tree decomposition, all bags need to be fully calibrated with each other. In particular, in Section B.3, we only calibrated bags in a bottom-up fashion where each bag was calibrated with its children bags. But here, we have to do full calibration consisting of two phases: One bottom-up phase where we calibrate each bag with its children, followed by a top-down phase where we calibrate each bag with its parent. These two calibrations are similar to the two phases of the Yannakakis algorithm (Yannakakis, 1981).
Theorem 3.1.
[…]
Proof of Theorem 3.1.
The case of $\text{IVM}^{+}_{\delta}[Q]$: We will use the same notation as in the proof of Theorem 3.1 for $\text{IVM}^{+}[Q]$ in Section B.3. In particular, here also we will fix an optimal tree decomposition $(\mathcal{T},\chi)$ of $Q$ and we will fix an orientation of $\mathcal{T}$ starting from some root node $r\in V(\mathcal{T})$. In addition to $Q^{\prime}_{t}$ and $P^{\prime}_{t}$ given by (41) and (42), we will maintain two additional materialized relations $Q_{t}^{\prime\prime}$ and $P_{t}^{\prime\prime}$ for every node $t\in V(\mathcal{T})$. However, this time, we will define them inductively starting from the root $r$ all the way to the leaves of $\mathcal{T}$: (Recall that for the root $r$, we have $\bm{Y}_{r}=\emptyset$.)
$Q^{\prime\prime}_{t}(\bm{X}_{t})$ =$\stackrel{{\scriptstyle\text{def}}}{{=}}Q^{\prime}_{t}(\bm{X}_{t})\wedge P^{\prime\prime}_{t}(\bm{Y}_{t})$
$P^{\prime\prime}_{t}(\bm{Y}_{t})$ =$\stackrel{{\scriptstyle\text{def}}}{{=}}\begin{cases}\text{\sf true}&\text{if $t=r$}\\ \pi_{\bm{Y}_{t}}Q^{\prime\prime}_{p}(\bm{X}_{p}),\text{\quad where $p=\text{\sf parent}(t)$}&\text{otherwise}\end{cases}$
By induction, we can now prove that for any node $t$,
(44) $Q^{\prime\prime}_{t}(\bm{X}_{t})=\pi_{\bm{X}_{t}}Q(\bm{X})$
(Compare the above to (43).) The induction now starts from the root $r$ and goes all the way to the leaves of $\mathcal{T}$. It is similar to the top-down phase of the Yannakakis algorithm (Yannakakis, 1981). The AGM bound for each $Q^{\prime\prime}_{t}$ is still upper bounded by $N^{\mathsf{w}(Q)}$. Hence, they can be maintained in a total time of $\mathcal{O}(N^{\mathsf{w}(Q)})$. The same goes for $P^{\prime\prime}_{t}$.
Now suppose we have an insert of a tuple $\bm{t}$ into a relation $R(\bm{Z})$ and we want to enumerate the new output tuples $\delta Q$ that are added due to this insert. We find the node $t$ in $\mathcal{T}$ where $\bm{Z}\subseteq\chi(t)$. The existence of such a node $t$ is guaranteed by the definition of a tree decomposition (Definition A.2). Now, we re-orient the tree $\mathcal{T}$ so that $t$ becomes the root. We enumerate output tuples $\bm{x}_{t}$ of $Q^{\prime\prime}_{t}$ that join with $\bm{t}$. For each such tuple $\bm{x}_{t}$, we enumerate the corresponding tuples $\bm{x}_{c}$ from $Q^{\prime\prime}_{c}$ for every child $c$ of $t$. (These are the children of $t$ under the new orientation of the tree $\mathcal{T}$). By (44), we guarantee that for each tuple $\bm{x}_{t}$ and each child $c$, there must be at least one tuple $x_{c}$ in $Q^{\prime\prime}_{c}$ that joins with $\bm{x}_{t}$. We continue this process inductively all the way to the leaves of $\mathcal{T}$. ∎
Example B.3 (for Theorem 3.1).
Consider again the query (9) from Example 5.3. Suppose now that we want to extend our solution from Example 5.3 to the $\text{IVM}^{+}_{\delta}[Q]$ problem. In addition to $Q_{1},P_{1}$ and $Q_{2}^{\prime}$ that are defined in Example 5.3, the proof of Theorem 3.1 for $\text{IVM}^{+}_{\delta}[Q]$ says that we also need to maintain the following relations:
$P_{1}^{\prime\prime}(B,C)$ $=Q_{2}^{\prime}(B,C,D)$
$Q_{1}^{\prime\prime}(A,B,C)$ $=Q_{1}(A,B,C)\wedge P_{1}^{\prime\prime}(B,C)$
The AGM bound of $Q_{1}^{\prime\prime}$ is still $N^{3/2}$, and so is the AGM bound of $P_{1}^{\prime\prime}$. Hence, we can maintain all inserts into them in a total time of $\mathcal{O}(N^{3/2})$. Moreover, note that by definition, $Q_{1}^{\prime\prime}$ must be the projection of $Q(A,B,C,D)$ onto $\{A,B,C\}$.
Now, suppose we have an insert of a tuple $\bm{t}$ into either one of the relations $S,U$ or $V$. Then, we can enumerate the new output tuples that are added due to this insert by starting from $Q_{2}^{\prime}$, just like we did in Example 5.3. However, if we have inserts into $R$ or $T$, then we have to start our enumeration from $Q_{1}^{\prime\prime}$. For each output tuple $(a,b,c)$ of $Q_{1}^{\prime\prime}$ that joins with $\bm{t}$, we enumerate the corresponding $D$-values from $Q_{2}^{\prime}$. There must be at least one $D$-value because $Q_{1}^{\prime\prime}$ is the projection of the output $Q$ onto $\{A,B,C\}$.
B.5. Lower bound for $\text{IVM}^{+}[Q]$ and $\text{IVM}^{+}_{\delta}[Q]$
In this section, we prove Theorem 3.3, which gives a lower bound on the amortized updated time for both $\text{IVM}^{+}[Q]$ and $\text{IVM}^{+}_{\delta}[Q]$ in terms of a lower bound on the static evaluation of $Q$.
Theorem 3.3.
[…]
Proof of Theorem 3.3.
Fix a query $Q$. We first consider $\text{IVM}^{+}[Q]$. Suppose, for sake of contradiction, that there is a constant $\gamma>0$ where $\text{IVM}^{+}[Q]$ can be solved with amortized update time $\widetilde{\mathcal{O}}(N^{\alpha})$ where $\alpha\stackrel{{\scriptstyle\text{def}}}{{=}}\omega(Q)-1-\gamma$. We will show how to use an oracle for $\text{IVM}^{+}[Q]$ to solve $\text{Eval}[Q]$ in time $\widetilde{\mathcal{O}}(|\mathcal{D}|^{\omega(Q)-\gamma}+|Q(\mathcal{D})|)$ on any database instance $\mathcal{D}$. This contradicts the definition of $\omega(Q)$ (Definition 2.2).
We construct a stream of inserts of length $N=|\mathcal{D}|$ where we insert one tuple of $\mathcal{D}$ at a time. We apply these inserts one by one in a total time of $\widetilde{\mathcal{O}}(N^{1+\alpha})$. After the last insert, we invoke constant-delay enumeration of $Q(\mathcal{D})$. Thus, the overall time is $\widetilde{\mathcal{O}}(N^{1+\alpha}+|Q(\mathcal{D})|)$.
For $\text{IVM}^{+}_{\delta}[Q]$, the proof is very similar except that we enumerate $\delta Q(\mathcal{D})$ after every insert. ∎
Appendix C Missing Details in Section 6
In this section, we review some necessary background on intersection joins in the static setting. Let $Q$ be a query, $\overline{Q}$ be its time extension, and $\widehat{Q}$ be its multivariate extension. Suppose we want to evaluate $\overline{Q}$ over a given database instance $\overline{\mathcal{D}}$. This evaluation problem can be reduced (Khamis et al., 2022) to the evaluation of $\widehat{Q}$ over some database instance, $\widehat{\mathcal{D}}$, that satisfies $|\widehat{\mathcal{D}}|=\widetilde{\mathcal{O}}(|\overline{\mathcal{D}}|)$ and can be constructed in time $\widetilde{\mathcal{O}}(|\overline{\mathcal{D}}|)$. Moreover, this reduction is optimal (up to a $\widetilde{\mathcal{O}}(1)$ factor). In particular, the query $\overline{Q}$ is exactly as hard as the hardest component $\widehat{Q}_{\bm{\sigma}}$ in $\widehat{Q}$. This is shown (Khamis et al., 2022) by a backward reduction from each component $\widehat{Q}_{\bm{\sigma}}$ to $\overline{Q}$. We explain both reductions in detail below. Recall that both reductions are in the context of static evaluation. However later on in Section D, we will rely on both reductions in order to prove upper and lower bounds on the IVM problem in the insert-delete setting.
C.1. A reduction from $\overline{Q}$ to $\widehat{Q}$
The reduction in (Khamis et al., 2022) from a time extension query $\overline{Q}$ to the corresponding multivariate extension query $\widehat{Q}$ relies on constructing a segment tree for the interval variable $[Z]$ in $\overline{Q}$.
Segment Trees
Given $N=2^{n}$ for some $n\in\mathbb{N}$, a (discrete) segment tree $\mathcal{T}_{N}$ over $N$ is a complete binary tree whose nodes are discrete intervals that are subsets of $[N]$. A segment tree has the following properties:
  • The leaves from left-to-right are the intervals $[1,1]$, …, $[N,N]$ consisting of single values.
  • Each inner node is the union of its child nodes.
Figure 2 depicts a segment tree $\mathcal{T}_{8}$ for $N=8$. Following (Khamis et al., 2022), we identify each node of a segment tree by a bitstring: The root is identified by the empty string $\varepsilon$; for each inner node that is identified by a bitstring $b$, the left and right child nodes are identified by the bitstrings $b\circ 0$ and respectively $b\circ 1$ (where $\circ$ denotes string concatenation). We denote by $V(\mathcal{T}_{N})$ the set of nodes of the segment tree $\mathcal{T}_{N}$ and by $\text{\sf parent}(v)$ the parent node of a node $v$. Given an interval $x\subseteq[N]$, its canonical partition with respect to $N$ is defined as the set of maximal nodes $y$ of $\mathcal{T}_{N}$ whose corresponding intervals are contained in $x$, i.e.
$\text{\sf CP}_{N}(x)\stackrel{{\scriptstyle\text{def}}}{{=}}\{y\in V(\mathcal{T}_{N})\mid y\subseteq x\text{ but }\text{\sf parent}(y)\not\subseteq x\}$
For any interval $x\subseteq[N]$, $\text{\sf CP}_{N}(x)$ is partition of $x$; see (Khamis et al., 2022). Moreover $|\text{\sf CP}_{N}(x)|=\mathcal{O}(\log N)$. For example, in the segment tree $\mathcal{T}_{8}$ from Figure 2, we have,
$\text{\sf CP}_{8}([1,8])$ =$=\{\varepsilon\}$
$\text{\sf CP}_{8}([2,5])$ =$=\{001,01,100\}$
$[1,8]$$[1,4]$$[5,8]$$[1,2]$$[3,4]$$[5,6]$$[7,8]$$[1,1]$$[2,2]$$[3,3]$$[4,4]$$[5,5]$$[6,6]$$[7,7]$$[8,8]$$\varepsilon$00111100000000010101011010101011111111000000000000001001001001010010010010011011011011100100100100101101101101110110110110111111111111
Figure 2. A segment tree $\mathcal{T}_{8}$. Each node is identified by a bitstring.
[…]
A segment tree $\mathcal{T}_{8}$. Each node is identified by a bitstring.
Lemma C.1 ((Khamis et al., 2022)).
Consider a set of discrete intervals =$\mathcal{I}=\{x_{1},\ldots,x_{k}\}$ with $x_{i}\subseteq[N]$ for some $N\in\mathbb{N}$ for $i\in[k]$. It holds that $\bigcap_{i\in[k]}x_{i}\neq\emptyset$ if and only if there is a permutation $\bm{\sigma}$ of $[k]$ and a tuple $(b_{1},\ldots,b_{k})$ of bitstrings such that
(45) $(b_{1}\circ\cdots\circ b_{j})\in\text{\sf CP}_{N}(x_{\sigma_{j}})\text{ for }j\in[k].$
Moreover, the set of bitstrings $b_{1}\circ\cdots\circ b_{k}$ for all tuples $(b_{1},\ldots,b_{k})$ that satisfy (45) for some permutation $\bm{\sigma}$ of $[k]$ is exactly the canonical partition of the intersection $\bigcap_{i\in[k]}x_{i}$:
[…]
(Recall that $\circ$ is the concatenation operator on strings.) The above lemma is the core of the reduction. It basically says that a set of $k$ intervals $\mathcal{I}$ intersect if and only if their canonical partitions contain $k$ nodes that fall on the same root-to-leaf path in the segment tree. The nodes on this path can appear in any order, and the order is captured by the permutation $\bm{\sigma}$. For example, consider again the segment tree from Figure 2. The intervals $([1,8],[2,5],[3,7])$ overlap. Indeed, the canonical partitions of these intervals contain the nodes $(\varepsilon,01,01)$ respectively, and these nodes fall on the path from the root $\varepsilon$ to the leaf $010$, thus witnessing the intersection. Alternatively, the intersection is also witnessed by the nodes $(\varepsilon,100,10)$, which fall on the path from the root $\varepsilon$ to the leaf $100$. These are the only two witness to the intersection, and they correspond to the two bitstrings $01$ and $100$. Note that $\{01,100\}$ is indeed the canonical partition of $[3,5]$, which is the intersection of the three intervals $([1,8],[2,5],[3,7])$. This is basically what the second part of the lemma says.
Example C.2 ().
Suppose we want to solve the query $\overline{Q}_{\triangle}$ from Example 6.3. We can start by constructing a segment tree for the interval variable $[Z]$ and replacing each value $[z]$ of $[Z]$ by the components in its canonical partition $\text{\sf CP}_{N}([z])$. We can then use Lemma C.1 to look for intersections. In particular, let’s first try to find triples of segment-tree nodes $(z^{R},z^{S},z^{T})$ from $R,S,T$ respectively that appear on the same root-to-leaf path in the specific order $(z^{R},z^{S},z^{T})$. This can only happen if the bitstring $z^{R}$ is a prefix of $z^{S}$ which in turn is a prefix of $z^{T}$. To test this prefix pattern, we can break each $z^{S}$ into a concatenation two strings $z^{S}=z^{S}_{1}\circ z^{S}_{1}$ in all possible ways and similarly break $z^{T}$ into a concatenation of three string $z^{T}=z^{T}_{1}\circ z^{T}_{2}\circ z^{T}_{3}$ and then check whether $z^{R}=z^{S}_{1}=z^{T}_{1}$ and $z^{S}_{2}=z^{T}_{2}$. In particular, we construct the following input relations:
(46) $\widehat{R}_{123}$ =$=\{(z_{1},a,b)\mid\exists[z]:([z],a,b)\in\overline{R}\wedge z_{1}\in\text{\sf CP}_{N}([z])\}$
$\widehat{S}_{123}$ =$=\{(z_{1},z_{2},b,c)\mid\exists[z]:([z],b,c)\in\overline{S}\wedge(z_{1}\circ z_{2})\in\text{\sf CP}_{N}([z])\}$
$\widehat{T}_{123}$ =$=\{(z_{1},z_{2},z_{3},a,c)\mid\exists z:([z],a,c)\in\overline{T}\wedge$
$\quad\quad\quad\quad(z_{1}\circ z_{2}\circ z_{3})\in\text{\sf CP}_{N}([z])\}$
Then we solve the query
(47) $\widehat{Q}_{123}(Z_{1},Z_{2},Z_{3},A,B,C)=\\ \widehat{R}_{123}(Z_{1},A,B)\wedge\widehat{S}_{123}(Z_{1},Z_{2},B,C)\wedge\widehat{T}_{123}(Z_{1},Z_{2},Z_{3},A,C)$
Note that $|\widehat{T}_{123}|=\mathcal{O}(|\overline{T}|\cdot\text{\sf polylog}(N))$ and the same goes for $\widehat{R}_{123}$ and $\widehat{S}_{123}$. This is because the height of the segment tree is $\mathcal{O}(\log N)$. Consider now the database instance depicted in Figure 1(d). Note that the corresponding relation $\widehat{R}_{123}$ above contains (among others) the tuple $(\varepsilon,a_{1},b_{1})$ while $\widehat{S}_{123}$ contains the tuples $(\varepsilon,01,b_{1},c_{1})$ and $\widehat{T}_{123}$ contains the tuple $(\varepsilon,01,\varepsilon,a_{1},c_{1})$. Those tuples join together producing the output tuple $(\varepsilon,01,\varepsilon,a_{1},b_{1},c_{1})$ of $\widehat{Q}_{123}$, which is a witness to the output tuple $([3,5],a_{1},b_{1},c_{1})$ of $\overline{Q}_{\triangle}$.
The query $\widehat{Q}_{123}$ above only considers the specific permutation $(R,S,T)$ of the input relations $\{R,S,T\}$. We have five other permutations corresponding to the following five queries:
$\widehat{Q}_{132}$ $=\widehat{R}_{132}(Z_{1},A,B)\wedge\widehat{S}_{132}(Z_{1},Z_{2},Z_{3},B,C)\wedge\widehat{T}_{132}(Z_{1},Z_{2},A,C)$
$\widehat{Q}_{213}$ $=\widehat{R}_{213}(Z_{1},Z_{2},A,B)\wedge\widehat{S}_{213}(Z_{1},B,C)\wedge\widehat{T}_{213}(Z_{1},Z_{2},Z_{3},A,C)$
$\widehat{Q}_{231}$ $=\widehat{R}_{231}(Z_{1},Z_{2},Z_{3},A,B)\wedge\widehat{S}_{231}(Z_{1},B,C)\wedge\widehat{T}_{231}(Z_{1},Z_{2},A,C)$
$\widehat{Q}_{312}$ $=\widehat{R}_{312}(Z_{1},Z_{2},A,B)\wedge\widehat{S}_{312}(Z_{1},Z_{2},Z_{3},B,C)\wedge\widehat{T}_{312}(Z_{1},A,C)$
(48) $\widehat{Q}_{321}$ $=\widehat{R}_{321}(Z_{1},Z_{2},Z_{3},A,B)\wedge\widehat{S}_{321}(Z_{1},Z_{2},B,C)\wedge\widehat{T}_{321}(Z_{1},A,C)$
Over the database instance depicted in Figure 1(d), note that $\widehat{R}_{132}$ contains the tuple $(\varepsilon,a_{1},b_{1})$ while $\widehat{T}_{132}$ contains $(\varepsilon,10,a_{1},c_{1})$ and $\widehat{S}_{132}$ contains $(\varepsilon,10,0,b_{1},c_{1})$. These tuples join together producing the output tuple $(\varepsilon,10,0,a_{1},b_{1},c_{1})$ of $\widehat{Q}_{132}$, which is a yet another witness to the output tuple $([3,5],a_{1},b_{1},c_{1})$ of $\overline{Q}_{\triangle}$. In particular, the output tuple $([3,5],a_{1},b_{1},c_{1})$ has two different witnesses in $\widehat{Q}_{\triangle}$ specifically because the canonical partition of $[3,5]$ consists of two nodes, namely $01$ and $100$. More generally, any output tuple $([z],a,b,c)$ of $\overline{Q}_{\triangle}$ has as many witnesses in $\widehat{Q}_{\triangle}$ as the number of nodes in the canonical partition of $[z]$. (Recall that for any interval $[z]$, we have $|\text{\sf CP}_{N}([z])|=\mathcal{O}(\log N)$.)
The answer to $\overline{Q}_{\triangle}$ can be retrieved from the union of the answers to the six queries above. The new relations $\widehat{R}_{123},\widehat{S}_{123},\ldots$ can be constructed in time $\mathcal{O}(N\cdot\text{\sf polylog}N)$ where $N$ is the input size to $\overline{Q}_{\triangle}$.
Extrapolating from the above example, we introduce the following definition:
Definition C.3 (Canonical partition of a relation $\text{\sf CP}_{N}^{(i)}(R)$).
Let $\bm{t}=([z],x_{1},\ldots,x_{m})$ be a tuple where $[z]$ is an interval and $x_{1},\ldots,x_{m}$ are points. Given natural numbers $N$ and $i$, the canonical partition of $\bm{t}$ is defined as the following set of tuples
(49) $\text{\sf CP}_{N}^{(i)}([z],x_{1},\ldots,x_{m})\stackrel{{\scriptstyle\text{def}}}{{=}}\\ \{(z_{1},\ldots,z_{i},x_{1},\ldots,x_{m})\mid(z_{1}\circ\cdots\circ z_{i})\in\text{\sf CP}_{N}([z])\}$
We lift the definition of a canonical partition $\text{\sf CP}_{N}^{(i)}$ to relations and define $\text{\sf CP}_{N}^{(i)}(R)$ as the union of the canonical partitions of the tuples in the given relation $R$. We also lift the definition to deltas and define $\text{\sf CP}_{N}^{(i)}(\delta R)$ as the corresponding delta for $\text{\sf CP}_{N}^{(i)}(R)$.
Definition C.4 (Canonical partition of a database instance $\text{\sf CP}_{N}(\overline{\mathcal{D}})$).
Let $Q=R_{1}(\bm{X}_{1})\wedge\cdots\wedge R_{k}(\bm{X}_{k})$ be a query and $\overline{Q}=\overline{R}_{1}([Z],\bm{X}_{1})\wedge\cdots\wedge\overline{R}_{k}([Z],\bm{X}_{k})$ its time extension. Given a permutation $\bm{\sigma}\in\Sigma_{k}$, let $\widehat{Q}_{\bm{\sigma}}$ be the corresponding component of its multivariate extension $\widehat{Q}$:
$\widehat{Q}_{\bm{\sigma}}=\widehat{R}_{\sigma_{1}}(Z_{1},\bm{X}_{\sigma_{1}})\wedge\widehat{R}_{\sigma_{2}}(Z_{1},Z_{2},\bm{X}_{\sigma_{2}})\wedge\cdots\wedge\widehat{R}_{\sigma_{k}}(Z_{1},\ldots,Z_{k},\bm{X}_{\sigma_{k}})$
Given a database instance $\overline{\mathcal{D}}$ for $\overline{Q}$, the canonical partition of $\overline{\mathcal{D}}$ using $\bm{\sigma}$, denoted by $\text{\sf CP}_{N}^{\bm{\sigma}}(\overline{\mathcal{D}})$, is a database instance $\widehat{\mathcal{D}}_{\bm{\sigma}}$ for $\widehat{Q}_{\bm{\sigma}}$ that is defined as:
(50) $\widehat{R}_{\sigma_{i}}\stackrel{{\scriptstyle\text{def}}}{{=}}\text{\sf CP}_{N}^{(i)}(\overline{R}_{\sigma_{i}}),\quad\quad\text{ for }i\in[k]$
Moreover, define $\text{\sf CP}_{N}(\overline{\mathcal{D}})$ to be a database instance for $\widehat{Q}$ which is a combination of $\text{\sf CP}_{N}^{\bm{\sigma}}(\overline{\mathcal{D}})$ for every $\bm{\sigma}\in\Sigma_{k}$:
$\text{\sf CP}_{N}(\overline{\mathcal{D}})\stackrel{{\scriptstyle\text{def}}}{{=}}\left(\text{\sf CP}_{N}^{\bm{\sigma}}(\overline{\mathcal{D}})\right)_{\bm{\sigma}\in\Sigma_{k}}$
We lift the definition of $\text{\sf CP}_{N}^{\bm{\sigma}}$ to deltas and define $\text{\sf CP}_{N}^{\bm{\sigma}}(\delta\overline{\mathcal{D}})$ to be the corresponding delta for $\text{\sf CP}_{N}^{\bm{\sigma}}(\overline{\mathcal{D}})$. We define $\text{\sf CP}_{N}(\delta\overline{\mathcal{D}})$ similarly.
(46) is a special case of (50) above. The following theorem basically says that evaluating $\overline{Q}(\overline{\mathcal{D}})$ is equivalent to evaluating $\widehat{Q}$ on the canonical partition of $\overline{\mathcal{D}}$.
Theorem C.5 (Implicit in (Khamis et al., 2022)).
Given a tuple $(z_{1},\ldots,z_{k},x_{1},\ldots,x_{m})$, define
[…]
We lift the definition of $G_{k}$ to relations and define $G_{k}(R)$ to be the relation resulting from applying $G$ to every tuple in $R$. Then,
(52) $\text{\sf CP}_{N}^{(1)}(\overline{Q}(\overline{\mathcal{D}}))=G_{k}(\widehat{Q}(\text{\sf CP}_{N}(\overline{\mathcal{D}})))$
C.2. A reduction from $\widehat{Q}$ to $\overline{Q}$
In the previous Section C.1, we saw how to reduce the static evaluation of $\overline{Q}$ to $\widehat{Q}$. We now summarize the backward reduction from the static evaluation of each component $\widehat{Q}_{\bm{\sigma}}$ of $\widehat{Q}$ back to $\overline{Q}$, based on (Khamis et al., 2022). Later on, we will use this backward reduction to prove lower bounds on the IVM problem in Section D.4.
Example C.6 ().
Continuing with Example 6.3, let’s take the query $\widehat{Q}_{123}$ from (47). Let $\widehat{\mathcal{D}}_{123}^{\prime}$ be an arbitrary database instance for $\widehat{Q}_{123}$. (Unlike Example 6.3, here we cannot make any assumption about how $\widehat{\mathcal{D}}_{123}^{\prime}$ was constructed.) We show how to use an oracle for the intersection query $\overline{Q}_{\triangle}$ from (11) in order to compute $\widehat{Q}_{123}(\widehat{\mathcal{D}}_{123}^{\prime})$ in the same time. WLOG we can assume that each value $x$ that appears in $\widehat{\mathcal{D}}_{123}^{\prime}$ is a bitstring of length exactly $\ell$ for some constant $\ell$. (If a bitstring has length less than $\ell$, we can pad it with zeros.) We construct a segment tree $\mathcal{T}_{N^{\prime}}$ where $N^{\prime}\stackrel{{\scriptstyle\text{def}}}{{=}}2^{3\ell}$, i.e. a segment tree of depth $3\ell$. (We chose $3$ in this example because it is the number of atoms in $\widehat{Q}_{123}$.) Each node $v$ in the segment tree corresponds to an interval that is contained in $[N^{\prime}]$ and is identified by a bitstring of length at most $3\ell$. Given a bitstring $b$ of length at most $3\ell$, let $\text{\sf seg}_{N^{\prime}}(b)$ denote the corresponding interval in the segment tree $\mathcal{T}_{N^{\prime}}$. We construct the following database instance $\overline{\mathcal{D}}^{\prime}$ for $\overline{Q}_{\triangle}$:
$\overline{R}^{\prime}$ =$=\{(\text{\sf seg}_{N^{\prime}}(z_{1}),a,b)\mid(z_{1},a,b)\in\widehat{R}_{123}^{\prime}\}$
$\overline{S}^{\prime}$ =$=\{(\text{\sf seg}_{N^{\prime}}(z_{1}\circ z_{2}),b,c)\mid(z_{1},z_{2},b,c)\in\widehat{S}_{123}^{\prime}\}$
$\overline{T}^{\prime}$ =$=\{(\text{\sf seg}_{N^{\prime}}(z_{1}\circ z_{2}\circ z_{3}),a,c)\mid(z_{1},z_{2},z_{3},a,c)\in\widehat{T}_{123}^{\prime}\}$
Following (Khamis et al., 2022), we can show that there is a one-to-one mapping between the output tuples of $\overline{Q}_{\triangle}(\overline{\mathcal{D}}^{\prime})$ and the output tuples of $\widehat{Q}_{123}(\widehat{\mathcal{D}}_{123}^{\prime})$. This follows from the observation that the three intervals $\text{\sf seg}_{N^{\prime}}(z_{1}^{R})$, $\text{\sf seg}_{N^{\prime}}(z_{1}^{S}\circ z_{2}^{S})$, $\text{\sf seg}_{N^{\prime}}(z_{1}^{T}\circ z_{2}^{T}\circ z_{3}^{T})$ overlap if and only if $z_{1}^{R}=z_{1}^{S}=z_{1}^{T}$ and $z_{2}^{S}=z_{2}^{T}$. (Recall that the bitstrings $z_{1}^{R},z_{1}^{S},\ldots$ all have the same length $\ell$.) Moreover note that $\overline{R}^{\prime},\overline{S}^{\prime},\overline{T}^{\prime}$ have the same sizes as $\widehat{R}_{123}^{\prime},\widehat{S}_{123}^{\prime},\widehat{\mathcal{T}}_{123}^{\prime}$ respectively and they can be constructed in linear time.
Extrapolating from the above example, we introduce the following concept:
Definition C.7 (Interval version of a database instance $\text{\sf IV}(\widehat{\mathcal{D}}_{\bm{\sigma}})$).
Let $Q=R_{1}(\bm{X}_{1})\wedge\cdots\wedge R_{k}(\bm{X}_{k})$ be a query, $\overline{Q}$ its time extension, $\bm{\sigma}$ be a permutation of $[k]$, and $\widehat{Q}_{\bm{\sigma}}$ be a component of the multivariate extension $\widehat{Q}$ of $Q$. Let $\widehat{\mathcal{D}}_{\bm{\sigma}}$ be an arbitrary database instance for $\widehat{Q}_{\bm{\sigma}}$. Define the interval version of $\widehat{\mathcal{D}}_{\bm{\sigma}}$, denoted by $\text{\sf IV}(\widehat{\mathcal{D}}_{\bm{\sigma}})$, to be a database instance $\overline{D}$ for $\overline{Q}$ that is constructed as follows: WLOG assume each value $x$ that appears in $\widehat{\mathcal{D}}_{\bm{\sigma}}$ is a bitstring of length exactly $\ell$ for some constant $\ell$ (pad with zeros if needed) and let $N\stackrel{{\scriptstyle\text{def}}}{{=}}2^{k\ell}$. Let $\text{\sf seg}_{N}(b)$ map a bitstring $b$ of length at most $k\ell$ to the corresponding interval in the segment tree $\mathcal{T}_{N}$. Then $\text{\sf IV}(\widehat{\mathcal{D}}_{\bm{\sigma}})$ consists of the following relations for $i\in[k]$:
$\overline{R}_{\sigma_{i}}\stackrel{{\scriptstyle\text{def}}}{{=}}\{(\text{\sf seg}_{N}(z_{1}\circ\cdots\circ z_{i}),x_{1},\ldots,x_{m_{i}})\mid\\ (z_{1},\ldots,z_{i},x_{1},\ldots,x_{m_{i}})\in\widehat{R}_{\sigma_{i}}\}$
We lift the definition of IV to deltas and define $\text{\sf IV}(\delta\widehat{\mathcal{D}}_{\bm{\sigma}})$ to be the corresponding delta for $\text{\sf IV}(\widehat{\mathcal{D}}_{\bm{\sigma}})$.
Proposition C.8 ().
For any database instance $\widehat{\mathcal{D}}_{\bm{\sigma}}$ for $\widehat{Q}_{\bm{\sigma}}$, we have
(53) $|\text{\sf IV}(\widehat{\mathcal{D}}_{\bm{\sigma}})|=|\widehat{\mathcal{D}}_{\bm{\sigma}}|$
Moreover, $\text{\sf IV}(\widehat{\mathcal{D}}_{\bm{\sigma}})$ can be computed in linear time in $|\widehat{\mathcal{D}}_{\bm{\sigma}}|$.
The following theorem basically says that evaluating $\widehat{Q}_{\bm{\sigma}}$ on $\widehat{\mathcal{D}}_{\bm{\sigma}}$ is equivalent to evaluating $\overline{Q}$ on the interval version of $\widehat{\mathcal{D}}_{\bm{\sigma}}$.
Theorem C.9 (Implicit in (Khamis et al., 2022)).
Given a tuple $(z_{1},\ldots,z_{k},x_{1},\ldots,x_{m})$, define
[…]
Moreover, we lift $H_{k}$ to relations and define $H_{k}$ of a relation $R$ to be the relation resulting from applying $H_{k}$ to every tuple in $R$. Then,
(55) $\widehat{Q}_{\bm{\sigma}}(\widehat{\mathcal{D}}_{\bm{\sigma}})$ $=H_{k}(\overline{Q}(\text{\sf IV}(\widehat{\mathcal{D}}_{\bm{\sigma}})))$
(56) $|\widehat{Q}_{\bm{\sigma}}(\widehat{\mathcal{D}}_{\bm{\sigma}})|$ $=|\overline{Q}(\text{\sf IV}(\widehat{\mathcal{D}}_{\bm{\sigma}}))|$
Appendix D Missing Details from Section 7
Theorem 3.4 gives an upper bound on the amortized update time for both the $\text{IVM}^{\pm}$ and $\text{IVM}^{\pm}_{\delta}$ problems in terms of $N$, the length of the update stream. Corollary 3.5 tightens this upper bound from $N$ to the current size of the database, $|\mathcal{D}^{(\tau)}|$. In contrast, Theorem 3.6 gives a lower bound on the amortized update time only for $\text{IVM}^{\pm}_{\delta}$ (in terms of $N$ and by extension $|\mathcal{D}^{(\tau)}|$). We prove Theorem 3.4 for $\text{IVM}^{\pm}$ in Section D.1 and for $\text{IVM}^{\pm}_{\delta}$ in Section D.2. We prove Corollary 3.5 in Section D.3. In Section D.4, we prove Theorem 3.6 (for $\text{IVM}^{\pm}_{\delta}$).
D.1. Upper Bound for $\text{IVM}^{\pm}[Q]$
Example 7.1 introduces the high-level idea of our algorithm for $\text{IVM}^{\pm}[Q]$. Algorithm 1 summarizes our algorithmic framework for $\text{IVM}^{\pm}[Q]$ that meets Theorem 3.4. Algorithm 1 assumes that the length $N$ of the update stream is known in advance. However, the following remark explains how to remove this assumption.
Remark D.1 ().
If the number of updates $N$ is not known in advance, we can initialize $N$ to $1$ and keep doubling $N$ every time we reach $N$ updates. We show in the proof of Lemma D.3 that amortized analysis maintains the same overall runtime.
Algorithm 1 Algorithm of $\text{IVM}^{\pm}$[Q]
Inputs
  • $Q=R_{1}(\bm{X}_{1})\wedge\cdots\wedge R_{k}(\bm{X}_{k})$triangleright Input query to $\text{IVM}^{\pm}$
  • An initially empty database $\mathcal{D}$triangleright Database instance for $Q$
  • $N$: The length of the update stream $\triangleright$ See Remark D.1
Initialization
  • Construct the time extension $\overline{Q}$triangleright Definition 6.2
  • Initialize $\overline{\mathcal{D}}$ to be empty $\triangleright$ Database instance for $\overline{Q}$
  • Construct the multivariate extension $\widehat{Q}$triangleright Eq. (3)
  • Initialize $\widehat{\mathcal{D}}$ to be empty $\triangleright$ Database instance for $\widehat{Q}$
  • For each component $\widehat{Q}_{\bm{\sigma}}$ of $\widehat{Q}$
    • Initialize a tree decomposition using Theorem 3.1
Invariants
  • $\widehat{\mathcal{D}}=\text{\sf CP}_{N}(\overline{\mathcal{D}})$$\triangleright$ Definition C.4
Insertion $\delta_{\tau}\mathcal{D}=\{+R_{j}(\bm{t})\}$normal-$\triangleright$ The $\tau$-th update inserts a tuple $\bm{t}$ into $R_{j}$ $\delta_{\tau}\overline{\mathcal{D}}\leftarrow\{+\overline{R}_{j}([\tau,\infty],\bm{t})\}$ normal-$\triangleright$ Insert $([\tau,\infty],\bm{t})$ into $\overline{R}_{j}$ $\overline{\mathcal{D}}\leftarrow\overline{\mathcal{D}}\biguplus\delta_{\tau}\overline{D}$normal-$\triangleright$ Apply $\delta_{\tau}\overline{\mathcal{D}}$ to $\overline{\mathcal{D}}$ $\delta_{\tau}\widehat{\mathcal{D}}\leftarrow\text{\sf CP}_{N}(\delta_{\tau}\overline{\mathcal{D}})$ normal-$\triangleright$ Definition C.4 $\widehat{\mathcal{D}}\leftarrow\widehat{\mathcal{D}}\biguplus\delta_{\tau}\widehat{\mathcal{D}}$ normal-$\triangleright$ Use Theorem 3.1
Deletion $\delta_{\tau}\mathcal{D}=\{-R_{j}(\bm{t})\}$normal-$\triangleright$ The $\tau$-th update deletes a tuple $\bm{t}$ from $R_{j}$ Find $([\tau^{\prime},\infty],\bm{t})\in\overline{R}_{j}$ $\delta_{\tau}\overline{\mathcal{D}}\leftarrow\{-\overline{R}_{j}([\tau^{\prime},\infty],\bm{t}),+\overline{R}_{j}([\tau^{\prime},\tau],\bm{t})\}$ normal-$\triangleright$ Replace $([\tau^{\prime},\infty],\bm{t})$ with $([\tau^{\prime},\tau],\bm{t})$ in $\overline{R}_{j}$ $\overline{\mathcal{D}}\leftarrow\overline{\mathcal{D}}\biguplus\delta_{\tau}\overline{D}$ $\delta_{\tau}\widehat{\mathcal{D}}\leftarrow\text{\sf CP}_{N}(\delta_{\tau}\overline{\mathcal{D}})$ $\widehat{\mathcal{D}}\leftarrow\widehat{\mathcal{D}}\biguplus\delta_{\tau}\widehat{\mathcal{D}}$ normal-$\triangleright$ Proposition D.2
Enumeration
  • Enumerate (using Theorem 3.1)
    $\pi_{\text{\sf vars}(Q)}\bigcup_{\begin{subarray}{c}z_{1},\ldots,z_{k+1}\in\{0,1\}^{*}\\ z_{1}\circ\cdots\circ z_{k+1}=\text{\sf CP}([\tau,\tau])\end{subarray}}\sigma_{Z_{1}=z_{1},\ldots,Z_{k}=z_{k}}\widehat{Q}(\widehat{\mathcal{D}})$
For the rest of this section, we prove that Algorithm 1 meets Theorem 3.4. Fix a query $Q$ together with its time extension $\overline{Q}$ (Definition 6.2) and its multivariate extension $\widehat{Q}$ (Eq. (3)). We split the proof of Theorem 3.4 for $\text{IVM}^{\pm}[Q]$ into three parts. First, we reduce the maintenance of $Q$ to the maintenance of $\overline{Q}$ (Lemma D.1). Then, we reduce the latter problem to the maintenance of $\widehat{Q}$ (Lemma D.3). Finally, we show that $\widehat{Q}$ can be maintained with $\widetilde{\mathcal{O}}(N^{{\mathsf{w}(\widehat{Q})}-1})$ amortized update time, where ${\mathsf{w}(\widehat{Q})}$ is the fractional hypertree width of $\widehat{Q}$ (Lemma D.4). Theorem 3.4 for $\text{IVM}^{\pm}[Q]$ then follows from Lemmas D.1, D.3, and D.4.
D.1.1. From the Maintenance of $Q$ to the Maintenance of $\overline{Q}$
We reduce the maintenance of $Q$ to the maintenance of $\overline{Q}$. Let $\mathcal{D}$ be the database instance for $Q$ and $\overline{\mathcal{D}}$ be the database instance for $\overline{Q}$. As mentioned before, the idea is to use the interval variable $[Z]$ in $\overline{Q}$ to describe the “lifespan” of the tuples in the database $\mathcal{D}$ for $Q$. Given an atom $\overline{R}([Z],\bm{X})$ in $\overline{Q}$ and a tuple $\bm{t}$ over $\bm{X}$, we introduce the following types of updates to relation $\overline{R}$:
  • $+\overline{R}([\tau,\infty],\bm{t})$: Intuitively, this single-tuple insert indicates that the tuple $\bm{t}$ lives in the database from time $\tau$ on.
  • $(-\overline{R}([\tau^{\prime},\infty],\bm{t}),+\overline{R}([\tau^{\prime},\tau],\bm{t})$) for some $\tau^{\prime}<\tau$. This compound update is called truncation and replaces $([\tau^{\prime},\infty],\bm{t})$ in $\overline{R}$ with $([\tau^{\prime},\tau],\bm{t})$. Intuitively, it signals that the lifespan of tuple $\bm{t}$ ends at time $\tau$.
We call an update stream $\delta_{1}\overline{\mathcal{D}},\delta_{2}\overline{\mathcal{D}},\ldots$ for $\overline{\mathcal{D}}$ restricted if each update $\delta_{\tau}\overline{\mathcal{D}}$ is an insert of the form $+\overline{R}([\tau,\infty],\bm{t})$ or a truncation of the form $(-\overline{R}([\tau^{\prime},\infty],\bm{t}),+\overline{R}([\tau^{\prime},\tau],\bm{t}))$. The restricted result of $\overline{Q}$ after update $\delta_{\tau}\overline{\mathcal{D}}$ is defined as all output tuples of $\overline{Q}$ whose $[Z]$-interval contains $\tau$:
(57) $\overline{Q}_{\tau}(\overline{\mathcal{D}}^{(\tau)})\stackrel{{\scriptstyle\text{def}}}{{=}}\sigma_{\tau\in[Z]}\overline{Q}(\overline{\mathcal{D}}^{(\tau)}),$
Note that by construction, we have
(58) $Q(\mathcal{D}^{(\tau)})=\pi_{\text{\sf vars}(Q)}\overline{Q}_{\tau}(\overline{\mathcal{D}}^{(\tau)})$
We introduce the maintenance problem for the intersection query $\overline{Q}$:
Problem: $\overline{\text{IVM}}[\overline{Q}]$ Parameter: A time extension query $\overline{Q}$ Given: Restricted update stream $\delta_{1}\overline{\mathcal{D}},\delta_{2}\overline{\mathcal{D}},\ldots$ Task: Ensure constant-delay enumeration of $\overline{Q}_{\tau}(\overline{\mathcal{D}}^{(\tau)})$ after each update $\delta_{\tau}\overline{\mathcal{D}}$ We make the reduction from $\text{IVM}^{\pm}[Q]$ to $\overline{\text{IVM}}[\overline{Q}]$ precise:
Lemma D.1 ().
[…]
Proof.
The reduction from $\text{IVM}^{\pm}[Q]$ to $\overline{\text{IVM}}[\overline{Q}]$ works as follows. Consider an update stream $\delta_{1}\mathcal{D},\delta_{2}\mathcal{D},\ldots$ for $Q$. For each $\tau\geq 1$, we pursue the following strategy. If $\delta_{\tau}\mathcal{D}$ is an insert $+R(\bm{t})$, we execute the insert $+\overline{R}([\tau,\infty],\bm{t})$ on the database $\overline{\mathcal{D}}$ of $\overline{Q}$. If it is a delete $-R(\bm{t})$, we apply the truncation $(-\overline{R}([\tau^{\prime},\infty],\bm{t}),+\overline{R}([\tau^{\prime},\tau],\bm{t}))$. For any enumeration request to $Q$ that comes immediately after some update $\delta_{\tau}\mathcal{D}$, we trigger the enumeration procedure for $\overline{Q}$, which enumerates the tuples in $\overline{Q}_{\tau}(\overline{\mathcal{D}}^{(\tau)})$ with constant delay. (Recall (58).) By construction, the tuples in $\overline{Q}_{\tau}(\overline{\mathcal{D}}^{(\tau)})$ must be distinct with respect to their projections onto their point variables (i.e. $\text{\sf vars}(Q)$). In particular,
(59) $|Q(\mathcal{D}^{(\tau)})|=|\overline{Q}_{\tau}(\overline{\mathcal{D}}^{(\tau)})|$
Hence, by reporting the projections of the tuples in $\overline{Q}_{\tau}(\overline{\mathcal{D}}^{(\tau)})$ onto $\text{\sf vars}(Q)$, we enumerate the output $Q(\mathcal{D}^{(\tau)})$ with constant delay. ∎
The following proposition will be useful later for the proof of Theorem 3.4. It basically says that when we truncate a tuple from $\overline{R}([\tau^{\prime},\infty],\bm{t})$ to $\overline{R}([\tau^{\prime},\tau],\bm{t})$, the truncated tuple is still going to join with all other tuples that it used to join with before the truncation.
Proposition D.2 ().
Consider a truncation update $\delta_{\tau}\overline{\mathcal{D}}=$$\{-\overline{R}_{j}([\tau^{\prime},\infty],\bm{t}_{j}),$ $+\overline{R}_{j}([\tau^{\prime},\tau],\bm{t}_{j})\}$. Suppose that the tuple before truncation $\overline{R}_{j}([\tau^{\prime},\infty],\bm{t}_{j})$ used to join with a set of tuples in $\{\overline{R}_{i}([\alpha_{i},\beta_{i}],\bm{t}_{i})\mid i\in S\}$ for some $S\subseteq[k]$, i.e.
$\Join_{i\in S}\overline{R}_{i}([\alpha_{i},\beta_{i}],\bm{t}_{i})\Join\overline{R}_{j}([\tau^{\prime},\infty],\bm{t}_{j})=([\alpha,\beta],\bm{t})$
Then, the following hold:
  1. (1)
    $\beta_{i}=\infty$ for all $i\in S$ and $\beta=\infty$
  2. (2)
    The truncated tuple $\overline{R}_{j}([\tau^{\prime},\tau],\bm{t}_{j})$ still joins with the same set of tuples. Namely
    $\Join_{i\in S}\overline{R}_{i}([\alpha_{i},\beta_{i}],\bm{t}_{i})\Join\overline{R}_{j}([\tau^{\prime},\tau],\bm{t}_{j})=([\alpha,\tau],\bm{t})$
Proof.
Right before time $\tau$ when we are about to truncate, the database $\overline{\mathcal{D}}^{(\tau-1)}$ does not yet contain any interval $[\alpha_{i},\beta_{i}]$ where $\alpha_{i}\geq\tau$. This is because $\tau$ is the current time, and those intervals $[\alpha_{i},\beta_{i}]$ where $\alpha_{i}\geq\tau$ start in the future. This proves part 2. Moreover, every interval $[\alpha_{i},\beta_{i}]$ in the database $\overline{\mathcal{D}}^{(\tau-1)}$ where $\alpha_{i}<\tau\leq\beta_{i}$ must satisfy $\beta_{i}=\infty$. This is because such an interval ends in the future, thus we could not have encountered its end yet. This proves part 1. ∎
D.1.2. From the Maintenance of $\overline{Q}$ to the Maintenance of $\widehat{Q}$
Now, we give a reduction from the maintenance of $\overline{Q}$ to the maintenance of $\widehat{Q}$. Let $\widehat{\mathcal{D}}$ be the database instance for $\widehat{Q}$. In order to translate an update $\delta_{\tau}\overline{\mathcal{D}}$ for $\overline{Q}$ into an update $\delta_{\tau}\widehat{\mathcal{D}}$ for $\widehat{Q}$, we take the canonical partition of $\delta_{\tau}\overline{\mathcal{D}}$, as given by Definition C.4. In particular, suppose that $N$, the length of the update stream, is a power of 2 and is known in advance. (See Remark D.1 about removing this assumption.) The corresponding update to $\widehat{Q}$ is
(60) $\delta_{\tau}\widehat{\mathcal{D}}\leftarrow\text{\sf CP}_{N}(\delta_{\tau}\overline{\mathcal{D}})$
We say that a stream $\delta_{1}\widehat{\mathcal{D}},\delta_{2}\widehat{\mathcal{D}},\ldots$ is a restricted canonical update stream for $\widehat{Q}$ if there is a restricted update stream $\delta_{1}\overline{\mathcal{D}},\delta_{2}\overline{\mathcal{D}},\ldots$ for $\overline{Q}$ such that each $\delta_{\tau}\widehat{\mathcal{D}}$ is the canonical partition of $\delta_{\tau}\overline{\mathcal{D}}$ with respect to $N$. The restricted result of $\widehat{Q}$ after update $\delta_{\tau}\widehat{\mathcal{D}}$ is defined as the set of all output tuples of $\widehat{Q}$ where the bitstring $Z_{1}\circ\cdots\circ Z_{k}$ (when interpreted as an interval in the segment tree $\mathcal{T}_{N}$) contains $\tau$. Formally,
(63) $\widehat{Q}_{\tau}(\widehat{\mathcal{D}}^{(\tau)})\stackrel{{\scriptstyle\text{def}}}{{=}}\bigcup_{\begin{subarray}{c}z_{1},\ldots,z_{k+1}\in\{0,1\}^{*}\\ z_{1}\circ\cdots\circ z_{k+1}=\text{\sf CP}([\tau,\tau])\end{subarray}}\sigma_{Z_{1}=z_{1},\ldots,Z_{k}=z_{k}}\widehat{Q}(\widehat{\mathcal{D}}^{(\tau)})$
We are ready to introduce the maintenance problem for $\widehat{Q}$:
Problem: $\widehat{\text{IVM}}[\widehat{Q}]$ Parameter: A multivariate extension query $\widehat{Q}$ Given: Restricted canonical update stream $\delta_{1}\widehat{\mathcal{D}},\delta_{2}\widehat{\mathcal{D}},\ldots$ Task: Ensure constant-delay enumeration of $\widehat{Q}_{\tau}(\widehat{\mathcal{D}}^{(\tau)})$ after each update $\delta_{\tau}\widehat{\mathcal{D}}$
We formalize the reduction from $\overline{\text{IVM}}[\overline{Q}]$ to $\widehat{\text{IVM}}[\widehat{Q}]$:
Lemma D.3 ().
[…]
Proof.
The reduction expressed in Lemma D.3 works as follows. Consider an update stream $\delta_{1}\overline{\mathcal{D}},\delta_{2}\overline{\mathcal{D}}\ldots$ for $\overline{Q}$. We transform each update $\delta_{\tau}\overline{\mathcal{D}}$ into its canonical partition $\delta_{\tau}\widehat{\mathcal{D}}$ with respect to $N$, and apply $\delta_{\tau}\widehat{\mathcal{D}}$ on the database $\widehat{\mathcal{D}}$ for $\widehat{Q}$. As a result, the algorithm maintains the invariant that the database $\widehat{\mathcal{D}}$ is always the canonical partition of the database $\overline{\mathcal{D}}$ with respect to $N$. Formally,
$\widehat{\mathcal{D}}^{(\tau)}=\text{\sf CP}_{N}(\overline{\mathcal{D}}^{(\tau)}),\quad\quad\text{for all $\tau\in[N]$}.$
Moreover, by Definition C.3, the number of single-tuple updates in $\delta_{\tau}\widehat{\mathcal{D}}$ is at most polylogarithmic in $N$ (which is $\widetilde{\mathcal{O}}(1)$).
Theorem C.5 connects the full outputs of $\overline{Q}(\overline{\mathcal{D}}^{(\tau)})$ and $\widehat{Q}(\widehat{\mathcal{D}}^{(\tau)})$ using the following relation:
(64) $\text{\sf CP}_{N}^{(1)}(\overline{Q}(\overline{\mathcal{D}}^{(\tau)}))=G_{k}(\widehat{Q}(\widehat{\mathcal{D}}^{(\tau)}))$
From the above equation along with (63) and (57), we conclude that:
(65) $\pi_{\text{\sf vars}(Q)}\widehat{Q}_{\tau}(\widehat{\mathcal{D}}^{(\tau)})$ $=\pi_{\text{\sf vars}(Q)}\overline{Q}_{\tau}(\overline{\mathcal{D}}^{(\tau)})$
(66) $|\widehat{Q}_{\tau}(\widehat{\mathcal{D}}^{(\tau)})|$ $=|\overline{Q}_{\tau}(\overline{\mathcal{D}}^{(\tau)})|$
Equation (66) follows from the fact that for any segment $[Z]$ where $\tau\in[Z]$, there is exactly one segment in the canonical partition of $[Z]$ that contains $\tau$. Hence, we can use constant-delay enumeration for $\widehat{Q}_{\tau}(\widehat{\mathcal{D}}^{(\tau)})$ in order to enumerate the output of $\overline{Q}_{\tau}(\overline{\mathcal{D}}^{(\tau)})$.
Finally, if the length of the update stream $N$ is not known in advance, then we use the doubling trick described in Remark D.1. In particular, we initialize $N$ to 1. Before we process $\tau$-th update for $\tau=N+1$, we replace the segment tree $\mathcal{T}_{N}$ by $\mathcal{T}_{2N}$ and replace $N$ by $2N$. Replacing the segment tree corresponds to creating a copy of the old tree and adding a new root node whose two children are the old segment tree and the copy. After updating the segment tree, we can redo the previous inserts into the new tree. The total update time would be
$\widetilde{\mathcal{O}}(1)\sum_{i\in[\lceil\log_{2}N\rceil]}2^{i}\cdot f(2^{i})=\widetilde{\mathcal{O}}(N\cdot f(N))$
(Recall that $f$ is monotonic.) Hence, the amortized update time is still $\widetilde{\mathcal{O}}(f(N))$. ∎
D.1.3. Update Time for $\widehat{Q}$
Note that (58),  (59), (65), and (66) imply:
$Q(\mathcal{D}^{(\tau)})$ $=\pi_{\text{\sf vars}(Q)}\widehat{Q}_{\tau}(\widehat{\mathcal{D}}^{(\tau)})$
$|Q(\mathcal{D}^{(\tau)})|$ $=|\widehat{Q}_{\tau}(\widehat{\mathcal{D}}^{(\tau)})|$
This basically means that we can do constant-delay enumeration for $Q(\mathcal{D}^{(\tau)})$ using constant-delay enumeration for $\widehat{Q}_{\tau}(\widehat{\mathcal{D}}^{(\tau)})$. Now, we prove the following statement:
Lemma D.4 ().
[…]
Before we prove the above lemma, we need some preliminaries:
Definition D.5 ().
A subset $\bm{Y}\subseteq\text{\sf vars}(\widehat{Q})$ is called $Z$-prefix-closed if it satisfies the following property: For every $Z_{i}\in\bm{Y}$, we must have $\{Z_{1},\ldots,Z_{i-1}\}\subseteq\bm{Y}$. An atom $R(\bm{Y})$ is $Z$-prefix-closed if its schema $\bm{Y}$ is $Z$-prefix-closed.
Recall from Appendix A that a tree decomposition $(\mathcal{T},\chi)$ of a query $Q$ is called optimal if it minimizes the fractional hypertree width of $Q$.
Proposition D.6 ().
Let $\widehat{Q}_{\bm{\sigma}}$ be a component of $\widehat{Q}$. There must exist an optimal tree decomposition $(\mathcal{T},\chi)$ for $\widehat{Q}_{\bm{\sigma}}$ where every bag of $(\mathcal{T},\chi)$ is $Z$-prefix-closed.
Proof of Proposition D.6.
The proof is based on the equivalence of tree decompositions and variable elimination orders. See Appendix A for some background. In particular, the proof works as follows. Pick some optimal variable elimination order $\bm{\mu}$. We inductively prove the following:
Claim D.2 ().
Throughout variable elimination, every atom is $Z$-prefix-closed.
Claim D.3 ().
For every variable $X\in\bm{\mu}$, the set $\bm{U}_{X}^{\bm{\mu}}$ is prefix-closed.
Initially, both claims above hold. Now, suppose we pick a variable $X$ to eliminate. By induction, all atoms that contain $X$ are $Z$-prefix-closed, hence $\bm{U}_{X}^{\bm{\mu}}$ is $Z$-prefix-closed. Now in order to show that the newly added atom $R_{X}^{\bm{\mu}}(\bm{U}_{X}^{\bm{\mu}}-\{X\})$ can be assumed to be $Z$-prefix closed, we recognize two cases. If $X$ is not in $\{Z_{1},\ldots,Z_{k}\}$, then $\bm{U}_{X}^{\bm{\mu}}-\{X\}$ is obviously $Z$-prefix-closed. On the other hand, suppose $X=Z_{i}$ for some $i\in[k]$. Let $j$ be the largest index where $Z_{j}\in\bm{U}_{X}^{\bm{\mu}}$. If $Z_{j}=Z_{i}$, then $\bm{U}_{X}^{\bm{\mu}}-\{X\}$ is $Z$-prefix closed. Otherwise, $\bm{U}_{X}^{\bm{\mu}}$ must contain $Z_{i},Z_{i+1},\ldots,Z_{j}$. Moreover, none of these variables can occur in any atom other than $R_{X}^{\bm{\mu}}(\bm{U}_{X}^{\bm{\mu}}-\{X\})$. Hence, there is an optimal variable elimination order $\bm{\mu}^{\prime}$ that eliminates them in any order. We choose the order $Z_{j},Z_{j-1},\ldots,Z_{i}$. ∎
Proof of Lemma D.4.
The proof of Lemma D.4 relies on the proof of Theorem 3.1 that is given in Section B.3. Consider an optimal tree decomposition $(\mathcal{T},\chi)$ for $\widehat{Q}$, i.e. where $\mathsf{w}(\mathcal{T},\chi)=\mathsf{w}(\widehat{Q})$. Let $\delta_{1}\widehat{\mathcal{D}},\delta_{2}\widehat{\mathcal{D}},\ldots$ be a restricted canonical update stream for $\widehat{Q}$. Recall that each $\delta_{\tau}\widehat{\mathcal{D}}$ is the canonical partition of a corresponding update $\delta_{\tau}\overline{\mathcal{D}}$ to $\overline{\mathcal{D}}$. Consider an update $\delta_{\tau}\widehat{\mathcal{D}}=\text{\sf CP}_{N}(\delta_{\tau}\overline{\mathcal{D}})$ in this stream. First, assume that $\delta_{\tau}\overline{\mathcal{D}}$ is an insert, i.e. $\delta_{\tau}\overline{\mathcal{D}}=\{+\overline{R}_{j}([\tau,\infty],\bm{t}_{j})\}$ for some $j\in[k]$ and tuple $\bm{t}_{j}$ with schema $\bm{X}_{j}$. In this case, the corresponding $\delta_{\tau}\widehat{\mathcal{D}}=\text{\sf CP}_{N}(\delta_{\tau}\overline{\mathcal{D}})$ is a set of single-tuple inserts, and we can apply these inserts to the materialized bags of the tree decomposition $(\mathcal{T},\chi)$, just as explained in the proof of Theorem 3.1. Truncation is more tricky. In particular, suppose that $\delta_{\tau}\overline{\mathcal{D}}$ is a truncation, i.e. $\delta_{\tau}\overline{\mathcal{D}}=\{-\overline{R}_{j}([\tau^{\prime},\infty],\bm{t}_{j}),+\overline{R}_{j}([\tau^{\prime},\tau],\bm{t}_{j})\}$ for some $\tau^{\prime}<\tau$, $j\in[m]$ and tuple $\bm{t}_{j}$ over schema $\bm{X}_{j}$. In this case, we update the materialized bags of $(\mathcal{T},\chi)$ using the updates given by $\text{\sf CP}_{N}(\delta_{\tau}\overline{\mathcal{D}})$. Some of these updates are inserts, namely $\text{\sf CP}_{N}(\{+\overline{R}_{j}([\tau^{\prime},\tau],\bm{t}_{j})\})$, and these are accounted for by the analysis from the proof of Theorem 3.1. In particular, they are covered by the $\widetilde{\mathcal{O}}(N^{{\mathsf{w}(\widehat{Q})}})$ total update time over all inserts combined. However, some of the updates in $\text{\sf CP}_{N}(\delta_{\tau}\overline{\mathcal{D}})$ are deletes, namely $\text{\sf CP}_{N}(\{-\overline{R}_{j}([\tau^{\prime},\infty],\bm{t}_{j})\})$, and in order to account for those, we will use amortized analysis. In particular, we will deposit extra credits with every insert so that we can use them later to pay for the deletes during truncation. We explain this argument in detail below.
We start with the following claim, which is very similar in spirit to Proposition D.2:
Claim D.4 ().
Let $\delta_{\tau}\overline{\mathcal{D}}=\{+\overline{R}_{j}([\tau,\infty],\bm{t}_{j})\}$ be an insert into $\overline{\mathcal{D}}$, for some $j\in[m]$ and tuple $\bm{t}_{j}$ over the schema $\bm{X}_{j}$. For some $i\in[k]-\{j\}$, let $([\alpha,\beta],\bm{t}_{i})$ be a tuple of another relation $\overline{R}_{i}$ that joins with the inserted tuple $([\tau,\infty],\bm{t}_{j})$ of $\overline{R}_{j}$. Then, we must have $\alpha<\tau$ and $\beta=\infty$. Namely, the interval $[\alpha,\beta]$ strictly contains $[\tau,\infty]$.
Proof of Claim D.4.
The above claim follows from the fact that $\tau$ is the current time, and any interval $[\alpha,\beta]$ where $\alpha\geq\tau$ would start in the future, hence it cannot exist yet in $\overline{\mathcal{D}}$. Moreover, in order for $[\alpha,\beta]$ to overlap with $[\tau,\infty]$, it must be the case that $\beta\geq\tau$, which implies that $\beta=\infty$ because it is in the future. This proves the claim. ∎
For the purpose of analyzing the algorithm, we distinguish between two types of tuples in $\overline{\mathcal{D}}$: finalized tuples and provisional tuples. Finalized tuples cannot be deleted once they are inserted, while provisional tuples can be deleted later. Also for the purpose of analyzing the algorithm, we slightly modify the way the algorithm handles inserts. In particular, whenever we have an insert $+\overline{R}_{j}([\tau,\infty],\bm{t}_{j})$ into $\overline{\mathcal{D}}$, we insert this tuple as provisional (since it might be deleted later due to truncation) and we insert an additional tuple $([\tau,\tau],\bm{t}_{j})$ into $\overline{R}_{j}$ that is finalized. In particular, we know for sure that the tuple $([\tau,\tau],\bm{t}_{j})$ will never be deleted later because the lifespan of the tuple $\bm{t}_{j}$ extends at least until the current time $\tau$. We also update $\widehat{\mathcal{D}}$ by applying the two corresponding inserts, namely $\text{\sf CP}_{N}(\{+\overline{R}_{j}([\tau,\infty],\bm{t}_{j})\})$ and $\text{\sf CP}_{N}(\{+\overline{R}_{j}([\tau,\tau],\bm{t}_{j})\})$. Tuples in $\widehat{\mathcal{D}}$ inherit the provisional/finalized property from the underlying tuples in $\overline{\mathcal{D}}$. We make the following claim:
Claim D.5 ().
The total time needed to apply the insert $\text{\sf CP}_{N}(\{+\overline{R}_{j}([\tau,\infty],\bm{t}_{j})\})$ into $\widehat{\mathcal{D}}$ is within a $\widetilde{\mathcal{O}}(1)$ factor of the total time needed to apply the insert $\text{\sf CP}_{N}(\{+\overline{R}_{j}([\tau,\tau],\bm{t}_{j})\})$.
The above claim follows from Claim D.4 along with the fact that every bag in the tree decomposition $(\mathcal{T},\chi)$ is $Z$-prefix closed (Proposition D.6). In particular, Claim D.4 basically says that the two intervals $[\tau,\infty]$ and $[\tau,\tau]$ are “indistinguishable” as far as joins are concerned. Proposition D.6 ensures that whenever we take a projection in the query plan corresponding to the tree decomposition $(\mathcal{T},\chi)$, we are basically replacing these two intervals with two new intervals that contain them. Hence, the new intervals remain indistinguishable.
Now, we can analyze the total update time as follows. Claim D.5 implies that the total time needed to apply provisional inserts is upper bounded by the total time needed to apply finalized inserts. Hence, the overall runtime is dominated by finalized inserts. Moreover, every time we apply a provisional insert, we can deposit one credit with every tuple that we insert into any bag of $(\mathcal{T},\chi)$, in order to pay for the potential delete of this tuple in the future. These credits cover the cost of all deletes of provisional tuples. (Recall that finalized tuples cannot be deleted.) Finally, the total time needed to apply all finalized inserts is bounded by $\widetilde{\mathcal{O}}(N^{\mathsf{w}(\widehat{Q})})$, by Theorem 3.1.
We are left with explaining how to support constant-delay enumeration of $\widehat{Q}_{\tau}(\widehat{\mathcal{D}}^{(\tau)})$. To that end, we enumerate selections $\sigma_{Z_{1}=z_{1},\ldots,Z_{k}=z_{k}}\widehat{Q}(\widehat{\mathcal{D}}^{(\tau)})$ for various assignments $Z_{1}=z_{1},$ $\ldots$,… , $Z_{k}=z_{k}$, as defined by (63). Luckily, for every component $\widehat{Q}_{\bm{\sigma}}$ of $\widehat{Q}$, there is one input relation $\widehat{R}_{\sigma_{k}}$ that contains the variables $\{Z_{1},\ldots,Z_{k}\}$; see (2). Hence, every tree decomposition of $\widehat{Q}_{\bm{\sigma}}$ must contain a bag $B$ that contains all the variables $Z_{1},\ldots,Z_{k}$. In order to enumerate a selection $\sigma_{Z_{1}=z_{1},\ldots,Z_{k}=z_{k}}\widehat{Q}(\widehat{\mathcal{D}}^{(\tau)})$ with constant delay, we can use the bag $B$ as the root of the tree decomposition. ∎
D.2. Upper Bound for $\text{IVM}^{\pm}_{\delta}[Q]$
In the previous section, we proved Theorem 3.4 for the $\text{IVM}^{\pm}[Q]$ problem for a given query $Q$. In this section, we extend the proof and the corresponding algorithm from the previous section in order to prove Theorem 3.4 for the $\text{IVM}^{\pm}_{\delta}[Q]$ problem.
Here are the main changes to Algorithm 1 in order to support $\text{IVM}^{\pm}_{\delta}[Q]$, i.e. support constant-delay enumeration of $\delta_{\tau}Q(\mathcal{D})$ after every update $\delta_{\tau}\mathcal{D}$:
  • Insertion $\delta_{\tau}\mathcal{D}=\{+R_{j}(\bm{t})\}$: In addition to applying the insert $\{+\overline{R}_{j}([\tau,\infty],\bm{t})\}$ into $\overline{\mathcal{D}}$, we now also insert $+\overline{R}_{j}([\tau,\tau],\bm{t})$.
  • Deletion $\delta_{\tau}\mathcal{D}=\{-R_{j}(\bm{t})\}$: In addition to applying the update $\{-\overline{R}_{j}([\tau^{\prime},\infty],\bm{t}),+\overline{R}_{j}([\tau^{\prime},\tau],\bm{t})\}$ to $\overline{\mathcal{D}}$, we now also apply the insert $+\overline{R}_{j}([\tau,\tau],\bm{t})$.
  • Enumeration Suppose we want to enumerate $\delta_{\tau}Q(\mathcal{D})$ after an update $\delta_{\tau}\mathcal{D}=sR_{j}(\bm{t})$ where $s\in\{+,-\}$. Then, we enumerate the following:
    (69) $\delta_{\tau}Q(\mathcal{D})=s\left(\pi_{\text{\sf vars}(Q)}\bigcup_{\begin{subarray}{c}z_{1},\ldots,z_{k}\in\{0,1\}^{*}\\ z_{1}\circ\cdots\circ z_{k}=\text{\sf CP}([\tau,\tau])\end{subarray}}\sigma_{Z_{1}=z_{1},\ldots,Z_{k}=z_{k}}\widehat{Q}(\widehat{\mathcal{D}})\right)$
We call the above modified version of Algorithm 1 the delta version of Algorithm 1. Before we prove that the delta version of Algorithm 1 proves Theorem 3.4 for $\text{IVM}^{\pm}_{\delta}[Q]$, we need some preliminaries.
Proposition D.7 ().
Given a discrete interval $[\alpha,\beta]$ for some $\alpha,\beta\in\mathbb{N}$, let $N$ be a power of 2 such that $N\geq\beta$. If $\text{\sf CP}_{N}([\alpha,\beta])$ contains a singleton interval $[x,x]$ for some $x\in\mathbb{N}$, then $x$ is an endpoint of $[\alpha,\beta]$. In particular, either $x=\alpha$ or $x=\beta$.
Proposition D.7 follows from the definition of a segment tree $\mathcal{T}_{N}$ and canonical partition $\text{\sf CP}_{N}$. For example, consider the segment tree $\mathcal{T}_{8}$ depicted in Figure 2. Let’s take the segment $[2,5]$. Its canonical partition consists of $\{[2,2],[3,4],[5,5]\}$. Note that only $[2,2]$ and $[5,5]$ are singleton intervals. But $2$ and $5$ are the endpoints of the interval $[2,5]$. The above proposition says that we cannot find in the canonical partition of any segment $[\alpha,\beta]$ any singleton intervals that are not end points of $[\alpha,\beta]$.
The following theorem implies Theorem 3.4 for $\text{IVM}^{\pm}_{\delta}[Q]$:
Theorem D.8 ().
[…]
Proof.
The proof is very similar to the proof of Theorem 3.4 from Section D.1 for $\text{IVM}^{\pm}[Q]$. For every insert or delete $\delta_{\tau}\mathcal{D}=\{sR_{j}(\bm{t})\}$ where $s\in\{+,-\}$, we now have to insert an extra tuple $+\overline{R}_{j}([\tau,\tau],\bm{t})$ into $\overline{\mathcal{D}}$. This only increases the size of $\overline{\mathcal{D}}$ by a constant factor. Hence, it doesn’t affect the overall runtime.
Next, we prove correctness. In particular, our target is to prove that (69) does indeed compute $\delta_{\tau}Q(\mathcal{D})$, as defined by (6). To that end, we define the following:
(70) $\overline{Q}_{\delta_{\tau}}(\overline{\mathcal{D}}^{(\tau)})$ =$\stackrel{{\scriptstyle\text{def}}}{{=}}\sigma_{[Z]=[\tau,\tau]}\overline{Q}(\overline{\mathcal{D}}^{(\tau)})$
(73) $\widehat{Q}_{\delta_{\tau}}(\widehat{\mathcal{D}}^{(\tau)})$ =$\stackrel{{\scriptstyle\text{def}}}{{=}}\bigcup_{\begin{subarray}{c}z_{1},\ldots,z_{k}\in\{0,1\}^{*}\\ z_{1}\circ\cdots\circ z_{k}=\text{\sf CP}([\tau,\tau])\end{subarray}}\sigma_{Z_{1}=z_{1},\ldots,Z_{k}=z_{k}}\widehat{Q}(\widehat{\mathcal{D}}^{(\tau)})$
Compare the above to (57) and (63) respectively. The correctness proof goes through two steps. First, we prove that
$\delta_{\tau}Q(\mathcal{D})=s\left(\pi_{\text{\sf vars}(Q)}\overline{Q}_{\delta_{\tau}}(\overline{\mathcal{D}}^{(\tau)})\right)$
where recall that $s\in\{+,-\}$ depending on the update $\delta_{\tau}\mathcal{D}=\{sR_{j}(\bm{t})\}$. This follows from the fact that $\tau$ is the unique timestamp for the update $\delta_{\tau}\mathcal{D}$. In particular, all corresponding updates $\overline{s}\overline{R}_{j}([\alpha,\beta],\bm{t})$ in $\delta_{\tau}\overline{\mathcal{D}}$ (where $\overline{s}\in\{+,-\}$) must have $\tau\in[\alpha,\beta]$. Moreover, no other tuple $([\alpha^{\prime},\beta^{\prime}],\bm{t}^{\prime})$ in $\overline{R}_{j}$ can have $\tau\in\{\alpha,\beta\}$ (because the timestamp $\tau$ is unique to this update).
The second step of the correctness proof is to prove that
(74) $\pi_{\text{\sf vars}(Q)}\overline{Q}_{\delta_{\tau}}(\overline{\mathcal{D}}^{(\tau)})=\pi_{\text{\sf vars}(Q)}\widehat{Q}_{\delta_{\tau}}(\widehat{\mathcal{D}}^{(\tau)})$
The containment $\subseteq$ in (74) holds thanks to the extra intervals$+\overline{R}_{j}([\tau,\tau],\bm{t})$ that we are inserting to $\overline{\mathcal{D}}$ in the delta version of Algorithm 1. The other containment $\supseteq$ in (74) follows from Proposition D.7. In particular, (73) only selects the leaf corresponding to $[\tau,\tau]$ in the segment tree, but Proposition D.7 tells us that this leaf could not belong to the canonical partition of any interval $[\alpha,\beta]$ unless $\tau\in\{\alpha,\beta\}$. Such intervals could only have been added to $\overline{\mathcal{D}}$ during the update $\delta_{\tau}\mathcal{D}$ because the timestamp $\tau$ is unique to this update.
Finally, we show that constant-delay enumeration for $\widehat{Q}_{\delta_{\tau}}(\widehat{\mathcal{D}}^{(\tau)})$ implies constant-delay enumeration for $\delta_{\tau}Q(\mathcal{D})$. To that end, we prove:
$|\widehat{Q}_{\delta_{\tau}}(\widehat{\mathcal{D}}^{(\tau)})|=|\delta_{\tau}Q(\mathcal{D})|$
We also prove that in two steps, namely
(75) $|\overline{Q}_{\delta_{\tau}}(\overline{\mathcal{D}}^{(\tau)})|$ $=|\delta_{\tau}Q(\mathcal{D})|$
(76) $|\widehat{Q}_{\delta_{\tau}}(\widehat{\mathcal{D}}^{(\tau)})|$ $=|\overline{Q}_{\delta_{\tau}}(\overline{\mathcal{D}}^{(\tau)})|$
Equation (75) follows from (70) along with the definition of $\delta_{\tau}Q(\mathcal{D})$. Equation (76) follows from the fact that the delta version of Algorithm 1 ensures the following: for every timestamp $\tau$, there can be at most two tuples in $\overline{\mathcal{D}}$ where $\tau$ appears as an end point of an interval. The two tuples must have the form $([\tau,\tau],\bm{t})$ and $([\alpha,\beta],\bm{t})$ for the same $\bm{t}$ where either $\alpha=\tau$ or $\beta=\tau$. When we take the canonical partition, both tuples are going to produce the same tuple in $\widehat{Q}_{\delta_{\tau}}(\widehat{\mathcal{D}}^{(\tau)})$. ∎
D.3. Upper bound in terms of $|\mathcal{D}|$ instead of $N$
The upper bound in Theorem 3.4 is written in terms of the length $N$ of the update stream. In this section, we prove Corollary 3.5, which gives a stronger bound where $N$ is replaced by the current database size $|\mathcal{D}^{(\tau)}|$ at the $\tau$-th update. Note that $|\mathcal{D}^{(\tau)}|\leq N$ and $|\mathcal{D}^{(\tau)}|$ could be unboundedly smaller than $N$ (specifically when the stream contains many deletes before the timestamp $\tau$).
Corollary 3.5.
[…]
Having an amortized update time of $\widetilde{\mathcal{O}}(|\mathcal{D}^{(\tau)}|^{\mathsf{w}(\widehat{Q})-1})$ means that all $N$ updates combined take time $\widetilde{\mathcal{O}}(\sum_{\tau\in[N]}|\mathcal{D}^{(\tau)}|^{\mathsf{w}(\widehat{Q})-1})$.
Proof.
The proof uses Theorem 3.4 as a black box. In particular, we will use the same maintenance algorithm from the proof of Theorem 3.4. However, every now and then, we will “reset” the algorithm by taking the current database $\mathcal{D}^{(\tau)}$, creating a new empty database from scratch, and inserting every tuple in $\mathcal{D}^{(\tau)}$ into the new database as a single-tuple update. After that, we process the next $\frac{1}{2}|\mathcal{D}^{(\tau)}|$ updates [^5] For simplicity, we assume that $|\mathcal{D}^{(\tau)}|$ is a positive even integer. from the original update sequence, and then we apply another reset. The total number of updates between the last two resets is thus $\frac{3}{2}|\mathcal{D}^{(\tau)}|$. By Theorem 3.4, these $\frac{3}{2}|\mathcal{D}^{(\tau)}|$ updates together take time $\widetilde{\mathcal{O}}(|\mathcal{D}^{(\tau)}|^{\mathsf{w}(\widehat{Q})})$. Moreover, note that between the two resets, the database size cannot change by more than a factor of 2. This is because we only have $\frac{1}{2}|\mathcal{D}^{(\tau)}|$ updates from the original stream that are applied to $\mathcal{D}^{(\tau)}$. In the two extreme cases, all of these updates are inserts or all of them are deletes.
Consider the original update stream of length $N$. Initially, $|\mathcal{D}^{(0)}|=0$. While using the above scheme to process the stream, suppose that we end up performing $k$ resets in total at time stamps $\tau_{1}\stackrel{{\scriptstyle\text{def}}}{{=}}0<\tau_{2}<\cdots<\tau_{k}\leq N$. WLOG assume that $\tau_{k}=N$. From the above construction, for any $i\in[k-1]$, we have the following:
(77) $\tau_{i+1}-\tau_{i}$ $=\frac{1}{2}|\mathcal{D}^{(\tau_{i})}|$
(78) $\frac{1}{2}|\mathcal{D}^{(\tau_{i})}|$ $\leq|\mathcal{D}^{(\tau)}|\leq\frac{3}{2}|\mathcal{D}^{(\tau_{i})}|,\quad\forall\tau\in[\tau_{i},\tau_{i+1})$
Eq. (78) holds because between $\tau_{i}$ and $\tau_{i+1}$, we are making $\frac{1}{2}|\mathcal{D}^{(\tau_{i})}|$ updates to $\mathcal{D}^{(\tau_{i})}$, which in the two extreme cases are either all inserts or all deletes. As noted before, from the beginning of the $\tau_{i}$-reset until the beginning of the $\tau_{i+1}$-reset, we process $\frac{3}{2}|\mathcal{D}^{(\tau_{i})}|$ updates, which together take time $\widetilde{\mathcal{O}}(|\mathcal{D}^{(\tau_{i})}|^{\mathsf{w}(\widehat{Q})})$, thanks to Theorem 3.4. To bound the total runtime, we sum up over $i\in[k-1]$: (We don’t actually need to perform the last reset at time $\tau_{k}=N$ since there are no subsequent updates anyhow.)
(79) $\sum_{i\in[k-1]}|\mathcal{D}^{(\tau_{i})}|^{\mathsf{w}(\widehat{Q})}$
(80) $=$ $\sum_{i\in[k-1]}\sum_{\tau\in[\tau_{i},\tau_{i+1})}\frac{1}{\tau_{i+1}-\tau_{i}}\cdot|\mathcal{D}^{(\tau_{i})}|^{\mathsf{w}(\widehat{Q})}$
(81) $=$ $2\cdot\sum_{i\in[k-1]}\sum_{\tau\in[\tau_{i},\tau_{i+1})}|\mathcal{D}^{(\tau_{i})}|^{\mathsf{w}(\widehat{Q})-1}$
(82) $\leq$ $2^{\mathsf{w}(\widehat{Q})}\cdot\sum_{i\in[k-1]}\sum_{\tau\in[\tau_{i},\tau_{i+1})}|\mathcal{D}^{(\tau)}|^{\mathsf{w}(\widehat{Q})-1}$
(83) $=$ $\mathcal{O}(1)\cdot\sum_{\tau\in[N]}|\mathcal{D}^{(\tau)}|^{\mathsf{w}(\widehat{Q})-1}$
Equality (81) follows from Eq. (77), while inequality (82) follows from Eq. (78). In particular, the total time needed to process all $N$ updates in the stream is $\widetilde{\mathcal{O}}(\sum_{\tau\in[N]}|\mathcal{D}^{(\tau)}|^{\mathsf{w}(\widehat{Q})-1})$. This implies that the amortized time per update is $\widetilde{\mathcal{O}}(|\mathcal{D}^{(\tau)}|^{\mathsf{w}(\widehat{Q})-1})$, as desired. ∎
D.4. Lower Bound for $\text{IVM}^{\pm}_{\delta}[Q]$
The goal of this section is to prove Theorem 3.6, which gives a lower bound on the $\text{IVM}^{\pm}_{\delta}$ update time of a query $Q$ in terms of a lower bound on the static evaluation time for any component $\widehat{Q}_{\bm{\sigma}}$ in its multivariate extension. To that end, we give a reduction from the static evaluation of $\widehat{Q}_{\bm{\sigma}}$ to the $\text{IVM}^{\pm}_{\delta}$ problem for $Q$. Example 7.2 introduces the main idea behind this reduction.
For the rest of this section, we fix a query $Q$, its time extension $\overline{Q}$ (Definition 6.2), and one component $\widehat{Q}_{\bm{\sigma}}$ in its multivariate extension $\widehat{Q}$ (Definition 2.3). Moreover, we fix an arbitrary database instance $\widehat{\mathcal{D}}_{\bm{\sigma}}$ for $\widehat{Q}_{\bm{\sigma}}$. Algorithm 2 describes our reduction from $\text{Eval}[\widehat{Q}_{\bm{\sigma}}]$ to $\text{IVM}^{\pm}_{\delta}[Q]$. Theorem C.9 already gives a reduction from $\text{Eval}[\widehat{Q}_{\bm{\sigma}}]$ to the (static) evaluation of $\overline{Q}$ over some database instance $\overline{\mathcal{D}}\stackrel{{\scriptstyle\text{def}}}{{=}}\text{\sf IV}(\widehat{\mathcal{D}}_{\bm{\sigma}})$ that is given by Definition C.7 and satisfies the following properties:
$|\overline{\mathcal{D}}|$ $=|\widehat{\mathcal{D}}_{\bm{\sigma}}|$
$H_{k}(\overline{Q}(\overline{\mathcal{D}}))$ $=\widehat{Q}_{\bm{\sigma}}(\widehat{\mathcal{D}}_{\bm{\sigma}})$
$|\overline{Q}(\overline{\mathcal{D}})|$ $=|\widehat{Q}_{\bm{\sigma}}(\widehat{\mathcal{D}}_{\bm{\sigma}})|$
(Recall the definition of $H_{k}$ from Theorem C.9.) The above properties follow from Proposition C.8 and Theorem C.9, which also imply that $\overline{\mathcal{D}}$ can be constructed in linear time in $|\widehat{\mathcal{D}}_{\bm{\sigma}}|$, and $H_{k}(\overline{Q}(\overline{\mathcal{D}}))$ can be computed in linear time in $|\overline{Q}(\overline{\mathcal{D}})|=|\widehat{Q}_{\bm{\sigma}}(\widehat{\mathcal{D}}_{\bm{\sigma}})|$. It remains to show how to reduce the (static) evaluation of $\overline{Q}$ over $\overline{\mathcal{D}}$ into $\text{IVM}^{\pm}_{\delta}[Q]$. This is shown by the following Lemma.
Algorithm 2 Reduction from $\text{Eval}[\widehat{Q}_{\bm{\sigma}}]$ to $\text{IVM}^{\pm}_{\delta}[Q]$
Inputs
  • A query $\widehat{Q}_{\bm{\sigma}}$triangleright Equation (2)
  • An arbitrary database instance $\widehat{\mathcal{D}}_{\sigma}$ for $\widehat{Q}_{\bm{\sigma}}$
Algorithm
  • Construct $\overline{Q}$triangleright Equation (10)
  • $\overline{\mathcal{D}}\leftarrow\text{\sf IV}(\widehat{\mathcal{D}}_{\bm{\sigma}})$triangleright Definition C.7
  • Initialize $\mathcal{D}\stackrel{{\scriptstyle\text{def}}}{{=}}\left(R_{j}\right)_{j\in[k]}$ to be empty $\triangleright$ Database for $Q$
  • $\triangleright$ Sort tuples in $\overline{\mathcal{D}}$ by their end points
    $L\leftarrow$$\{(\alpha,([\alpha,\beta],\bm{t}_{j}))\mid([\alpha,\beta],\bm{t}_{j})\in\overline{R}_{j}\text{ for some $j\in[k]$}\}$
    $\cup$ $\{(\beta,([\alpha,\beta],\bm{t}_{j}))\mid([\alpha,\beta],\bm{t}_{j})\in\overline{R}_{j}\text{ for some $j\in[k]$}\}$
  • $\Delta\leftarrow\emptyset$
  • For each $(\tau,([\alpha,\beta],\bm{t}_{j}))\in L$ ordered by $\tau$:
    • if $\tau=\alpha$,
      • insert $\bm{t}_{j}$ into $R_{j}$
    • Otherwise,
      • delete $\bm{t}_{j}$ from $R_{j}$
    • $\Delta\leftarrow\Delta\cup(\{\tau\}\times\delta_{\tau}Q(\mathcal{D}))$$\triangleright$ Invoke constant-delay enumeration from $\text{IVM}^{\pm}_{\delta}[Q]$ to get $\delta_{\tau}Q(\mathcal{D})$
  • $\overline{Q}(\overline{\mathcal{D}})\leftarrow\{([\tau,\tau^{\prime}],\bm{t})\mid(\tau,+Q(\bm{t}))\in\Delta\wedge(\tau^{\prime},-Q(\bm{t}))\in\Delta\}$
  • return $\widehat{Q}_{\bm{\sigma}}(\widehat{\mathcal{D}}_{\bm{\sigma}})=H_{k}(\overline{Q}(\overline{\mathcal{D}}))$triangleright Theorem C.9
Lemma D.9 ().
[…]
Assuming Lemma D.9 holds, Theorem 3.6 follows by taking $f(N)=N^{\omega(\widehat{Q}_{\bm{\sigma}})-1-\gamma}$ for some constant $\gamma>0$. (Recall Definition 2.2.) By Lemma D.9, this would lead into an algorithm for $\text{Eval}[\overline{Q}]$ with runtime $\widetilde{\mathcal{O}}(|\overline{\mathcal{D}}|^{\omega(\widehat{Q}_{\bm{\sigma}})-\gamma}+|\overline{Q}(\overline{\mathcal{D}})|)$, and by extension, an algorithm for $\text{Eval}[\widehat{Q}_{\bm{\sigma}}]$ with runtime $\widetilde{\mathcal{O}}(|\widehat{\mathcal{D}}_{\bm{\sigma}}|^{\omega(\widehat{Q}_{\bm{\sigma}})-\gamma}+|\widehat{Q}_{\bm{\sigma}}(\widehat{\mathcal{D}}_{\bm{\sigma}})|)$. This would contradict the definition of $\omega(\widehat{Q}_{\bm{\sigma}})$, thus proving that our choice of $f(N)$ is not possible. Next, we prove Lemma D.9.
Proof of Lemma D.9.
Following Algorithm 2, the reduction works as follows. We collect a list $L$ of all pairs $(\alpha,([\alpha,\beta],\bm{t}_{j}))$ and $(\beta,([\alpha,\beta],\bm{t}_{j}))$ for every tuple $([\alpha,\beta],\bm{t}_{j})$ that occurs in any relation $\overline{R}_{j}$ in $\overline{\mathcal{D}}$. We go through pairs $(\tau,([\alpha,\beta],\bm{t}_{j}))$ in the list $L$ in order of $\tau$. If $\tau=\alpha$, then we insert $\bm{t}_{j}$ into $R_{j}$. Otherwise, we delete $\bm{t}_{j}$ from $R_{j}$. In total, we have $2|\overline{\mathcal{D}}|$ updates and the amortized time per update is $f(2|\overline{\mathcal{D}}|)$.
After every insert or delete, we invoke the constant-delay enumeration procedure for $\text{IVM}^{\pm}_{\delta}[Q]$ in order to retrieve $\delta_{\tau}Q(\mathcal{D})$, i.e. the output tuples of $Q$ that are inserted or deleted due to the insert or delete in $R_{j}$. Whenever constant-delay enumeration reports a delete $-Q(\bm{t})$ at time $\tau^{\prime}$ where the the same tuple $\bm{t}$ was reportedly inserted $+Q(\bm{t})$ at a previous time $\tau$, we report the tuple $([\tau,\tau^{\prime}],\bm{t})$ as an output tuple of $\overline{Q}$. The overall time we spend on enumeration is $\widetilde{\mathcal{O}}(|\overline{Q}(\overline{\mathcal{D}})|)$. ∎
We conclude this section with the following remark about the prospects of extending Theorem 3.6 from $\text{IVM}^{\pm}_{\delta}[Q]$ to $\text{IVM}^{\pm}[Q]$.
Remark D.6 ().
It is not clear how to extend the proof of Theorem 3.6, which holds for the $\text{IVM}^{\pm}_{\delta}[Q]$ problem, so that it also works for the $\text{IVM}^{\pm}[Q]$ problem. The main issue lies in the enumeration step. In particular, suppose that the IVM oracle that we have in the proof is for $\text{IVM}^{\pm}[Q]$. Per the above reduction, we invoke constant-delay enumeration after every update in $Q$ in order to extract the output of $\overline{Q}$. However, now the oracle enumerates the full output of $Q$ after every update. As a result, the same output tuple of $Q$ might be reported many times throughout its lifetime. This in turn would prevent us from bounding the total enumeration time with $\widetilde{\mathcal{O}}(|\overline{Q}(\overline{\mathcal{D}})|)$, as we did in the original proof. Proving an analog of Theorem 3.6 for $\text{IVM}^{\pm}[Q]$ remains an open problem.
Appendix E Missing Details in Section 3
In this section, we first prove the following statements:
Proposition E.1 ().
A query $Q$ is hierarchical if and only if its multivariate extension $\widehat{Q}$ is acyclic.
Proposition E.2 ().
For any non-hierarchical query $Q$, $\mathsf{w}(\widehat{Q})\geq\frac{3}{2}$.
Proposition E.3 ().
For any Loomis-Whitney query $Q$, $\mathsf{w}(\widehat{Q})=\frac{3}{2}$.
Proposition E.4 ().
For any query $Q$, $\mathsf{w}(Q)\leq\mathsf{w}(\widehat{Q})\leq\mathsf{w}(Q)+1.$
Proposition E.5 ().
[…]
E.1. Proof of Proposition E.1
Proposition E.1.
A query is hierarchical if and only if its multivariate extension is acyclic.
It follows from Proposition E.2 that the multivariate extension $\widehat{Q}$ of any non-hierarchical query $Q$ is not acyclic. This is because acyclic queries have a fractional hypertree width of $1$. It remains to prove:
Lemma E.6 ().
The multivariate extension of any hierarchical query is acyclic.
First, we introduce some simple notions. Consider a query
$Q=R_{1}(\bm{X}_{1})\wedge\cdots\wedge R_{k}(\bm{X}_{k}).$
A variable $X$ is called root variable if it is contained in every atom, i.e., $X\in\bigcap_{i\in[k]}\bm{X}_{i}$. It is a leaf variable if it is contained in exactly one atom, i.e, $X\in\bm{X}_{i}$ for some $i\in[k]$ and $X\notin\bm{X}_{j}$ for all $j\in[k]\setminus\{i\}$. We say that an atom $R_{i}(\bm{X}_{i})$ is subsumed by another atom $R_{j}(\bm{X}_{j})$ if $\bm{X}_{i}\subseteq\bm{X}_{j}$. A query $Q^{\prime}$ results from $Q$ by the elimination of variable $X$ if it is of the form
$Q^{\prime}=R_{1}(\bm{X}_{1}\setminus\{X\})\wedge\cdots\wedge R_{k}(\bm{X}_{k}\setminus\{X\}).$
It results by the elimination of an atom $R_{i}(\bm{X}_{i})$ if it has the form
$Q^{\prime}=R_{1}(\bm{X}_{1})\wedge\cdots\wedge R_{i-1}(\bm{X}_{i-1})\wedge R_{i+1}(\bm{X}_{i+1})\wedge\cdots\wedge R_{k}(\bm{X}_{k}).$
Observe that the elimination of a leaf variable and the elimination of a subsumed atom are exactly the two steps of the GYO-reduction (Yu and Ozsoyoglu, 1979). The following claim follows directly from the definition of acyclic queries and the correctness of the GYO-reduction:
Claim E.1 ().
Let $Q$ and $Q^{\prime}$ be join queries such that $Q^{\prime}$ results from $Q$ by the elimination of a root variable, a leaf variable, or a subsumed atom. The query $Q$ is acyclic if and only if the query $Q^{\prime}$ is acyclic.
Proof of Claim E.1.
We distinguish on how $Q^{\prime}$ is obtained from $Q$.
$Q^{\prime}$ is obtained by the elimination of a leaf variable or a subsumed atom
In this case, the claim follows directly from the correctness of the GYO-reduction (Yu and Ozsoyoglu, 1979), which says that a query is acyclic if and only if we obtain the trivial query consisting of a single atom with empty schema by repeatedly applying leaf variable and subsumed atom elimination. Indeed, $Q$ is GYO-reducible if and only if $Q^{\prime}$ is GYO-reducible.
$Q^{\prime}$ is obtained by root variable elimination
Assume that $Q^{\prime}$ is obtained from $Q$ by removing a root variable $X$. Consider a join tree $\mathcal{T}$ for $Q$, and let $\mathcal{T}^{\prime}$ be the tree obtained by removing the variable $X$ from every node in $\mathcal{T}$. There is a one-to-one mapping between the nodes in $\mathcal{T}^{\prime}$ and the atoms in $Q^{\prime}$. Moreover, removing $X$ from all nodes in $\mathcal{T}$ does not affect the connectivity property in the join tree. Hence, $\mathcal{T}^{\prime}$ must be a join tree for $Q^{\prime}$.
Analogously, any join tree for $Q^{\prime}$ turns into a join tree for $Q$ by adding the same variable $X$ to all nodes. We conclude that $Q$ is acyclic if and only if $Q^{\prime}$ is acyclic. ∎
We are ready to prove Lemma E.6. Consider a hierarchical query $Q=R_{1}(\bm{X}_{1})\wedge\cdots\wedge R_{k}(\bm{X}_{k})$ and its multivariate extension $\widehat{Q}=\bigvee_{\bm{\sigma}\in\Sigma_{k}}\widehat{Q}_{\bm{\sigma}}$ (Eq (3)). We show that each component $\widehat{Q}_{\bm{\sigma}}$ of $\widehat{Q}$ is acyclic. WLOG we only consider the component for $\bm{\sigma}=(1,\ldots,k)$:
$\widehat{Q}_{\bm{\sigma}}=\widehat{R}_{1}(Z_{1}\bm{X}_{1})\wedge\widehat{R}_{2}(Z_{1},Z_{2},\bm{X}_{2})\wedge\cdots\wedge\widehat{R}_{k}(Z_{1},\ldots,Z_{k},\bm{X}_{k}).$
The other components are symmetric.
The proof is by induction on the number $k$ of atoms in $\widehat{Q}_{\bm{\sigma}}$.
Base Case
Assume that $k=1$, that is, $\widehat{Q}_{\bm{\sigma}}$ consists of a single atom. In this case, the query is obviously acyclic.
Induction Step
Assume now that $\widehat{Q}_{\bm{\sigma}}$ has $k>1$ atoms. The variable $Z_{1}$ is a root variable. We eliminate it and obtain
$\widehat{Q}_{\bm{\sigma}}^{\prime}=\widehat{R}_{1}^{\prime}(\bm{X}_{1})\wedge\widehat{R}_{2}^{\prime}(Z_{2},\bm{X}_{2})\wedge\cdots\wedge\widehat{R}_{k}^{\prime}(Z_{2},\ldots,Z_{k},\bm{X}_{k}).$
The atom $\widehat{R}_{1}^{\prime}(\bm{X}_{1})$ has now the same schema as the atom ${R}_{1}(\bm{X}_{1})$ in the original hierarchical query $Q$. We remove from $\widehat{R}_{1}^{\prime}(\bm{X}_{1})$ all variables that are leaves in $\widehat{Q}_{\bm{\sigma}}^{\prime}$ and obtain the atom $\widehat{R}_{1}^{\prime\prime}(\bm{X}_{1}^{\prime})$. It follows from the definition of hierarchical queries that the schema $\bm{X}_{1}^{\prime}$ must be subsumed by the schema of one of the atoms $\widehat{R}_{2}^{\prime}(Z_{2},\bm{X}_{2}),$ $\ldots$,… , $\widehat{R}_{k}^{\prime}(Z_{2},\ldots,Z_{k},\bm{X}_{k})$. We remove the atom $\widehat{R}_{1}^{\prime\prime}(\bm{X}_{1}^{\prime})$ from $\widehat{Q}_{\bm{\sigma}}^{\prime}$ and obtain the query
$\widehat{Q}_{\bm{\sigma}}^{\prime\prime}=\widehat{R}_{2}^{\prime}(Z_{2},\bm{X}_{2})\wedge\cdots\wedge\widehat{R}_{k}^{\prime}(Z_{2},\ldots,Z_{k},\bm{X}_{k}).$
This query has $k-1$ atoms and contains the variables $Z_{2},\ldots,Z_{k}$ but not $Z_{1}$. Observe that $\widehat{Q}_{\bm{\sigma}}^{\prime\prime}$ is the multivariate extension of a query $Q^{\prime}$ that results from $Q$ by skipping the atom $R_{1}(\bm{X}_{1})$. Hierarchical queries remain hierarchical under atom elimination. Hence, $Q^{\prime}$ is hierarchical. By induction hypothesis, $\widehat{Q}_{\bm{\sigma}}^{\prime\prime}$ is acyclic. Since we obtained $\widehat{Q}_{\bm{\sigma}}^{\prime\prime}$ from $\widehat{Q}_{\bm{\sigma}}$ by eliminating root variables, leaf variables, or subsumed atoms, it follows from Claim E.1 that $\widehat{Q}_{\bm{\sigma}}$ must be acyclic. This completes the induction step.
E.2. Proof of Proposition E.2
Proposition E.2.
For any non-hierarchical query $Q$, $\mathsf{w}(\widehat{Q})\geq\frac{3}{2}$.
Proof.
Consider a non-hierarchical query $Q$. By definition, $Q$ must contain two variables $X$ and $Y$ and three atoms $R_{1}(\bm{X}_{1})$, $R_{2}(\bm{X}_{2})$, and $R_{3}(\bm{X}_{3})$ such that
  • $X\in\bm{X}_{1}$ but $X\notin\bm{X}_{3}$,
  • $Y\in\bm{X}_{3}$ but $Y\notin\bm{X}_{1}$, and
  • $X,Y\in\bm{X}_{2}$.
The multivariate extension $\widehat{Q}$ of $Q$ must have a component $\widehat{Q}_{\bm{\sigma}}$ that contains three atoms of the form
$\widehat{R}_{1}(Z_{1},\ldots,Z_{k-1},\bm{X}_{1}),\widehat{R}_{2}(Z_{1},\ldots,Z_{k-2},\bm{X}_{2}),\widehat{R}_{3}(Z_{1},\ldots,Z_{k},\bm{X}_{3})$
where $\{Z_{1},\ldots,Z_{k}\}$ are fresh variables. Moreover, no other atoms in $\widehat{Q}_{\bm{\sigma}}$ contains any of the the three variables $\{Z_{k-2},Z_{k-1},Z_{k}\}$. Let $\bm{Y}=\{X,Y,Z_{k-1}\}$. When restricted to the variables in $\bm{Y}$, the resulting query, which we call $\widehat{Q}_{\bm{\sigma}}^{\prime}$, is a triangle query. In particular, let $\widehat{Q}_{\bm{\sigma}}^{\prime}\stackrel{{\scriptstyle\text{def}}}{{=}}(\widehat{Q}_{\bm{\sigma}})_{\bm{Y}}$, where the notation $(\widehat{Q}_{\bm{\sigma}})_{\bm{Y}}$ was defined in (14). Then,
$\widehat{Q}_{\bm{\sigma}}^{\prime}=\widehat{R}_{1}(Z_{k-1},X),\widehat{R}_{2}(X,Y),\widehat{R}_{3}(Z_{k-1},Y)$
Therefore, $\rho^{*}(\widehat{Q}_{\bm{\sigma}}^{\prime})=3/2$. We use this fact to to show that $\mathsf{w}(\widehat{Q}_{\bm{\sigma}})$ must be at least $3/2$. To that end, we use variable elimination orders (see Appendix A). In particular, we use the definition of fractional hypertree width in terms of variable elimination orders, as given by Eq. (16). Let $\bm{\mu}$ be an optimal variable elimination order for $\widehat{Q}_{\bm{\sigma}}$, i.e., $\mathsf{w}(\bm{\mu})=\mathsf{w}(\widehat{Q}_{\bm{\sigma}})$. We make the following claim:
Claim E.2 ().
There must be a variable $W\in\bm{\mu}$ where $\bm{U}_{W}^{\bm{\mu}}\supseteq\bm{Y}$.
We prove the above claim as follows. Initially, the restricted query $(\widehat{Q}_{\bm{\sigma}})_{\bm{Y}}$ is a triangle query. Every time we eliminate a variable $W$ that does not satisfy $\bm{U}_{W}^{\bm{\mu}}\supseteq\bm{Y}$, the restricted query $(\widehat{Q}_{\bm{\sigma}})_{\bm{Y}}$ remains a triangle query. Hence, elimination could not have terminated because we haven’t yet eliminated any of the variables in $\bm{Y}$. Therefore, there must be a variable $W\in\bm{\mu}$ where $\bm{U}_{W}^{\bm{\mu}}\supseteq\bm{Y}$. Now, we have:
$\mathsf{w}(\widehat{Q}_{\bm{\sigma}})=\mathsf{w}(\bm{\mu})\geq\rho^{*}(\bm{U}_{W}^{\bm{\mu}})\geq\rho^{*}(\bm{Y})=3/2.$
The first equality above follows from our assumption that $\bm{\mu}$ is an optimal variable elimination order for $\widehat{Q}_{\bm{\sigma}}$. The second inequality above follows from the monotonicity of the fractional edge cover number $\rho^{*}$ (see Appendix A). ∎
E.3. Proof of Proposition E.3
Proposition E.3.
For any Loomis-Whitney query $Q$, $\mathsf{w}(\widehat{Q})=\frac{3}{2}$.
Consider an arbitrary Loomis-Whitney query $Q$ of some degree $k\geq 3$ and its multivariate extension $\widehat{Q}$ with fresh variables $Z_{1},\ldots,Z_{k}$. Proposition E.2 implies that the fractional hypertree width of $\widehat{Q}$ is at least $\frac{3}{2}$. Let $\widehat{Q}_{\bm{\sigma}}$ be an arbitrary component of $\widehat{Q}$. Next, we show that the fractional hypertree width of $\widehat{Q}_{\bm{\sigma}}$ is at most $\frac{3}{2}$, which concludes the proof.
Let $\widehat{Q}_{\bm{\sigma}}^{\prime}$ be the query that we obtain from $\widehat{Q}_{\bm{\sigma}}$ by eliminating the variable $Z_{k}$. Since this variable occurs in only a single atom in $\widehat{Q}_{\bm{\sigma}}$, the fractional hypertree width of $\widehat{Q}_{\bm{\sigma}}$ and $\widehat{Q}_{\bm{\sigma}}^{\prime}$ must be the same. This follows from the equivalence between variable elimination orders and tree decompositions, as explained in Section A. It remains to show that $\widehat{Q}_{\bm{\sigma}}^{\prime}$ has a fractional hypertree width of $\frac{3}{2}$. By definition of a multivariate extension, $\widehat{Q}_{\bm{\sigma}}^{\prime}$ must contain three atoms of the form:
$\widehat{R}_{1}(Z_{1},\ldots,Z_{k-2},\bm{X}_{1}),\widehat{R}_{2}(Z_{1},\ldots,Z_{k-1},\bm{X}_{2}),\widehat{R}_{3}(Z_{1},\ldots,Z_{k-1},\bm{X}_{3})$
Note that each of the variables $\{Z_{1},\ldots,Z_{k-1}\}$ is contained in at least two of the atoms $\widehat{R}_{1},\widehat{R}_{2},\widehat{R}_{3}$ above. Moreover, because $\bm{X}_{1},\bm{X}_{2},\bm{X}_{3}$ are the schemas of three atoms in a Loomis-Whitney query $Q$, each variable $X\in\text{\sf vars}(Q)$ must occur in at least two of the three atoms $\widehat{R}_{1},\widehat{R}_{2},\widehat{R}_{3}$ above. Consider now a fractional edge cover $\lambda_{\widehat{R}}=\lambda_{\widehat{S}}=\lambda_{\widehat{T}}=\frac{1}{2}$. This is indeed a valid fractional edge cover for $\widehat{Q}_{\bm{\sigma}}^{\prime}$, and it proves that $\rho^{*}(\widehat{Q}_{\bm{\sigma}}^{\prime})\leq\frac{3}{2}$. This implies that $\mathsf{w}(\widehat{Q}_{\bm{\sigma}})\leq\frac{3}{2}$.
E.4. Proof of Proposition E.4
Proposition E.4.
For any query $Q$, $\mathsf{w}(Q)\leq\mathsf{w}(\widehat{Q})\leq\mathsf{w}(Q)+1$.
We consider a query $Q=R_{1}(\bm{X}_{1})\wedge\ldots\wedge R_{k}(\bm{X}_{k})$ and its multivariate extension $\widehat{Q}=\bigvee_{\bm{\sigma}\in\Sigma_{k}}\widehat{Q}_{\bm{\sigma}}$ with fresh variables $Z_{1},\ldots,Z_{k}$ (Eq. (3)).
Proof of $\mathsf{w}(Q)\leq\mathsf{w}(\widehat{Q})$
For the sake of contradiction, assume that $\mathsf{w}(Q)>\mathsf{w}(\widehat{Q})$. Consider an arbitrary component $\widehat{Q}_{\bm{\sigma}}$ of $\widehat{Q}$. It follows from the definition of the fractional hypertree width that $\mathsf{w}(\widehat{Q}_{\bm{\sigma}})\leq\mathsf{w}(\widehat{Q})$. Hence, $\mathsf{w}(Q)>\mathsf{w}(\widehat{Q}_{\bm{\sigma}})$. Let $(\mathcal{T}_{\bm{\sigma}},\chi_{\bm{\sigma}})$ be a tree decomposition of $\widehat{Q}_{\bm{\sigma}}$ with $\mathsf{w}(\mathcal{T}_{\bm{\sigma}},\chi_{\bm{\sigma}})$ $=$ $\mathsf{w}(\widehat{Q}_{\bm{\sigma}})$. We will construct from $(\mathcal{T}_{\bm{\sigma}},\chi_{\bm{\sigma}})$ a tree decomposition $(\mathcal{T},\chi)$ for $Q$ with $\mathsf{w}(\mathcal{T},\chi)\leq\mathsf{w}(\mathcal{T}_{\bm{\sigma}},\chi_{\bm{\sigma}})$. This implies that $\mathsf{w}(\mathcal{T},\chi)\leq\mathsf{w}(\widehat{Q})$, hence $\mathsf{w}(Q)\leq\mathsf{w}(\widehat{Q})$, which is a contradiction.
Without loss of generality, assume that $\widehat{Q}_{\bm{\sigma}}$ is of the form $\widehat{Q}_{\bm{\sigma}}=\widehat{R}_{1}(Z_{1},\bm{X}_{1}),\ldots,\widehat{R}_{k}(Z_{1},\ldots,Z_{k},\bm{X}_{k})$. For all other components of $\widehat{Q}$, the argument is completely analogous. Let $\widehat{B}_{1},\ldots,\widehat{B}_{n}$ be the bags of $(\mathcal{T}_{\bm{\sigma}},\chi_{\bm{\sigma}})$. The pair $(\mathcal{T},\chi)$ results from $(\mathcal{T}_{\bm{\sigma}},\chi_{\bm{\sigma}})$ by eliminating the variables $Z_{1},\ldots,Z_{k}$ from all bags. Let $B_{1},\ldots,B_{n}$ denote the bags of $(\mathcal{T},\chi)$. The pair $(\mathcal{T},\chi)$ constitutes a tree decomposition for $Q$, since (a) for each atom $R_{i}(\bm{X}_{i})$ of $Q$ there must still be one bag that contains $\bm{X}_{i}$ and (b) for each variable $X\in\bigcup_{i\in[k]}\bm{X}_{i}$ the bags in $(\mathcal{T},\chi)$ containing $X$ build a connected subtree. Given $i\in[n]$, let $\widehat{Q}_{\bm{\sigma}}^{i}$ be the query that we obtain from $\widehat{Q}_{\bm{\sigma}}$ by restricting the schema of each atom to the variables in $\widehat{B}_{i}$. Similarly, let $Q^{i}$ be the query obtained from $Q$ by restricting the schema of each atom to the variables in the bag $B_{i}$. Consider now a fractional edge cover $\big{(}\lambda_{\widehat{R}_{i}(\bm{X}_{i})}\big{)}_{i\in[k]}$ for $\widehat{Q}_{\bm{\sigma}}^{i}$. This tuple is also a fractional edge cover for $Q^{i}$, since $B_{i}\subseteq\widehat{B}_{i}$. This means that $\rho^{*}(Q^{i})\leq\rho^{*}(\widehat{Q}_{\bm{\sigma}}^{i})$. Hence, $\mathsf{w}(\mathcal{T},\chi)$leq $\mathsf{w}(\mathcal{T}_{\bm{\sigma}},\chi_{\bm{\sigma}})$.
Proof of $\widehat{\mathsf{w}}(Q)\leq\mathsf{w}(Q)+1$
Let $\widehat{Q}_{\bm{\sigma}}$ be an arbitrary component of $\widehat{Q}$. Consider a tree decomposition $(\mathcal{T},\chi)$ of $Q$ with fractional hypertree width $\mathsf{w}$. We will transform $(\mathcal{T},\chi)$ into a tree decomposition $(\mathcal{T}_{\bm{\sigma}},\chi_{\bm{\sigma}})$ for $\widehat{Q}_{\bm{\sigma}}$ with fractional hypertree width of at most $\mathsf{w}(Q)+1$. This implies that $\mathsf{w}(\widehat{Q})$ is at most $\mathsf{w}(Q)+1$.
Without loss of generality, assume that $\widehat{Q}_{\bm{\sigma}}$ is of the form $\widehat{Q}_{\bm{\sigma}}=\widehat{R}_{1}(Z_{1},\bm{X}_{1}),\ldots,\widehat{R}_{k}(Z_{1},\ldots,Z_{k},\bm{X}_{k})$. For other components of $\widehat{Q}$, the argument is analogous. Let $B_{1},\ldots,B_{n}$ be the bags of $(\mathcal{T},\chi)$. The pair $(\mathcal{T}_{\bm{\sigma}},\chi_{\bm{\sigma}})$ results from $(\mathcal{T},\chi)$ by adding to each bag the variables $Z_{1},\ldots,Z_{k}$. Let $\widehat{B}_{1}=B_{1}\cup\{Z_{1},\ldots,Z_{k}\},\ldots,\widehat{B}_{n}\cup\{Z_{1},\ldots,Z_{k}\}$ be the bags of $(\mathcal{T}_{\bm{\sigma}},\chi_{\bm{\sigma}})$. By construction, $(\mathcal{T}_{\bm{\sigma}},\chi_{\bm{\sigma}})$ constitutes a tree decomposition for $\widehat{Q}_{\bm{\sigma}}$, since (a) for each $i\in[k]$, the variables $\{Z_{1},\ldots,Z_{i}\}\cup\bm{X}_{i}$ must be included in one bag and (b) for each variable $X\in\{Z_{1},\ldots,Z_{k}\}\cup\bigcup_{i\in[k]}\bm{X}_{i}$, the bags that include $X$ still build a connected subtree. It remains to show that $\mathsf{w}(\mathcal{T}_{\bm{\sigma}},\chi_{\bm{\sigma}})\leq\mathsf{w}(Q)+1$. For some $i\in[n]$, let $Q^{i}$ be the query obtained from $Q$ by restricting each atom to the variables in $B_{i}$ and let $\widehat{Q}_{\bm{\sigma}}^{i}$ be the query obtained from $\widehat{Q}_{\bm{\sigma}}$ by restricting each atom to the variables in $\widehat{B}_{i}$. Consider a fractional edge cover $\big{(}\lambda_{R_{i}(\bm{X}_{i})}\big{)}_{i\in[k]}$ for $Q^{i}$. We can turn it into a fractional edge cover for $\widehat{Q}^{i}$ by setting $\lambda_{R_{k}(\bm{X}_{k})}$ to $1$. This means that $\rho^{*}(\widehat{Q}^{i}_{\bm{\sigma}})\leq\rho^{*}(Q^{i})+1$. This implies $\mathsf{w}(\mathcal{T}_{\bm{\sigma}},\chi_{\bm{\sigma}})$leq $\mathsf{w}(\mathcal{T},\chi)+1$.
E.5. Proof of Proposition E.5
Proposition E.5.
[…]
The proof is based on the Online Vector-Matrix-Vector Multiplication (OuMv) Conjecture, which is implied by the Online Matrix-Vector Multiplication (OMv) Conjecture (Henzinger et al., 2015). See Appendix A.
Consider the simplest non-hierarchical query $Q_{nh}=R(A)\wedge S(A,B)\wedge T(B)$. The proof of Proposition E.5 consists of three parts. First, we show hardness for the problem $\text{IVM}^{\pm}[Q_{nh}]$ (Section E.5.1). Afterwards, we show how the hardness proof for $\text{IVM}^{\pm}[Q_{nh}]$ extends to a hardness proof for $\text{IVM}^{\pm}_{\delta}[Q_{nh}]$ (Section E.5.2). Finally, we explain how these two results generalize to arbitrary non-hierarchical queries without self-joins (Section E.5.3). The technique used in the latter part is standard in the literature (see, e.g. (Berkholz et al., 2017)), so we only give the high-level idea.
E.5.1. Hardness of $\text{IVM}^{\pm}[Q_{\text{nh}}]$
We reduce the OuMv problem to $\text{IVM}^{\pm}[Q_{\text{nh}}]$. Assume that there is some $\gamma>0$ and an algorithm $\mathcal{A}$ that solves $\text{IVM}^{\pm}[Q_{\text{nh}}]$ with $\widetilde{\mathcal{O}}(N^{1/2-\gamma})$ amortized update time for update streams of length $N$. We show how to use $\mathcal{A}$ to solve the OuMv problem in sub-cubic time, which is a contradiction to the OuMv Conjecture. Let the $n\times n$ matrix $\bm{M}$ and the $n\times 1$ column vectors $(\bm{u}_{1},\bm{v}_{1}),\ldots,(\bm{u}_{n},\bm{v}_{n})$ be the input to the OuMv problem. For $i,j\in[n]$, we denote by $\bm{M}(i,j)$ the Boolean value in $\bm{M}$ at position $(i,j)$ and by $\bm{u}(i)$ (and $\bm{v}(i)$) the value in $\bm{u}(i)$ (and $\bm{v}(i)$) at position $i$. The idea is to use relation $S$ to encode the matrix $\bm{M}$ and the relations $R$ and $T$ to encode the vectors $\bm{u}_{r}$ and $\bm{v}_{r}$ for $r\in[n]$. The encoding closely follows prior work (Berkholz et al., 2017). Starting from the empty database, we execute at most $n^{2}$ inserts to $S$ such that $(i,j)\in S\Leftrightarrow\bm{M}(i,j)=1$. In each round $r\in[n]$, we do the following. We use at most $4n$ updates to delete the tuples in $R$ and $T$ and to insert new tuples such that $i\in\bm{R}\Leftrightarrow\bm{u}_{r}(i)=1$ and $j\in\bm{T}\Leftrightarrow\bm{v}_{r}(j)=1$. At the end of round $r$ we request the enumeration of the result of $Q_{\text{nh}}$. It holds $\bm{u}_{r}^{\top}\bm{M}\bm{v}_{r}=1$ if and only if the result of $Q_{\text{nh}}$ contains at least one tuple. So, we can stop the enumeration process after constant time. Summed up over $n$ rounds, we execute $n^{2}+4n^{2}=\mathcal{O}(n^{2})$ updates, some of which are followed by enumeration requests that do not require more than constant time. Hence, using the update mechanism of algorithm $\mathcal{A}$, the overall time to solve the OuMv problem takes $\mathcal{O}(\text{\sf polylog}(n^{2})\cdot n^{2}\cdot(n^{2})^{1/2-\gamma})$ $=$ $\mathcal{O}(\text{\sf polylog}(n^{2})\cdot n^{3-2\gamma})$. As the polylogarithmic factor $\text{\sf polylog}(n^{2})$ can be asymptotically upper-bounded by $n^{\gamma^{\prime}}$ for arbitrarily small $\gamma^{\prime}$, we conclude that the OuMv problem can be solved in time $\mathcal{O}(n^{3-\gamma^{\prime\prime}})$ for some $\gamma^{\prime\prime}>0$. This contradicts the OuMv Conjecture.
E.5.2. Hardness of $\text{IVM}^{\pm}_{\delta}[Q_{\text{nh}}]$
We reduce the OuMv problem to $\text{IVM}^{\pm}_{\delta}[Q_{\text{nh}}]$. The encoding of the matrix and the vectors is exactly as in Section E.5.1. The only difference is how we obtain the result of $\bm{u}_{r}^{\top}\bm{M}\bm{v}_{r}$ in each round $r\in[n]$. We first execute the necessary updates to encode $\bm{u}_{r}$ in $R$. Then, we delete the contents of $T$. Afterwards, we insert, one by one, each value $j$ with $\bm{v}_{r}(j)=1$ to $T$. After each insert, we request the enumeration of the delta in the result of $Q_{\text{nh}}$ and check whether the delta contains at lease one tuple. The main observation is that $\bm{u}^{\top}\bm{M}\bm{v}=1$ if and only if one of the enumeration requests in round $r$ reports at least one tuple. The overall number of updates is as in Section E.5.1. The additional cost caused by each enumeration request is constant.
E.5.3. Hardness for Arbitrary Non-Hierarchical Queries
Given any non-hierarchical query $Q$ without self-joins, we can reduce the problems $\text{IVM}^{\pm}[Q_{\text{nh}}]$ and $\text{IVM}^{\pm}_{\delta}[Q_{\text{nh}}]$to $\text{IVM}^{\pm}[Q]$ and respectively $\text{IVM}^{\pm}_{\delta}[Q]$. We give the high-level idea of the reduction. Since $Q$ is non-hierarchical, it must contain two variables $A^{\prime}$ and $B^{\prime}$ and three atoms $R^{\prime}(\bm{X}_{R})$, $S^{\prime}(\bm{X}_{S})$, and $T^{\prime}(\bm{X}_{T})$ such that $A^{\prime}\in\bm{X}_{R}$, $B^{\prime}\notin\bm{X}_{R}$, $A^{\prime},B^{\prime}\in\bm{X}_{S}$, $B^{\prime}\in\bm{X}_{T}$, and $A^{\prime}\notin\bm{X}_{T}$. We use $R^{\prime}(\bm{X}_{R})$, $S^{\prime}(\bm{X}_{S})$, and $T^{\prime}(\bm{X}_{T})$ to encode $R(A)$, $S(A,B)$, and respectively $T(B)$. While the variables $A^{\prime}$ and $B^{\prime}$ take on the $A$- and respectively $B$-values from the database for $Q_{\text{nh}}$, all other variables in $Q$ take on a fixed dummy value. Hence, we can easily translate updates and enumeration requests to $Q_{\text{nh}}$ to updates and enumeration requests to $Q$.